A Bibliography of Publications on Unicode

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

01 May 2024 Version 0.214

Title word cross-reference

#10 [DW06, The21k]. #14 [The21d]. #24 [DF06]. #29 [The21e]. #31 [The21f]. #38 [JC08, The21g]. #39 [The21l]. #40 [SD06]. #41 [Dan08]. #42 [Mul08]. #44 [DW08]. #45 [Jen08, The21h]. #46 [The21m].

$60 [Pog03]. α [PT00]. ϵ [PT00, PT00]. η [PT00]. κ [PT00]. λ [PT00]. μ [PT00]. ρ [PT00]. σ [PT00].

0 [Pog03].

1 [Cra98, Pog03, SD02, SD06]. 1-932026-01-0 [Pog03]. 1.0 [DYI^{+}04a, DYI^{+}04b, The90]. 1.1 [Dav97a, Eck97, MBH97, Uni91]. 1.2 [Gil98a]. 1.7 [Ned04]. 10646 [Bri96, Bri97a, Ksa96, Ksa97a, Ksa98a, Ksa98b, Yer03, Wu01, Wu02, Yer96, Yer98a, Zha97b]. 10646-1 [N^{+}05a, N^{+}05b, Ngô05]. 106461 [LaB00]. 11th [FCJ04]. 12th [IEE94]. 13.0 [Ano20b]. 14 [AH05]. 14.0.0 [The21c]. 16 [Ano06c, Phi01, Sch04]. 16-bit [Ano06c, Fei94]. 1641 [GD94a]. 1642 [GD94b]. 18 [Cri05]. 18030 [Mey01]. 18th [Wei07]. 1993 [N^{+}05b, SHA93]. 1995 [AGH95, BM98c, N^{+}05b, Sui98].

2 [Atk96, Fre97f, Hip94]. 2.0 [Cla97, Hoh97]. 2.1 [Fre98b]. 2.6 [Jun97]. 2000 [N^{+}05b, Ngô05]. 2002 [Ano02b]. 2004 [FCJ04]. 2010 [DD11]. 2015 [Pro15]. 2016 [Pro16]. 2044 [Yer96]. 2070 [YNAD97]. 21 [Ali05]. 2152 [GD97]. 21st [Ano02b]. 2253 [WKH97]. 2277 [Alv98]. 2278 [FP98]. 2279 [Yer98a]. 2482 [WA99, WA98a]. 2e [Lem97].

3 [Buc96a, Buc96b, Buc96c]. 3.0 [Man01]. 3.1 [DEFM01]. 3.2 [Ano01, Uni02d]. 3.5 [Ric95]. 30 [The16a]. 32 [Dav99d, Dav01e]. 3492 [Cos03]. 3629 [Yer03]. 3718 [McG04]. 373 [McC72]. 39 [The21a].

4 [Atk96]. 4.0 [N^{+}05b, Ngô05]. 4.1.0 [Uni05c]. 400 [Shi98a]. 4042 [Cri05]. 4GL [Tex98a, Tex98b].

5.0 [AKW07, BBM98, Des08]. 500 [RC98]. 512 [CL23]. 5th [USE01].

6 [Yas05]. 639 [Con02]. 69 [The21b].

7 [GD94b, GD97, J^{+}12]. 7.02 [AH12]. 7.4 [Ast15].

8 [Ano03b, Due97b, Dür97b, Jun97, KL21, Kuh99b, McC96, Phi01, Pik03, Pik12, Yer03, RGN16, Sun00b, WKH97, Yer96, Yer98a, Zif01]. 8-Bit [Bar97, Phi01]. 8601 [Tex01].

9 [Ano94c, Fen96, Cri05]. ’93 [BJ93]. ’94 [BJ94, Tan94]. 95 [Hal98a, H^{+}95, Kan95, Nie96, Nie97b, Ric95]. ’98 [SC98].

= [WG05, Yun00].

abacus [MVM09b]. ABAP [AH12, Ast15, Woo10]. Abecedarium [Fir93]. Abjad [MVM09a]. Abode [Yas05]. Abode-Japan1-6 [Yas05]. accelerating [Che97]. Access [McG04, SS98, WKH97, ME94]. Accessibility [Bre98, Lun98a, BEE95]. accordance [N^{+}05a, N^{+}05b, Ngô05]. account [Jen70]. accurate [Dic22]. Achievements [Ksa98a, Ksa98b]. Acrobat [Rub96]. Across [Ano96c, PLYZ95]. Activities [Bos97a, Bos97b]. ADAMAS [RM16]. Adaptive [PB98]. Added [Oet12]. Address [Bos98, Bre98, Nak98, Sch97]. adoption [Che97]. Advanced [Hel92, Hel93, KH98, Ric94, Ric95, Uns96]. Advancement [Tan94]. Advantages [Rob10]. Advent [MH95]. Adventure [Lei96a, Lei96e, Lei96b]. Adventures [Wri15]. Aesthetic [Mil98a]. African [BMCM06, Osb10]. Age [Pog03, SJ00, BBB^{+}02, Osb10]. Agents [Dem98]. Ago [Oet12]. Aho [NK07]. Aleph [Ram22]. Aleph-bet [Ram22]. Algorithm [DW88, Dav98a, Dav00a, DW06, PB98, PHE17, Unixxa, The21d, The21k, Wri20, Tau16]. Algorithms [Atk01]. ALI [Mil03]. ALI-BABA [Mil03]. Aligned [Nic98b]. Alley [PB98]. Alphabet [Fir93, MVM09a, Bie22, DR68]. Alphabets [Aki80, Gid01]. Ambiguities [Cla97]. American [Whi05a]. Amsterdam [Uni00b]. Analysis [Con02, Gil98a, SD95, MVM09b]. analyst [ZV97]. Ancient [BG97, Per09]. Andries [Des08]. Angeles [IEE90]. Animation [PM17]. Annex [AH05, Dan08, DF06, DW08, JC08, Mul08, The21d, The21e, The21f, The21g, The21h]. Annexes [Dan08]. annotated [Bie23]. Annotation [Uns96]. Annotations [Anoxxa, Unixxb]. Announcing [Ano19a, Ano20a]. Annual [BJ93, BJ94, USE01]. anog [J^{+}09]. Anomalies [FMW06]. Anpassungen [Ano07a]. Answers [Ano95a]. anywhere [Ano22]. API [Bro98, Ric95]. APL [BHT93]. Apple [Ten97a]. Applet [For96]. Applets [McF97, McF98a]. Appliances [Wel97b, Wel97a]. Application [Ber97, Cob98, Ish01, Li98, Mor98, MP93, Nie96, Nie97a, Nie97b, Nie98, RM97, Sunxxa, Tan94, Ten98, DGK96, LL94a, Ric94]. Application-Independent [RM97]. Applications [AKW07, Ams97, Atk97b, Bar98, Bin97, Bro97, Cai94, Dav97a, Fow96, Gav96, Lu01, May97, McF98a, Cos03, SHA93, Sch02, Won98, Woo99, Coa93, DG95, LL98, Tay92].

Applications/Applets [McF98a]. apply [Sui98]. Apprendre [Swi09]. Approach [Cap98, Gil98b, Mud97, J^{+}12, Kim92a]. April [SHA93, SC98, Uni94a, Uni96a, Uni98a, Uni01a, Uni04a, Uni05b, SSB97]. AquaBrowser [Ano05a]. Arabetic [Abu04]. Arabic [MVM09b, BM97a, Bec87, Har95, Har97a, Har14, Mac86, MVM09a, Mil96, Mil97b, Mil97a, Mil98a, Mil98b, PHE17, Saa08, Wed96]. Arbitrary [McC72]. Architectural [Cap98]. Architecture [Ano96b, Dud97, Fre96a, Fre97d, Fre97e, Gri98a, Hiu98, LL98, MNSJ99a, MNSJ99b, ZV97]. arithmetic [MVM09b]. Armaturenhersteller [Ano06a]. Armenia [You98]. Arpanet [PB21]. art [CMD^{+}97, PCM^{+}97]. AS/400 [Shi98a]. ASCII [Kuh99a, LBD08, Pet93a, Vac91]. Asia [Har07b, Uni96a, Uni98a]. Asian [Ano94a, Fre88b, Fre00a, Hah97, PB21, Sun98b, Sunxxa]. Asian-Language [Sun98b]. Aslib [Ano95b]. Aspects [Gid01]. Assessment [BJ93]. Assisted [BJ93]. associate [AH12, Ast15]. Association [Ano96a, Pog03, SD95]. Atlanta [Uni03a]. Atlas [MAD^{+}06]. ATypI [Pog03]. Aufwand [Ano07a]. August [BM98c, ME94, Uni99a]. austauschen [Ano07b]. Authentic [Mil98a, Mil98b]. automata [NK07]. Automated [Age91, ME94]. Automation [Bar97, Cle88, SSB97]. Avoiding [DLCS22]. AVX [CL23]. AVX-512 [CL23]. away [Swo05].

BABA [Mil03]. Babel [Wag06, Lem97]. back [MR08]. Backgrounds [Mil98b]. Bad [Fen96]. Bangla [Saj19]. Barcode [Hoh97]. Barrier [Ber09, Che97, Vac91]. base [N^{+}05a, N^{+}05b, Ngô05]. base62 [Wu01]. Based [Car93, Che98, CPJ^{+}98, FDW06, Gav96, Gil97, Gil98b, Mar98, Nic97c, Nie97b, RSLO04, UY98, Bin97, Hag13, Ish01, Nic96, Nie97a, Nie98, Noj01, PWC12, ZV97]. Bases [Fly98, Swi09]. Basic [Kap01c]. be [Ano22, Fow97, Kin12]. beachten [Ano06b]. become [Swo05]. beginner [SK18]. beginners [Bul05]. Beginning [SP87]. behind [Jor02]. Benefits [CPJ^{+}98]. Bergen [Ano96a]. Berlin [Uni05b]. Berry [Pog03]. beseitigt [Ano06b]. Best [Uni04b]. bet [Ram22]. betreibt [Ano07b]. Betriebssystem [Czy98]. Between [Hon96, Hon98b]. Beyond [Mul05, PDR17, Lem97]. Bibliographic [Cob98, MfFH07, CLLL98, Cha02]. Bibliothèques [Ali10]. Bibulous — a [Hag13]. Bidirectional [Abr98, Atk01, Dav98a, Dav00a, Gid01, Unixxa]. billions [LM22]. Binary [DS01]. Binary-Ordered [DS01]. biographical [Wed96]. Bit [Bar97, Phi01, Ano06c, Fei94]. bitmap [PLYZ95]. Bits [Dav01a]. BMP [Mul05]. BOCU [DS01, SD02, SD06]. BOCU-1 [SD02, SD06]. BOF [Ano96d, Ano96e, Ano97g, Ano97d, Ano97e, Ano97f]. Book [Ano89, Pog03, Ano07]. Bookshelf [FV97]. Boost.Unicode [WG05]. Bootstring [Cos03]. Border [MJF07]. Boston [Uni99b]. bother [Fai95]. bottle [MR08]. Boundaries [Dav01d, Dav03b]. Boundary [Gil98a]. Breaking [Ber09, Che97, Fre88a, Fre98a, Fre00b, The21d, AH05]. Breaks [Vac91]. Brief [Nak98, Fre06]. Browsers [Woo99]. Browsing [Bun97]. Bruce [KBKHF16]. Budapest [SSB97]. Budget [GN98a, GN98b]. Buffers [Kue05]. Bug [CPJ^{+}98]. Bug-Free [CPJ^{+}98]. Buggy [CPJ^{+}98]. Builder [Atk98a, Atk98b]. Building [Bar98, Had01, Ten98, Uni04a]. Bulletin [Gav97]. Burlington [Ano02b]. Burmese [DAA^{+}92]. Business [Che01, Mor98, Uni03a]. byte [KL21, Van91]. Bytes [Mee97]. Byzantine [Nic06].

C [Hip94, Pla99, J^{+}09, Dav97b, Dür94, Gwy93, Nes94, Pet93b, WG05]. C# [DD11].

C/C [Pla99, Dür94]. CA [IEE90, Uni96b, USE01]. Calendar [Kum98]. CALICO [BJ93, BJ94]. California [Uni94b, Uni95a, Uni96c, Uni97b, Uni98b, Uni99a, Uni00a, USE93, Uni01b, Uni04b, Uni06, Uni02c]. CALIS [Rev93]. calligraphy [MVM09a]. Can [Lev98, Ano22]. cannot [BEE95]. Canonical [Sch02]. CAPTCHA [RM16]. caracteres [Cha07, AGH95, Ano20b, Har04]. Care [RSLO04]. Caret [Abr98]. Cascading [Lil97, Lil98a]. Case [Ano97a, Bin97, BMCM06, Dav99a, Dür98b, GN98a, GN98b, Gre98a, Gre98b, Had01, Har99, Het98, Kap01c, Keb97, KT97, McK96a, McK97c, McK97a, Saa08, Wri20, Wri22]. Case-history-to-date [McK97c]. Cataloging [Age91]. Catching [Pou02]. Categories [Whi11]. Centaur [KBKHF16]. Central [SSB97]. Centric [Ham96]. certified [AH12, Ast15]. CESU [Phi01]. CESU-8 [Phi01]. Challenge [Ho96]. Challenges [Cav97a, Mil98a, TJ97, Uni05a, Osb10, Rob10]. Changes [BM98a, BM98b]. changing [Wri20, Wri22]. Character [Alv98, Ano96c, Bar97, BB93, BB97, Che94, CY98, Coy06, Dav99b, DS05a, DW08, DW12, Dür98a, DYI^{+}02, DYI^{+}04a, DYI^{+}04b, Edb97a, Edb97b, EHJ98, Eij08, Fre88b, Fre02a, Fre02b, FMW06, Gol97a, Gol98c, Gol98b, Hah97, Hon96, Hon97, Hon98b, Ish97c, Lun96, Mar98, McC72, Mey01, Mic99, Mul08, Mur96, Par16, PBA92, Ran97b, Saa08, SSB^{+}05, Sid05, SP87, Sun96, The91, Uni95b, The06a, WD98a, WD99, Whi05b, Whi11, YY97, Yan98, Yas05, BHT93, BM98c, Cal91, Che97, Cle88, Czy98, Fei94, JBCG97, Kim92b, Kwa95b, Leg91, Lun93b, N^{+}05a, N^{+}05b, Noe94, PLYZ95, Ram22, Shi98b, Ngô05]. Characteristics [Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c]. Characters [BH98, DW02, EHJ98, ML91, Pet93b, WA98b, BP11, CL23, LM22, LBD08, PWC12]. characters — [Mil03].

CharMapML [DS05a]. Charset [FP98]. charts [Bie23]. CHASE [Bri97b, Bri97a]. chat [MVM09a]. Checking [Liu97]. China [Eme01b, Mey01]. Chinese [Tan94, Zha97a, Cha97, Che98, Eme01a, HH89, Kun94, Liu97, Lu01, Lun99, MfFH07, PZ92]. Chon [PB21]. CJK [KH98, Lem97, Sun96, Tan13, Zha97b]. CJKV [Lun98b, Lun99]. Clara [Uni94b, Ano94b]. Clash [Sas98]. Class [McK97e, McK97b, McK98b]. classical [Per09]. Classification [RSLO04, Lon00]. CLDR [The16a, The21a]. Client [Nic96, Nic97c, Yon97]. co [Ano02b]. co-operation [Ano02b]. COBOL [Swo05]. Codage [AGH95, Cha07]. codages [Har04]. Code [Ano89, BB93, DLCS22, Hal97a, Lun98b, Par16, Pol91, Ano07, Box95, Cal91, Che97, Kim92a, Kin12, Lua90, Nes94, Ram22, Van91]. Coded [Mey01, N^{+}05a, N^{+}05b, Ngô05]. Codepoint [Mik98]. codes [Kim92b, Yas05]. Coding [Hah97, Keb97, KT97, PB98]. Coincidence [Sas98]. Collaboratory [RSLO04]. Collation [DW88, Dav01b, DW06, Kap01b, Noj01, Tau16, The21k, WD98b, Wis02]. color [Pog03, RM16]. Colorful [Hag16]. Coloring [WFD08]. combination [Bee00]. Combining [McG02]. come — Finally [Bee00]. Comes [Moo97, Ano91]. Comic [Mun19]. Commands [Ric98]. comments [Bee15]. Commerce [Gre98a, Gre98b]. Commercial [Ano89, Ano07]. Committee [Wei07]. Common [Bro98, Dan08, The16a, Kim92a]. CommonPoint [DGK96]. Communicating [MH95]. Communication [BP16, Cha12, PDR17]. Communicator [TJ97, Tan97]. Community [Bri96]. Comparing [O’C98]. Comparison [Ull21]. Compatibility [DS12, Phi01, The21m]. Compatible [SD02, SD06, Hor12]. compatibles [Tre94]. compiler [BVG97, Dür94]. Complex [Bis97b, Bis97a, BBM98, Cor00, Edb98, Gol98a, Jen97c, Jen97a]. compliant [Swo05]. Component [Sug96]. Components [Sun00a, The21b]. Composition [Lu01]. Compression [DS01, Ewe04b, FB98, Fre97b, Ran97b, SD02, SD06, SC98, Wol96, Wol97, WWW^{+}97, W^{+}00, Cra98, Ewe04a]. Computer [Ano95b, BJ93, Hiu97b, Hiu98, Mac86, Mil97a, MNSJ99b, Pol91, CPJ^{+}98]. Computers [Ano96a, Dic22, ML91]. Computing [Ano96a, Car93, Ham96, HH89, Lun99, McK96b, McK97d, Tan94, Uni05b, Uns96, Lua90, Osb10, Wel01]. Conception [Lun96]. concepts [Mil03]. Conceptual [Bun00]. Conference [Ano95b, Ano02b, IEE90, IEE94, SC98, Tan94, Uni95a, Uni96a, Uni96c, Uni97a, Uni97b, Uni98a, Uni98b, Uni99b, Uni99a, Uni00b, Uni00a, Uni01a, USE93, USE01, Uni01b, Uni02b, Uni02a, Uni02c, Uni03b, Uni03a, Uni04a, Uni04b, Uni05b, Uni05a, Uni06, Ano96a, SSB97, Uni96b]. Conformance [WF04]. Congress [FCJ04]. Connection [Uni99b, Uni99a, Uni00b, Uni00a, Uni01a, Uni01b, Uni02b, Uni02a, Uni03b]. Considerations [DS05b, Shi98b]. Consortium [BJ93, Bos97b, Dan93, McG04, Uni07]. Content [Ali05, Nic96, Nic97c, Zif01, Fow97]. Content-enabled [Fow97]. Continued [Ano96b, Ano97e, Ano97f, BMN98b, Fre97e, Hal98c, Ish98e, Ksa96, Ksa97a, Ksa97b, Man98c, Pan98b, Pan98d, Pra98b, Ano96e]. Control [Ano05b, SP87]. Controversies [Ste96]. Controversy [Par16]. Convergence [Ksa96, Ksa97a, Ksa97b]. conversation [PB21]. Conversion [Ano09, C^{+}09, Edb98, Lun98b, McK96a, McK97c, Sid06, YY97, Che04a, Che04b, Ned04]. convert [LBD08]. Converter [Bot97a, Bot97b]. Converting [McK97a, Zif01]. cookbook [Woo10]. Corasick [NK07]. Core [Uni11, Uni12, The16b, The16c, The16d]. Cork [MR08, Vie08]. corners [Eev15]. Corpora [Saa08]. Corporate [Cav97a]. correctly [Ull21]. Creating [AwRMW05, Bie23, DGK96, Dav97a, Ho96, TS96]. creator [Ano03a]. Cross [Lun98b, Nas96]. Cross-locale [Lun98b]. Cross-Platform [Nas96]. CSS [Lil98b]. CSS3 [Sui01]. Cultural [Ish97a, Ish98a, Uni05b]. Culture [Pog03, BBB^{+}02]. Cutting [SJ00]. Cutting-Edge [SJ00]. CW [Ano06b]. CW-Serie [Ano06b]. cxterm [PZ92]. Cybernetics [IEE90]. Cypher [Ano89, Ano07]. Cyrillic [Whi06]. Czech [Uni03b].

D [Pog03, BAP20]. dād [Har14]. dafur [Ano07b]. Dagaare [BMCM06]. dahs [Ram22]. daijiten [KRE98]. Dakelh [Dic22]. Dangerous [Pot05]. dans [Cha07]. Dark [Eev15]. d’ASCII [Ali10, AGH95]. Data [BI95, Dav04b, Dav05b, Fly98, Gra96, Hon98a, Kin12, LL94b, Noj01, SC98, The06a, The16a, PWC12, Spr96, Wed96, Wri23]. Database [Ano09, DW08, DW12, HM06, JC08, MfFH07, MKG97, McK97e, McK97b, McK98b, Mul08, Tex98a, Tex98b, The06a, The21g, Ano08, Cha02, C^{+}09]. Databases [Buc96a, Buc96b, Buc96c, BM97c, Buc97, Buc98, Gon98, Noj01]. Date [McK96a, McK97a, Tex01, McK97c]. Day [PF97]. DB [Nor96]. DB2 [Ano09, C^{+}09]. DC [Uni02a, Uni04a]. DCC [SC98]. Decode [BP11, Bur01]. Decodeunicode [BP11]. Defined [Yan98]. Definition [YY97]. demystified [Gil03]. dependent [Pla93]. depth [Rau14]. Derived [The06a]. describes [Eml06]. Design [Bre98, Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c, Lei96c, Lei96f, Lei96d, Lil97, Pog03, Pol91, Ran97c, Sun96, BBB^{+}02, BH93, Lil98a, Mic99, Wed96]. Designed [Pog03]. Designer [Dic22, May97]. Designing [Lun96, McG96, Kim92a].

Desktop [HM04, Nie97b, Nie98]. desperate [Lev98]. Desrichard [Ali10]. Details [HT04]. Detection [FDW06]. Deterministic [Dav03a]. Develop [Ams97]. developed [Ano22]. Developer [McK96a, McK97c, McK97a, Sunxxa, TS96, Ric94, Ric95]. Developers [Cav97a, Sunxxb, Box95]. Developing [Atk97b, Bin97, Dr.03, DLLT95, Ish97a, Kan95, Kum98, Li98, May97, MBH97, McF97, McF98a, McF98b, Nie96, Nie97a, Nie97b, Nie98, Won98, Tay92]. Development [Bar97, CY98, CPJ^{+}98, SP87, Uni96c, Uni97b, AH12, Ast15, Cle88, CMD^{+}97, LL94a, PCM^{+}97, Uni96b]. Developments [Dür97c, DYN98, Ran97a]. diacritics [MVM09a, Wel01]. Dictionaries [Fre06]. Dictionary [Bre93, Ano08, Yun00]. Diego [USE93]. Difficult [Due96, Dür96]. Digital [STS03, Osb10, Sug96]. direction [Kim92b]. Directions [Ksa98a, Ksa98b]. Directory [Ano95a, Bet97b, Kri97, WKH97, Tre94]. Discussion [Ano96d, Ano96e, Ano97d, Ano97e, Ano97f]. Dispelling [Jen97b, Kid98]. Displaying [Sar98b]. Distinguished [WKH97]. Distributed [Bet97a, Hiu98, Ish01, LL98, McK96b, McK97d, RSLO04, SHA93]. Districts [MVM09b]. dits [Ram22]. dits-and-dahs [Ram22]. Diversity [Uni05b]. Do [Ber97, Vie08, Gra99, J^{+}09, Lev98]. Document [SP87, SD95, YLM^{+}05, Nic98a]. Documentation [CPJ^{+}98, SS98]. Documents [Cha97, Lil98b, Har14]. Documentum [Zif01]. Does [Ish96b]. Domain [Cos03, Top01]. Domestic [Ano89, Ano07]. Domino [Cav97b]. Don [Box95]. done [Box95]. données [Swi09]. Dossier [Cha07]. Double [Mee97]. Downlevel [WK01]. Downloading [TK01]. Draft [Ano01, DE16, Jen08, PHE17, SD06]. Driver [TK01, See13]. drop [Hag13]. drop-in [Hag13]. DTP [Nie96, Nie97a]. Dublin [Ano02b, Uni02b]. Due [Pol91]. Dynamic [Dud97, Gri98a]. Dynaweb [Nic96, Nic97c].

e-Business [Che01]. Easier [van98]. easily [Box95]. East [Fre88b, Fre00a, Hah97]. Eastern [SSB97]. EBCDIC [Uma98, UY98, Uma02]. EBCDIC-Friendly [Uma98]. écriture [AH03]. Écritures [Ali10]. Edge [SJ00]. Editing [CS01, RSLO04, Sar98b]. Editing/Translating/Viewing [RSLO04]. Editor [DM97, For96, IF97, Ritxx, Pog03, Rev93]. Editorial [AGH95, Bee15]. Efficient [Hon98a, NK07, PB98, Cri05, Ran97b]. efforts [Jen70]. eight [Uni05a]. Eighteenth [Uni01a]. Eighth [Uni96a]. einheitliche [Ano07b]. Electronic [Bre93, SP87, Zha97a, BEE95]. Eleventh [Uni97b]. elsewhere [MR16]. Embedded [GN98a, GN98b]. Emly [Eml06]. Emoji [Dav15, DE16, PDR17]. emojis [Hag16]. Emulator [PZ92]. enabled [Fow97, JBCG97, JBC98]. Enabler [Kli98a, Kli98b]. Enabling [Tan98]. Encoding [Bun00, CPJ^{+}98, Edb98, Hoh97, Mil96, Mil97b, Mud97, Phi01, Sar06, SP87, Enc93, The91, Uni95b, WD98a, WD99, Whi06, Zha97b, Mik98, Eij08, Gil03, Leg91, Lun93b, Cos03, SK18, Vie08]. Encodings [Edb97a, Eme01a, Gol98c, HH07, Har07b]. encyclopaedia [KRE98]. Endian [Gil94]. Endian-neutral [Gil94]. Engine [Cor00]. Engineering [Yer98b]. engineers [Ess98]. engines [Mit21b, Saj19, Wri23, ZV97]. English [Czy98, Box95, Che97, KRE98, Whi05a]. enhance [RM16]. Enterprise [MJF07]. Entire [Tor96]. entitled [BJ94]. Entries [Ano95a]. Environment [Ali05, Liu97, Sun98b, BM98c, Shi98b, Sug96]. Environments [McK96b, McK97d, Sun00b].

Equivalence [Dür98b, Sch02]. ère [Har04]. ERP [Ano07b, Ano07a]. ERP-KONSOLIDIERUNG [Ano07b]. ERP-System [Ano07b]. Errors [Liu97]. Ethernet [Ano95a]. Ethernet-II [Ano95a]. Ethiopian [DAA^{+}92]. Etruscan [MVM09b]. Eudora [Res98]. Europe [SHA93, SSB97, Uni97a]. European [BGI94, Bri96, Bri97b, Keb97, KT97]. evaluation [Chi17]. événements [Swi09]. every [Hip94]. everyday [Woo10]. everyone [Hor12]. Everywhere [RGN16]. Evolution [Gra96, Uni02c]. Evolves [Fow97]. Examples [HT04]. Experience [Rob98b]. Experiences [McF97, Saj19]. Experiments [Bri97b]. Explained [Kor06, MRC11]. explains [Box95]. Exploiting [MLJ13, MLJ17]. Exploits [MJF07]. Exploratory [Uni92]. Explorer [BB97, Suz98]. Exploring [Dür98c]. Exposing [WFD08]. Expression [Anoxxb, Dav99c, Chi17]. Expressions [Dav03c, Dav04c, Dav21, Dav22]. Expressive [PDR17]. Extend [AKW07]. Extended [Har95, Lun96]. Extending [BHT93, Kin96]. Extensible [Cor00, HNTT96]. Extension [Lil97, Lil98a]. Extensions [Sui01, JBCG97, Tan13]. extensively [Pog03].

factors [BJ94]. Fall [Ano06b]. Familiar [Ano89, Ano07]. family [BGI94]. FAQ [Kuh99b, LaB00]. Features [BM97b, HM01]. February [Uni02a]. Fifteenth [Uni99a]. fifth [Uni04a]. file [PLYZ95, SK18]. Files [Bre93, Pla99, FB98]. Filtering [Cha97, Kue05]. Fingerspelled [PM17]. First [Hod98, Uni02b]. Flagstaff [BJ94]. Flashcard [RC98]. Flexible [Mik98, Sar98a]. Florida [Uni05a]. fluent [Che97]. Foldings [Fre02a]. Font [Dud97, Gri98a, HM01, Lil97, Lil98a, Pet93c, SP87, TK01, Wel96, Bie22, Bie23, BH93, Kew05b, N^{+}05a, N^{+}05b, PLYZ95, Ngô05, Vie08].

Fontes [Har04]. Fonts

[Ano97b, HH07, Man97b, Man97a, Man98c, Man98a, Man98b, Woo99, Bee07]. Forces [Pol91]. Foreign [Gav97, Ano89, Ano07, Cle94, Tre94]. ForeignDesk [Nas96]. Forge [Ler03, Ler13]. Format [Ano06c, GD94b, GD97, Hos24, Uma98, Yer03, Wu01, Wu02, Yer96, Yer98a]. Formats [Cri05, Tex01, PLYZ95]. Forms [Dav98c, DD00]. Fortran [MRC11]. Forum [Ano96d, Ano96e, Ano97g, Ano97d, Ano97e, Ano97f]. Fourteenth [Uni99b]. fourth [Uni03a]. Framework [Ish01, Lin98, OL98]. Francisco [FCJ04, Uni06]. Fraud [FDW06]. Free [CPJ^{+}98, Abu04]. Frequency [CY98]. Friendly [Uma98, Abu04, Swo05]. Fully [JBGC97]. Fundamentals [DYI^{+}04a]. fusion [MLJ13, MLJ17]. Future [Jen98, Ste96, Ten97b, Agr93, Kim92b, MR16].

Gaultney [Ano03a]. GB [Mey01]. GDP [Dav04a]. General [Bec98, McG02]. generalized [MLJ13, MLJ17]. Generation [Nic96, Nic97c]. Generic [Nic97b]. Genève [AGH95]. Gengogaku [KRE98]. Gentium [Ano03a]. Georgia [Uni03a]. German [BP11, See13]. Germany [SHA93, Uni97a, Uni05b]. Geschafte [Ano07b]. get [Dic22, Top01]. GIS [Li98]. Global [AKW07, Bet97a, Bos98, Dav97a, DLLT95, May97, McK98a, PF97, SS98, SJ00, Sun98a, Tay92, Uni96c, Uni97a, Uni97b, Uni98a, Uni98b, Uni99b, Uni99a, Uni00b, Uni00a, Uni01a, Uni01b, Uni02b, Uni02a, Uni03b, Uni03a, Ano07b, Che97, Czy98, DGK96, Uni96a, Uni96b]. Globalization [Fow97, Van94]. Globalizing [Gol97b, Gol97c, PF97]. Globally [MH95]. Globalzeichensatz [Czy98]. Glossary [ZV97]. Glossary-based [ZV97]. Glue [Keb97, KT97]. Glyph [Ado98, Hud00]. Glyphes [Har04]. Glyphs [BH98, EHJ98, Har95, Lun98a]. go [Swo05]. godfather [PB21]. godine [J^{+}09]. Goes [Fre97a]. Going [SJ00, Uni96c, Uni97a, Uni97b, Uni98a, Uni98b, Uni96a, Uni96b, Hag07]. Good [Fen96]. Government [Uni04a]. Governmental [CPJ^{+}98]. Graphical [Atk98a, Atk98b, Har07b]. Graphics [MNSJ99b, dC00]. Graphite [Cor00]. Greek [Har99, Whi06]. Green [CPJ^{+}98]. Guide [BGD06, Eme01a, Sunxxa, Sunxxb, Sun00b, TS96, Ess98, Ess00, Gil03, H^{+}95, Rau14, Ric94, Ric95, SK18, Tre94]. Guidelines [Anoxxb, Dav98b, Dav99c, Dav00b]. Gurmukhi [Sid05, Sid06]. Guts [Dav01b].

hacker [Lev98]. Hamburg [SHA93]. Han [JC08, Mar98, N^{+}05b, PLYZ95, Shi98b, The21g, Whi06, N^{+}05a, Ngô05]. Han-Character [Mar98]. hand [Ano95a]. Handle [Sun98a, Ull21]. handles [Huf17]. Handling [Dür98c, Gol98a, Hag11, Jen97c, Jen97a, WWW^{+}91, Cle94, DGK96]. Hangul [Kim92a]. Hansgrohe [Ano06a]. Hardbound [Pog03]. Headers [GC96]. Hebrew [Ram22]. Held [Ano94b]. Hello [PT93, PT00]. help [Dic22]. Here [Ano91, Pet93a]. heterogeneous [CLLL98]. hiding [PWC12]. Hindi [Wag06]. History [McK96a, McK97a, MVM09a, Nak98, DR68, KBKHF16, Lun93b, McK97c, Mun19, Pik03, Ritxx, Wed96]. Hitchhiker [Eme01a]. HK [Mey01]. Holland [Uni00b]. home [Uni07]. Homepage [Ho96]. Homograph [WFD08]. Hong [Uni01a, Che01, Uni96a]. hopes [Dic22]. Hotel [Ano02b, IEE90]. HP [Kli98a, Kli98b, Tiv97]. HP-UX [Kli98a, Kli98b]. HTML [DYN98, Woo99]. Huffman [PB98]. Human [BJ94]. Humanities [Ano96a, Uns96, Uns96]. Hypertext [Ben96, YNAD97]. Hyphenation [MR16]. hytung [Ano08].

I18N [Bec98]. IANA [FP98]. IAPR [IEE94]. IBM [Atk97a, Atk98a, Atk98b, Shi98a, Tre94]. ICCC94 [Tan94]. ICPC [RSLO04]. ICU [The21b]. Identification [Con02]. Identifier [Dav05a, The21f]. Identifiers [Due97a, Dür97a, Dür98b, Rob15, Wu01]. Identity [FDW06]. Ideograph [Dür98c]. Ideographic [HM06, Lu01]. Ideographs [Dür98c, Jen98, Jen08, Jen12, The21h]. IDN [Pot05]. IDNA [Cos03, DS12, The21m]. IEC [N^{+}05a, Ali94, Kim92b, Ksa97a, LaB00, N^{+}05b, Ngô05, Zha97b]. IETF [Alv98]. IFLA [ME94]. II [Ano95a, Hal94b]. IIIMF [Hiu98]. illustrated [Pog03]. im [Ano06b, Czy98]. Image [Uns96]. IME [Sar98a]. Impact [Car93, Che94, Ish97b, Ish98c]. Impacts [Ish96a, Ish97d, Ish98d, Ish98e]. Implementation [Dav98a, Due96, DNYA96a, DYN97a, Dür96, DNYA96b, DYN97c, DYN97b, Kun94, Lun96, Mey01, Ran97a, Ran97c, Sar98a, Sun96, Tiv97, Unixxa, NK07, Tau16]. Implementations [Abr98, Atk01, Dür98d, Hah97]. implementiert [Ano07b]. Implementing [Fow96, Lei96a, Lei96e, Lei96b, Lun98b, Tan96]. Implementor [Ano94b, Ada94]. Implementors [Uni94a, Uni94b, UIW92]. Implications [Bos98, Kwa95a, Van91]. Importance [Che01]. important [Box95]. improve [Cra98]. Improving [Age91, MR08]. in-depth [Rau14]. incompatible [PLYZ95]. incomplete [Ritxx]. Increased [Ano94a]. Independent [Hiu97b, Hiu97a, RM97, Wel96]. India [MNSJ99b]. Indian [Cha02, Mur98]. Indic [Ish03, MNSJ99a, MNSJ99b, Wis02]. indigenous [Osb10]. Industry [Nak98, Ano02b]. Informatics [FCJ04]. Information [Cle94, DDG96, Eme01b, Gav96, Ho97, Ish97a, Ish98a, Lun93a, Lun99, McK98a, Wel97b, Wel97a, You98, BEE95, Cal91, Noe94, Wed96, Yun00]. Informationen [Ano07b]. Infrastructure [Che01, Uni04a]. Inland [Ano89, Ano07]. Input [Gav96, Hiu97b, Hiu97a, Hiu98, Lei96c, Lei96f, Lei96d, Leo98, Lin98, Mar98, MKG97, Mat96, OL98, Mil03, Shi98b]. Insider [H^{+}95]. insights [Saj19]. Institute [Uns96]. Instruction [BJ93, KL21]. instructions [CL23, LM22, MLJ13, MLJ17]. Integrating [Mah09]. Integration [CPJ^{+}98, CMD^{+}97, PCM^{+}97]. Intentions [WFD08]. Interchange [SP87, Spr96]. Interface [Bec98, Ric94]. Interfaces [dN96]. Interlinear [Anoxxa, Unixxb]. International [Ano02b, Ber97, Che01, DG95, Dom96, Dr.03, IEE90, IEE94, Kan95, Kum98, Lil97, Pog03, RSLO04, SSB97, SD95, Sun00a, Sun00b, Tan94, TJ97, TS96, Uni95a, Uni96a, Uni96b, Uni96c, Uni97a, Uni97b, Uni98a, Uni98b, Uni99b, Uni99a, Uni00b, Uni00a, Uni01a, Uni01b, Uni02b, Uni02a, Uni02c, The21b, dN96, Ano96a, BBB^{+}02, Cal91, Kim92b, Tay92, YLM^{+}05, Lil98a]. Internationale [Pog03]. Internationalisation [Ish97c]. Internationalisierung [Ano06a, Ano06a]. Internationalization [Ano96b, Ano97g, Ben97, Bet97b, Bos97a, Bos97b, Cap98, DC01, DLLT95, Fre96a, Fre97d, Fre97e, Gri98b, Gri98c, Hal94a, Hal94b, Hal98b, Hal98c, Hiu96, Hiu97c, Ish01, Kri97, McC96, Pan98a, Pan98b, Ric98, Sun98a, Sunxxb, Suz98, Tan97, Uni03b, Uni03a, Uni04a, Uni04b, Uni05b, Uni05a, Uni06, UHP93, YNAD97]. Internationalized [MBH97, McF97, McF98a, McF98b, Cos03]. Internationalizing [Due97a, Dür97a, Res98, Coa93]. Internet [Uni96b, Uni98b, BB97, Due97a, Dür97a, Dür97c, Edb97b, Fow96, Gol97a, Gol98b, Gre98a, Gre98b, Hiu97b, Hiu97a, Hiu98, Mur96, Oud97, PB21, Pan98a, Pan98b, Suz98, Uni96a, Uni96c, Uni97a, Uni97b, Uni98a]. Interoperability [GC96, Kli98a, Kli98b]. Interview [Ano03a]. Interweaving [RM16].

Introduction [Ada94, Fre98b, Hal97a, HH89, Ish03, Ree17, UHP93]. Involve [Ish96b]. IO [Eck97]. ‘i’rab [MVM09a]. Ireland [Ano02b, Uni02b]. irilic [J^{+}09]. ISBN [Pog03]. ISCII [Sid06]. Islamic [MVM09a]. ISO [N^{+}05a, Ada94, Ali94, Bri96, Bri97a, Con02, Kim92b, Ksa96, Ksa97a, Ksa97b, Ksa98a, Ksa98b, LaB00, N^{+}05b, Yer03, Tex01, Ngô05, Uni96a, Wu01, Wu02, Yer96, Yer98a, Zha97b]. ISO/IEC [N^{+}05a, Ali94, Kim92b, Ksa97a, LaB00, N^{+}05b, Ngô05, Zha97b]. ISO10646 [Uni98a]. Israel [IEE94]. issue [Pro15, Pro16]. Issues [Coa93, Eme01b, Hon96, Hon97, Hon98b, Ish97a, Ish98a, Mur98, Sid06, Wis02, BM98c]. Istanbul [BM98c]. Itaiji [MKG97]. Items [Age91]. IUC12 [Uni98a]. IUC16 [Uni00b]. IUC17 [Uni00a]. IUC18 [Uni01a]. IUC19 [Uni01b]. IUC20 [Uni02a]. IUC21 [Uni02b]. IUC22 [Uni02c]. IUC23 [Uni03b]. IUC24 [Uni03a]. IUC25 [Uni04a]. IUC26 [Uni04b]. IUC27 [Uni05b, Uni06]. IUC28 [Uni05a].

January [USE93, Uni02a, Pro16]. Japan [Uni94a, Uni98a, Pan98e, Yas05]. Japan1 [Yas05]. Japanese [ACE^{+}08, Bre93, CY98, Har07a, HH89, Kun94, Lun93b, Lun93a, Lun96, Lun99, Nak98]. Java [Ano96b, Bec97, Bec98, Ber97, BVG97, Bin97, Cav97b, Che97, Cob98, Dav97b, DD03, DD07, DC01, Eck97, Fre96a, Fre97d, Fre97e, GJS96, G^{+}xx, GJSB00, Ham96, Hiu97b, Hiu97a, Hiu98, Hor12, Ish01, J^{+}12, Kum98, Leo98, Lin98, McF97, McF98a, McF98b, O’C98, OL98, Ran97a, Ran97c, RC98, Ste96, Won98]. Java-based [Bin97]. javar [BVG97]. JavaScript [Rau14]. JDK [Dav97a, Gil98a, MBH97]. jedoch [Ano07b]. Jerusalem [IEE94]. JIS [Sui98]. Jitai [Sas98]. John [Pog03]. Join [Pol91]. Joint [Ano96a]. Jose [Uni95a, Uni96b, Uni96c, Uni97b, Uni98b, Uni99a, Uni00a, Uni01b, Uni02c, Uni04b].

Joseph [Wri15]. Journée [AGH95]. Julia [DL19]. June [Ano96a]. Junicode [Bie22]. Just [PDR17, Ish97c].

Kangxi [Ano08]. Kanji [RC98, CY98, MKG97, Nak98, RC98]. Keeping [Jor02]. key [DR68]. Keyboard [BMCM06, Dic22, Kim92a, See13]. Keynote [Bos98, Bre98, Nak98, Sch97]. Keys [Noj01]. Kharosthi [MVM09b]. Khmer [DAA^{+}92]. Kilnam [PB21]. klassebibliotek [WG05]. Klingon [Box95]. Know [Har97b, Hip94]. knowledge [N^{+}05a, N^{+}05b, Ngô05]. knowledge-base [N^{+}05a, N^{+}05b, Ngô05]. Known [FMW06]. kog [J^{+}09]. kompatible [Ano07b]. Kong [Uni96a, Uni01a, Che01]. konnen [Ano07b]. KONSOLIDIERUNG [Ano07b]. Korean [HH89, Jun97, Kun94, Lun99, Par16].

lacks [Che97]. LAN [Ano95a]. land [Wri22]. Langages [Cha07]. Language [Atk96, Atk97b, Atk97a, BBB^{+}02, Ber09, Buc96a, Buc96b, Buc96c, BM97c, BM97b, Buc97, Che97, Cle88, Con02, Dav04a, Dav04b, DS05a, GJS96, G^{+}xx, GJSB00, Nic97a, Nic97b, Pla93, PM17, WA99, Sun98b, Sun00b, Tan98, WA98b, WA98a, Wis02, YNAD97, Yer98b, Box95, BAP20, Dic22, JBCG97, Lon00, Osb10, Ste96, Tre94, Dav05b, Pog03]. Language-dependent [Pla93]. Languages [Alv98, BB97, BMCM06, BG97, CPJ^{+}98, DDHF99, DF02, Gid01, Hal94b, Jor02, Mur96, Mur98, Box95, Cle94, MAD^{+}06, Osb10, Wu00]. Langues [Cha07]. Large [Cha97, Het98, Zha97a]. Large-Scale [Zha97a]. Largest [Mor98]. LaserHebrew [Ano02a]. Latest [Tan94]. Latin [IF97, Har14, Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c, Whi06]. Latin-X [IF97]. Latin/Arabic [Har14]. Law [CPJ^{+}98]. layer [LL94b]. Layout [Abr98, ACE^{+}08, Bec98, BBM98, Ete05, Ish13]. LDAP [Bet97b, Kri97]. LDML [Dav04b, Dav05b]. Leaks [Eve02]. Learning [BJ93]. learns [Kew05a]. left [Jor02]. legacy [Che97, Har07b]. Lending [Ano95a]. less [KL21]. Letter [IF97, Rob15]. Letters [CPJ^{+}98, DM97, Fir93]. Level [BM97a]. Lexical [LBD08]. Libraries [Bri97b, Noe94, Sug96]. Library [Bar97, Bri96, Coy06, Gav97, SSB97, Wei07, ME94, WG05]. Lifecycle [Cap98]. Ligature [Mil96, Mil97b]. ligatures [Pla93]. light [N^{+}05a, N^{+}05b, Ngô05]. Lightweight [WKH97]. like [Bie23]. Line [AH05, Fre88a, Fre00b, The21d, Zif01]. Lines [Fre98a]. lingual [Rev93, Shi98b]. Linguistic [Ano96a]. linguistics [KRE98]. Linking [Wed96]. Linus [CPJ^{+}98]. Linux [USE01, Kuh99b, Noj01]. Lion [Kew05a, Kew05b]. LISA [Ann97]. Lisbon [Wei07]. Lisp [Ste96]. List [CY98, Whi05a]. Literary [Ano96a]. Literature [Gav97]. Locale [Atk98a, Atk98b, Jun97, Noj01, O’C98, Sun98c, The16a, Lun98b, Dav04b, Dav05b]. Locales [Ano97a, Hal94b, Noj01]. Localisation [Ish96b, Ano02b]. Localization [Cap98, Cav97a, Mur98, Uni02b, UHP93, DGK96, Ess98, Ess00, Fow97]. Localizing [Nas96]. logical [Har07b]. London [Ano95b]. look [Fei94]. Lotus [Cav97b]. LRC [Ano02b]. LSCEP [Zha97a]. lualatex [Bee20]. LZ [Cra98].

Mac [See13, McC96, See13, Ten98]. Macintosh [Ten97b]. Madrid [ME94]. Mail [Bot97a, Bot97b, Dem98, GD94b, GD97]. Mail-Safe [GD94b, GD97]. Maintaining [MJF07]. Mainz [Uni97a]. Making [Hos24, Ish98b, Nic97a, Yon97, YLM^{+}05]. Man [IEE90, Jen70, PB21]. Management [Gra96, Het98, Zif01]. managers [Ess98]. Manifesto [RGN16]. mankind [DR68]. mannered [Dav97b]. Manual [N^{+}05a, N^{+}05b, Ngô05]. Map [Ten97a]. Mapping [Dav99b, DS05a, Hon96, Hon97, Hon98b]. Mappings [Dav99a]. MARC [MfFH07, Ali05]. March [BJ94, SC98, Uni99b, Uni00b, Uni03b, Uni04a, Uni06, Wei07, BJ93, Uni97a]. Mark [PHE17, Top01]. market [Tay92]. Markets [DLLT95]. Marks [Kuh99a, McG02, WWW^{+}91]. Markup [Dav04b, DS05a, Dav05b, DDHF99, DF02, Nic97a, Nic97b, YNAD97, YLM^{+}05]. mars [AGH95]. Mastering [MH93]. matching [NK07]. material [BEE95]. materials [ME94]. math [Bee00, Hag11, Mah09, Vie08]. Mathematics [BFS02, BP16, Sar06, Rob10, Rob15]. matters [Fow97]. Maxim [Pog03]. May [Ano02b, Uni02b, Fow97]. me [J^{+}09]. mean [Ano22]. means [Nic98a]. Mechanism [HNTT96, Mik98]. Mechanisms [DS05c, DS05d, DE16, The21l]. Medical [FCJ04]. medieval [Per09]. Medinfo [FCJ04]. Meeting [ME94, SHA93, Wei07]. Meets [Ano97c, Rub96, Kew05b]. Megafonts [Had01]. Memory [Bun97, PB98]. Memory-Efficient [PB98]. Merrimack [UIW92]. Method [Hiu97b, Hiu97a, Hiu98, Lei96c, Lei96f, Lei96d, Lin98, Mar98, MKG97, McG02, OL98, PWC12]. Methods [Hon96, Hon97, Hon98b, Leo98]. Mexico [DO00]. Michael [Eml06]. Microsoft [Ber98, Bis97b, Bis97a, Fre97a, Kap01a, Kap01b, Li98, Pra97b, Pra97a, Pra98a, Pra98b, Pra98c]. MIGRATE [Ano95a]. Migrating [DW02, Nes94]. Migration [Ano09, C^{+}09, JBGC97, Tex98a, Tex98b]. migrierte [Ano06a]. MIME [GD94a, SD02, SD06]. MIME-Compatible [SD02, SD06]. Misunderstood [DM97]. mixed [Har14]. MkIV [Hag07]. MlBibTEX [Huf17]. Model [CPJ^{+}98, Con02, Dür98a, DYI^{+}02, DYI^{+}04a, DYI^{+}04b, Fre02b, Mat96, Sar96, Sar97, WD98a, WD99, WF04, Spr96, Nic98a]. Model-Based [CPJ^{+}98]. Modeling [PDR17]. Modern [MRC11]. monde [AH03]. Mongolian [BDA^{+}92]. Monitor [Tiv97]. Move [Pet93a]. Movement [Abr98]. Movie [Fen96]. Mozilla [Tan98]. MT [ZV97]. MULE [Mik98]. MULE-Universal [Mik98]. Multi [BM97a, BM98c, Cha02, Rev93, Rob15, Shi98b, Van91]. multi-byte [Van91]. multi-character [BM98c]. multi-letter [Rob15]. Multi-Level [BM97a]. multi-lingual [Rev93, Shi98b]. Multi-script [BM98c, Cha02]. Multicode [Mud97]. Multilanguage [BBM98, Mat96]. Multilingual [Ano96d, Ano96e, Ano97d, Ano97e, Ano97f, Bar98, Bec84, Ben96, CS01, Dem98, DNYA96a, DYN97a, DNYA96b, DYN97c, DYN97b, Dür98d, Gav96, Gav97, Gil97, Gre98a, Gre98b, HNTT96, Har97b, Ho97, JBC98, Kew05a, Kew05b, Lem97, Lil98b, May97, Mud97, Nas96, Nic98b, RSLO04, RM97, SJ00, Tan98, Uni04a, Uni05b, Uni05a, Wan97, Won98, Woo99, You98, Bee19, Bee20, BM98c, Cal91, CLLL98, ES97, Kew06, LL94a, LL98, Lua90, ME94, Top01, Tre94, Wel01, Wu00, Wu01, ZV97]. Multilingualism [Ben97, Oud97]. Multinational [Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, McK96a, McK97c, McK97a]. Multiple [Ano96c, BB97, Edb97b, Gol97a, Gol98b, Mur96, Noe94]. Multiscript [Ano97a, Jun97, La 93, ME94]. multithreading [Swi09]. Musical [Nic06]. muss [Ano07b]. Mutamathil [Abu04]. Myanmar [HT04]. Mysap [Ano06a]. Myths [Jen97b, Kid98].

Nôm [N^{+}05a, Ngô05]. Name

[Sun98a, Top01]. Named [Whi05b]. Names [Ado98, Dav01c, DF05, DF06, FMW06, Cos03, Sid05, WKH97, Whi05a]. National [Buc96a, Buc96b, Buc96c, Buc97, Par16, Wei07, N^{+}05a, N^{+}05b, Ngô05, Wu00]. Nattilingmiutut [Ano22]. Natural [Nic97a, Lon00]. nauc [J^{+}09]. Navigation [van98]. Navigator [Tan96]. NC [Hiu97b, Hiu97a]. Nearly [Sar06]. Need [Har97b, Vie08]. Net [Ham96, Gav97]. Net-Centric [Ham96]. Netscape [Dud97, Gri98a, Gri98b, Gri98c, Tan96, TJ97, Tan97]. NetWeaver [AH12, Ast15]. Network [Het98, Hiu97b, Hiu98]. Networked [Uni04b]. networks [PB21]. neutral [Gil94]. Newline [Dav98b, Dav00b]. Newmath [Küs05]. News [Saa08, Pro15, Pro16]. Next [Bec97, Bec98, McG96, OL98, Wis01]. NH [UIW92]. nicht [Ano07a]. Nineteenth [Uni01b]. Ninth [Uni96b, Uni96c, Uni06]. njegova [J^{+}09]. No [Age91, CPJ^{+}98, CPJ^{+}98]. noblest [KBKHF16]. nog [J^{+}09]. Nom [N^{+}05b]. Non [Eml06, Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c, WWW^{+}91]. Non-Latin [Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c]. Non-Roman [Eml06]. Non-Spacing [WWW^{+}91]. Nonromanization [Age91]. Nonsense [Mil96, Mil97b]. Normalization [Dav98c, DD00, Dür98b, DYI^{+}04b]. Norway [Ano96a]. Notation [Nic06, MVM09b]. Note [The21a]. Notes [Fir93, Ber15, Ber23, Cav97b, May97]. Noticing [Oet12]. November [Ano95b, IEE90, USE01]. NSCII [N^{+}05a, Ngô05, N^{+}05b]. NT [BBM98, Ric95, Agr93, Bis97b, Bis97a, Hal94a, Hal94b, Hal98a, Kan95, MP93, MH93, Nie96, Nie97b, Nie98, Pet93c, Ric94]. NT/95 [Nie96]. Number [Gil98b, PM17]. Numbers [Kap05]. numeral [MVM09b]. Numerals [MVM09b, MVM09a]. numérique [Har04]. Nunavut [Ano22]. Nutshell [HM04]. o [PT00]. Oakland [USE01]. Obfuscation [WFD08]. Object [Ish01, Nic98a, Sar96, Sar97, Swo05]. Object-based [Ish01]. objects [Dav97b]. objet [Swi09]. Occurrences [Saa08]. October [IEE94, Pro15]. od [J^{+}09]. ODBC [Nor96]. odrz [J^{+}09]. off [Jor02]. Office2000 [Pra98a, Pra98b]. Office97 [Pra97b, Pra97a, Pra98c]. oktobra [J^{+}09]. old [Pik12]. OLE [Nor96]. Omega [AGH95, Fai95, Pla94]. On-Screen [Ish97a, Ish98a]. One [PF97, KL21, Gri98b, Gri98c]. ones [Ram22]. online [BM98c]. Only [Hod98]. Open [Ano96d, Ano96e, Ano97g, Ano97d, Ano97e, Ano97f, CPJ^{+}98, Noj01]. OpenType [Hag16, Hos24, Mah09]. Operating [Fen96, Sun98b, Czy98, Wu00]. operation [Ano02b]. operations [Shi98b]. Opinion [Age91]. opportunities [Osb10]. Optical [SSB^{+}05]. Optimising [Chi17]. Optimization [Ano09, C^{+}09]. Optimized [Abr98]. Options [Ano05a, ES97]. Oracle8 [Won98]. Ordered [DS01]. Ordering [PHE17, La 93]. Ordinary [Ano89, Ano07]. oriented [Swo05]. Origins [BAP20]. Orlando [Uni05a]. Orthographic [Saa08, Wel01]. Orthographies [AwRMW05]. OS/2 [Atk96]. OSI [LL94b, LL98]. Other [Age91, Hon96, Hon98b, Woo99, Box95, DDHF99, DF02, Fre06]. output [Mil03, Shi98b]. Overcoming [Cla97]. Overdog [CPJ^{+}98]. Overview [Ali94].

P.O. [Res98]. pace [Che97]. package [Lem97, Mit21a]. packages [Uns96]. page [BGI94, BI95, Uni07, Wu02]. page-shift [Wu02]. Pages [Gil97, Hal97a, Pog03, Van91]. Pan [Eme01b]. Pan-China [Eme01b]. Paperback [DLLT95]. Papers [Age91, Ano95b]. Parallel [Ben96, Nic98b]. Parkosz [Bie22]. Part [Bec98, Bre93, CMD^{+}97, Coy06, Hal94a, Hal94b, OL98, PCM^{+}97, Gil94]. Passeport [Des05]. Past [Bun00, MR16]. Patient [Tiv97]. Patrick [Des08]. Pattern [Dav05a, IEE94, SD95, The21f, NK07]. PC [Agr93, Tre94]. PCs [Tre94]. pdflatex [Bee19]. Pen [Car93]. Pen-Based [Car93]. Perl [Lev98, Oet12]. Permanent [Wei07]. Persistent [Sun98a]. Personalized [Cha97]. persons [BEE95]. Perspective [Pan98c, Pan98d, Cha02]. Petit [Cha07]. PHIGS [Coa93]. phones [Dic22]. Phonetic [Fre06]. Phrase [Ano89, Ano07]. Phrase-Book [Ano89, Ano07]. Phrases [Ano89, Ano07]. Picking [CPJ^{+}98]. PicoVue [Tiv97]. Pictograms [Aki80]. Pipeline [Eve02]. pisma [J^{+}09]. Plain [WA99, Sar06, WA98a]. Plain-Text [Sar06]. Plan [Fen96]. Plane [WA98b]. Planning [CPJ^{+}98]. Platform [Hiu97b, Hiu97a, Ho97, McG98, Nas96, Wel96, Agr93, LL94a]. Platforms [Lei96a, Lei96e, Lei96b, O’C98]. Policies [McG04]. Policy [Alv98]. poor [PB21]. Porting [Box95, Dav97b]. Portugal [Wei07]. Positional [MVM09b]. PostScript [TK01]. Potentials [Dür98c]. pour [Des05]. Power [Swo05]. Powering [Uni03a]. Practical [Bun00, Cai94, Ess98, Ess00, Gil03]. Practices [Uni04b]. Prague [Uni03b]. pratique [Des08]. Pre [Uni96b, Cra98]. Pre-conference [Uni96b]. pre-processors [Cra98]. precis [Cha07]. Preface [Hof96]. Preference [Ish01]. Preferences [Nic96, Nic97c]. preparing [Dür94]. prerequisite [BEE95]. presentation [ES97, LL94b]. Preventing [Mit21b]. previews [Chi17]. Primary [RSLO04]. primer [Gra00]. primitives [Wri20]. principles [Har07b]. printed [BEE95, SSB^{+}05]. Printer [TK01]. Printing [Bet97a, Nak98]. Problem [Hol03, J^{+}12, Mac86]. problem-solution [J^{+}12]. Problems [Bun00, Edb98, Gol98a, Jen97c, Jen97a, Fow97]. Procedures [FP98, McG04]. Proceedings [FCJ04, IEE94, SHA93, Tan94, Uni94a, Uni94b, UIW92, USE93, USE01, BM98c, IEE90, ME94, Uni95a, Uni96b, Uni96a]. process [Box95]. Processing [Bec84, DS12, HNTT96, Har97b, Hud00, Lu01, Lun93a, Lun99, Ran97a, Ran97c, Sch04, The21m, You98, Bec87, Cal91, GV92]. processors [Cra98]. Product [Cap98, Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c, Zif01]. Production [Ber15, Ber23]. Products [GN98a, GN98b, Ish98b, Kap01a, Tre94]. PRODUKTE [Ano06b]. Profession [Hol03]. program [DD03, DD07, DD11]. programmation [Swi09]. Programmer [FV97, Ree17, Gil03, Hip94, Swi09]. programmers [Rau14]. Programming [Gar12, Hal94a, Hel92, O’D94, Van91, BAP20, Hel93, H^{+}95, MP93, MH93, Ric94, Ste96, Woo10, Wu00]. Progress [Fei94, Pla94, Zha97b]. Project [Bre93, FV97, Noj01, Pla94, Tan98, Ess98, Bee07, Bri97b, Bri97a, CMD^{+}97, PCM^{+}97]. Promises [Due97b, Dür97b]. Properties [Due97b, Dür97b, Fre88a, Fre00b, AH05]. Property [Fre88b, Fre02b, Hon98a, The06a]. proposal [Lua90]. Proposals [Uni92, Wed96]. Proposed [Ano01, BM98a, BM98b, DF06, DW06, DE16, PHE17, SD06]. Prospects [Age91]. Protocol [Hiu97b, Hiu97a, WKH97]. prototype [BVG97]. provides [Mac86]. Providing [Ano03a]. provision [JBCG97]. pTEX [Tan13]. Public [McG04]. Publications [Zha97a]. Publish [Ho97]. Publisher [Fre97a]. Publishing [Bos98, Nie97b, Nie98]. Punycode [Cos03]. Purpose [Dür98a]. Putting [MR08, Nic98b]. Python [Swi09, Tau16]. pyuca [Tau16].

Q [Ber97]. QED [Ritxx]. Qt [TK01].

Qt/X11 [TK01]. Quest [GC96]. Questions [CPJ^{+}98]. Quick [HM04]. Quotation [Kuh99a].

R [Ano06a, Ano07a, Uni98b]. R/3 [Ano06a, Ano07a]. race [Top01]. radova [J^{+}09]. Rare [Dür98c]. RCSU [Wol96, Wol97]. Reach [AKW07]. reaches [Wri22]. read [Ano22]. readers [BGI94, BI95]. Ready [Bec97, Bec98, Bro97, OL98]. Real [Ull21, PB21]. Real-world [Ull21]. really [Fow97]. Receipts [Gav97]. recipes [J^{+}12, Woo10]. Recognition [IEE94, SSB^{+}05, SD95]. Recommendations [AwRMW05]. reconstruct [Wed96]. Records [Ali05]. Reference [Dav98a, HM04, Spr96, Unixxa, Kim92b]. References [Dan08]. REGAP [FDW06]. registracija [J^{+}09]. Registration [FP98]. Regular [Anoxxb, Dav99c, Dav03c, Dav04c, Dav21, Dav22, Chi17]. Relational [Buc96a, Buc96b, Buc96c, BM97c, Buc97, Tex98a, Tex98b]. Release [The21a, Wis01]. Render [Yon97]. Rendering [BM97a, Cor00, Gol98a, Har97b, Jen97c, Jen97a, Lil98b, McG02]. Reordering [Atk01]. replacement [Hag13]. Report [Ben97, Fre97a, Jen08]. Repository [The16a]. Representation [RM97, Tex01, WKH97, LL94b]. Representing [Bie22, HT04, Ram22]. Republic [Uni03b, You98]. Requirements [ACE^{+}08]. Research [DO00]. réseau [Swi09]. Resources [Hal94b]. Rest [Hos24]. restructuring [BVG97]. Retrieval [Che98, DO00, Eme01b, Wan97]. Retrieve [Ho97]. retrospective [Hag11]. Reuters [Wol96, Wol97]. Review [Pog03]. revision [Ano08]. revitalize [Dic22]. Revolution [Moo97, Uni02c]. RFC [Alv98, FP98, GD94b, GD94a, GD97, McC72, WA99, Cos03, Yer03, McG04, Cri05, WKH97, WA98a, Yer96, YNAD97, Yer98a].

Rhapsody [McG98]. RichEdit [CS01]. Right [Ams97]. Rivals [Pol91]. Road [Ten97a]. Robust [Ete05]. Rogers [KBKHF16]. Role [Bar97, Lon00]. Roman [Eml06, MVM09b, KBKHF16]. Romanization [MVM09a]. Rule [Gil98b]. Rule-Based [Gil98b]. Russian [Gav97].

sa [J^{+}09]. Safe [GD94b, GD97]. Sample [Whi05a]. San [FCJ04, Uni95a, Uni96b, Uni96c, Uni97b, Uni98b, Uni99a, Uni00a, USE93, Uni01b, Uni02c, Uni04b, Uni06]. Sanseido [KRE98, Yun00]. Santa [Uni94b, Ano94b]. SAP [Ano06b, Ano09, Ano07a, AH12, Ast15, C^{+}09]. SAP-Tuning [Ano06b]. Satellite [ME94]. Scale [Zha97a]. Scheme [Fre97b, Lu01, Phi01, Wol96, Wol97, WWW^{+}97, W^{+}00]. Schriftzeichen [BP11]. SCHWERPUNKT [Ano07b]. Scientific [GV92]. Screen [Ish97a, Ish98a]. Script [BM97a, Bis97b, Bis97a, Dav01c, DF05, DF06, Eve02, HM01, Har99, BM98c, Cha02, Jen70, Saj19]. Script-specific [HM01]. Scripts [BBM98, Bro98, Bun00, Due96, Dür96, Eml06, Gid01, Har07a, Ish03, Man01, Cle94, Har07b, Mac86]. SCS [Mey01]. Search [Cla97, Mar98, Wan97, Shi98b]. Searching [Ano96c, CLLL98, Cha97, Hod98]. Second [ME94, UIW92, LM22, Uni02c]. Security [DS05b, DS05c, DS05d, MJF07, The21l, RM16]. Segmentation [The21e]. sein [Ano07b, Ano07b]. seine [Ano07b]. Semantic [PDR17]. Semantics [BN98]. Semicode [See13]. Sense [Mil96, Mil97b]. September [Ano94c, FCJ04, Uni94b, Uni95a, Uni96b, Uni96c, Uni97b, Uni98b, Uni99a, Uni00a, Uni01b, Uni02c, Uni03a, Uni04b, Uni05a]. Sequences [Whi05b, Ull21]. Serie [Ano06b]. Server [Ber98, JBGC97, JBC98, Yon97, JBCG97, Kap01b]. Service [Sun98a]. Services [Bet97b, Kri97, Tor96, Noe94]. Set [Che94, Coy06, Har95, Lun96, Sun96, YY97, Yan98, BHT93, Che97, Cle88, Czy98, Fei94, Kwa95b, Lun93b, N^{+}05a, N^{+}05b, Ngô05, Top01]. Sets [Alv98, Ano96c, Bar97, BB97, Edb97b, EHJ98, Gol97a, Gol98b, Hon96, Hon98b, Ish97c, McC72, Mur96, Noe94, PBA92]. Seventeenth [Uni00a]. Seventh [Uni95a, Uni05b]. SGML [Uns96]. shape [Ano02b]. shaping [MNSJ99a, MNSJ99b]. SHARE [SHA93]. sharing [PLYZ95]. Sheets [Lil97, Lil98a]. Sheraton [IEE90]. Shift [Ham96, Wu02]. should [Hip94]. Showcase [USE01]. Shrink [Pra96a, Pra96b]. Shrink-Wrapping [Pra96a, Pra96b]. Sign [Jen70, PM17]. signatures [Pog03]. significantly [Cra98]. Signs [PM17]. SILE [Coz17]. SIMD [LM22]. similar [Mac86]. Similarity [PDR17]. Simple [Cra98, Due96, Dür96, Nic97b, Har14]. Single [Rob15]. Single- [Rob15]. Sinhala [BDA^{+}92, PM17]. Sixteenth [Uni00b]. sixth [Uni04b, Ano94b]. skupa [J^{+}09]. Slow [CPJ^{+}98]. smart [Kew05b]. Snowbird [SC98]. Society [CPJ^{+}98]. Software [CPJ^{+}98, DW02, Dom96, Dr.03, DLLT95, FV97, Fow97, Gol97b, Gol97c, Hah97, Kan95, MBH97, McF98b, McK96a, McK97c, McK97a, Mor98, Mur98, Nas96, PF97, Tay92, TS96, Uni96b, Uni96c, Uni97a, Uni97b, Uni98a, Uni98b, Uns96, UHP93, Van94, Zif01, DG95, DGK96, Ess98, Gil94, Uni96a]. softwares [CMD^{+}97, PCM^{+}97]. Solaris [Ano97a, Jun97, Sun98b, Sunxxb, Sun00b, Sun98c, TS96]. sollte [Ano07a]. Solution [Eme01b, J^{+}12, Mac86, PBA92]. Solutions [Bun00, Edb98, Uni05a, Woo10]. Some [Fir93, Keb97, KT97, Pog03]. soon [Kin12]. sorgfaltig [Ano07b]. sorgt [Ano07b]. Sorting [Dav03a]. Source [CPJ^{+}98, DLCS22, Noj01, Box95, Jen08, Jen12, The21h]. sources [CLLL98]. South [Har07b]. space [PWC12]. Spacing [WWW^{+}91]. speak [Kew05a]. Speaking [Rau14]. speaks [Wag06]. Special [Ano22, Kim92b]. specific [HM01]. Specification [GJS96, G^{+}xx, GJSB00, The16b, The16c, The16d, JBCG97, Uni11, Uni12]. Spellout [Gil98b]. Spoofing [DLCS22]. Sprachbarrieren [Ano06b, Ano07b]. Sprache [Ano07b]. Spring [SHA93]. SQL [Kap01b, Ber98, BM97c, BM97b, Buc97, BM98a, BM98b]. SQL2 [Buc96a, Buc96b, Buc96c]. SQL2/3 [Buc96a, Buc96b, Buc96c]. Stability [McG04]. Standard [AH05, BM98a, BM98b, Dan08, DW06, DW08, Fre96b, Fre97b, Fre97c, Fre97f, HK98, Hos24, JC08, JBCG97, Moo98a, Mul08, Pol91, SD06, SP87, Uni91, The91, Uni95b, Uni11, Uni12, The16b, The16c, The16d, Uni20, The21d, The21e, The21f, The21g, The21h, The21k, The21l, The21m, WWW^{+}97, W^{+}00, Cal91, Dan93, ES97, Gil03, LL94b, LL94a, PBA92, Ano94a, Ano05b, The21i, The21j, DF06, Nee00, Pla99, The96, The00, The03, The06b, The10, Ano19a, Ano20a]. Standardizacija [J^{+}09]. Standardization [BEE95, Lua90]. standardized [Dic22]. Standardizing [GC96, Kim92b]. Standards [Ano02b, BM97b, Che01, DNYA96a, DYN97a, DNYA96b, DYN97c, DYN97b, Dür98d, Fei94, GC96, Hos24, Mey01, Spr96, Cle88, CMD^{+}97, Mic99, N^{+}05a, N^{+}05b, Nor96, PCM^{+}97, Ngô05, Wu00, Gwy93]. Standorten [Ano07b]. staroslovenskog [J^{+}09]. State [CMD^{+}97, DO00, PCM^{+}97]. State-of-the-art [CMD^{+}97, PCM^{+}97]. Status [Dür97c, Dür98a, DYN98]. Staying [Swo05]. Steganographic [MJF07]. Step [Bec98, Hod98, OL98]. Steps [Bec97]. STIX [Bee07]. Storage [Hon98a]. Storing [Wri23]. Strategies [Gol97c, SJ00]. Stream [Kue05, MLJ13, MLJ17]. Streams [Ran97b, Eck97]. String [Ull21, WKH97, PDR17]. Strings [WFD08].

Structure [Ali05]. structures [BI95]. Studies [Had01, Keb97, KT97, Per09]. Study [Ano97a, Bin97, GN98a, GN98b, Gre98a, Gre98b, Har99, Het98, Kap01c, Saa08]. Style [Abu04, Lil97, Lil98a, Hag13]. Subsetting [TK01]. Success [SJ00]. Such [CPJ^{+}98]. Sudoc [Cha07]. Summary [McG04]. Supplemental [dC00]. Supplementary [DW02]. Support [Ano94a, Anoxxa, Atk96, Atk97b, Atk97a, BFS02, Bis97b, Bis97a, Buc96a, Buc96b, Buc96c, BM97c, BM97b, Buc97, Dem98, Gid01, Gri98b, Gri98c, Jun97, Kap01a, Kap01b, Lei96a, Lei96e, Lei96b, O’C98, Oet12, PDR17, Shi98a, Sun98b, Sun98c, Ten97b, UY98, Unixxb, Wel96, Woo99, Yan98, DG95, Lem97, MR08, Per09, Wu00]. Supported [JBGC97]. Supporting [BB97, BBM98, Edb97b, Gol97a, Gol98b, Mur96, Ran96]. Surrogate [Kap01a]. Survey [Edb97a, Ewe04b, Gol98c, Ewe04a, Fre06]. Syllabaries [Aki80]. syllabics [Ano22, Dic22]. symbol [Jen70]. Symbology [Hoh97]. symbols [Fre06]. Symposium [Ben97, BJ93, BJ94]. Syntax [BN98, Dav05a, The21f]. System [Ano07b, Fen96, Hiu96, McG96, Ran97c, Ran96, RSLO04, Sun96, Sun98a, UY98, Coz17, Czy98, DGK96, Dic22, Har14, Wed96, Hiu97c, Mat96, Mik98, PZ92]. Systems [Age91, Aki80, Ano19b, Cor00, Coy06, Het98, IEE90, Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c, McK97b, McK98a, Ran97a, SD95, Cha12, CMD^{+}97, DB96, DG95, ME94, PCM^{+}97, Wu00].

Tˆ óng [Ngô05, N^{+}05a]. TAB [Che04a]. Tableaux [Ano20b]. Tables [Dav99b]. tablets [Dic22]. Tagging [WA99, WA98a]. Tags [WA98b]. Taiwan [MfFH07]. Taiwanese [Jor02]. takes [Ano02b]. Taligent [DGK96]. TAM [Che04b]. Tamil [Kap05, SSB^{+}05]. Tastaturtreiber [See13]. TCVN [N^{+}05a, N^{+}05b, Ngô05]. Teach [Ano05b]. Teach-in [Ano05b]. Technical [DW06, Jen08, Mil98a, SD06, SS98, The21k, The21l, The21m]. Techniques [Ano09, C^{+}09]. Technological [Tan94]. technologically [Abu04]. technologically-friendly [Abu04]. TECHNOLOGIEN [Ano06b]. Technologies [Dür98c, Sug96, Kew05b]. Technology [DDG96, Uns96]. Teil [Ano06b]. telegrams [Ano89, Ano07]. Telegraphic [Ano89, Ano07]. Template [Gil97]. templates [Hag13]. Tenth [Uni97a]. Teradata [Rob98b]. Terminal [PZ92, dC00]. Testing [Dem98, Dom96]. TEXShop [SK18]. Text [Abr98, ACE^{+}08, Ano96c, Bec98, BBM98, Che98, CPJ^{+}98, DO00, Dav01d, Dav03b, Ete05, For96, Gil98a, HNTT96, Har97b, Hod98, Hos24, Ish13, Liu97, Lu01, McG96, Mud97, Ned04, PDR17, WA99, Ritxx, RM97, Sar96, Sar97, Sar98b, Sar06, Sui01, The21e, WA98a, Yon97, BEE95, Bee19, Bee20, Box95, DGK96, ES97, GV92, PWC12, SSB^{+}05]. text-based [PWC12]. Texts [Nic98b, MNSJ99a, MNSJ99b]. Thai [Kun94]. their [DNYA96a, DYN97a, DNYA96b, DYN97c, DYN97b, Dür98d, Mey01]. There [CPJ^{+}98]. these [Gid01]. Thing [CPJ^{+}98]. Think [DL19]. third [Uni03b]. Thirteenth [Uni98b]. Tibetan [BDA^{+}92]. Time [Tex01, Bee00]. Tips [Sar98b, Dav97b]. Tipster [DO00]. To-Date [McK96a, McK97a]. tofu [Mit21b]. Tokyo [Uni94a, Uni98a]. TOM [Sar96, Sar97]. Tong [N^{+}05b]. Tool [FDW06]. Tools [Gol97b, LBD08]. Topics [KH98]. Traditional [Har95, Har97a]. Transcoding [CL23, LM22]. Transformation [Ano06c, GD94b, GD97, Cri05, Uma98, Yer03, Wu01, Wu02, Yer96, Yer98a]. Transition [DDG96, Wu00]. Translatable [Ish98b]. Translating [Ano95b, RSLO04, SJ00]. Translation [Mik98, Tor96, Whi05a]. translator [Kwa95a]. translators [Ess98]. Trenches [Fre97a, Rob98a]. Trends [Ste96]. Tricks [Sar98b]. Trouble [Mee97]. true [N^{+}05a, Ngô05, N^{+}05b]. TrueType [Pet93c]. Truly [Mud97]. Truth [Kid98]. TSCII [Ned04]. Tuning [Ano06b]. Turkey [BM98c]. Turku [MVM09b]. turned [Pik12]. Tutorial [Fre97c, HK98, Lei96f, Ada94]. tutorials [Uni96b]. Twelfth [Uni98a]. Twentieth [Uni02a]. Twenty [Uni02b, Uni03b, Uni03a, Uni04a, Uni04b, Uni02c, Uni05b, Uni05a, Uni06]. Twenty-eight [Uni05a]. Twenty-fifth [Uni04a]. Twenty-first [Uni02b]. Twenty-fourth [Uni03a]. Twenty-ninth [Uni06]. Twenty-second [Uni02c]. Twenty-seventh [Uni05b]. Twenty-sixth [Uni04b]. Twenty-third [Uni03b]. Twi [BMCM06]. Two [Ano97a, Mey01]. Type [Abu04, Ano03a, Pog03, BBB^{+}02, N^{+}05a, Ngô05, N^{+}05b]. Typecase [Har95, Har97a]. types [Abu04, KBKHF16]. Typesetting [Bee19, Bee20, Mac86, Nak98, Saj19, Coz17, Har14, Kew06]. Typographique [Pog03]. Typography [Har99, Mil97a, Sui98, STS03].

u [J^{+}09, Jen08, Jen12, The21h]. U-source [Jen08, Jen12, The21h]. uberwindet [Ano07b]. UCS [Uma98]. UK [Ano95b]. Ultracode [Hoh97]. umsteigt [Ano07a]. Umstellung [Ano06b]. unarodnog [J^{+}09]. underpin [CMD^{+}97, PCM^{+}97]. Understanding [Lun93a, Nor96, Tex01]. Unencoded [Lun98a]. Uni [Bur01]. UNICODE [Wei07, Ada94, Ano94a, Ano94b, Ano02b, Ano06a, Ano06b, Ano06c, Ano07b, Ano07a, Ano09, Ano19a, Ano20a, BP11, Che94, Czy98, Kew05b, Mil03, Pla99, See13, Uni94a, Uni94b, Uni95a, Uni96b, Uni97b, Uni98a, Uni98b, Uni99b, Uni99a, Uni00b, Uni00a, Uni01a, UIW92, Uni01b, Uni02b, Uni02a, Uni02c, Uni03b, Uni03a, Uni04a, Uni04b, Uni05b, Uni05a, Uni06, Uni11, Uni12, Ano89, Ano94c, Ber09, Har07a, Kwa95a, La 93, MfFH07, Swo05, Ada94, Ado98, Ali94, Ali05, Ali10, AKW07, Ams97, ABP04, AGH95, AH03, Ann97, Ano07, Ano91, Ano95a, Ano96c, Ano97b, Ano97c, Anoxxb, Ano01, Ano02a, Ano05a, Ano05b, Ano08, Ano20b, AH05, Bar97, BGD06, Bee00, BFS02, Bee07, BBB^{+}02, Bet97a, BB93, Bet97b, Bie23, BH93, Bin97, BMCM06, Bos98, BMN98a, BMN98b, Box95, Bri96, Bri97b]. Unicode [Bri97a, Bro98, Buc96a, Buc96b, Buc96c, BM97c, BM97b, Buc97, Buc98, BM98a, BM98b, Bul05, BG97, Cai94, Cal91, Cav97a, Cha07, Che04a, Che04b, Che98, Che97, Chi17, C^{+}09, Cla97, CL23, CMD^{+}97, Cob98, The21i, The21j, Coy06, Dan93, Dan08, DW88, DC90, Dav98c, Dav98b, Dav99c, DD00, Dav00b, DS01, Dav01a, DEFM01, Dav03c, Dav04c, DS05b, DS05c, DS05d, DF06, DW06, DW08, DW12, DS12, Dav15, DE16, Dav21, Dav22, Des05, Des08, DM97, Dür94, Dür97c, DDHF99, DF02, Eev15, Eml06, ES97, Eve02, Ewe04a, Ewe04b, Fai95, Fei94, FB98, FV97, Fly98, For96, Fow96, Fow97, Fre88b, Fre96b, Fre97a, Fre97b, Fre97c, Fre97f, Fre98b, Fre02b, FMW06, FDW06, Gar12, DYN98, Gav96, Gil97, Gil03, GN98a, GN98b, GD94b, GD94a, GD97, Gol98a]. Unicode [Gon98, Gra00, Gri98b, Gri98c, Hag16, Hal94a, Hal96a, Hal96b, Hal97a, Hal97b, Hal98b, Hal98c, Hal98a, Har95, Har97a, Har07b, Het98, Hiu96, Hiu97c, HK98, Ho96, Hod98, Hoh97, Hol03, Hon96, Hon98a, Hon98b, HT04, Hos24, Huf17, Ish13, Jen97c, Jen97a, Jen98, Jen08, JC08, JBGC97, JBC98, Jun97, Kap01b, Kap01c, Keb97, KT97, Kew05a, Kew06, Kim92b, Kin96, Kin12, Kli98a, Kli98b, Kor06, Kri97, Ksa96, Ksa97a, Ksa97b, Ksa98a, Ksa98b, Kue05, Kuh99a, Kuh99b, Kun94, Küs05, Kwa95b, LaB00, Leg91, Lei96a, Lei96e, Lei96b, LM22, Leo98, Ler03, Ler13, Li98, LL94b, LL94a, Liu97, Lon00, LBD08, MJF07, Mah09, Man97b, Man97a, Man98c, Man98a, Man98b, Man01, Mar98, MH95, McC96, McK96a, McK96b, McK97c, McK97a, McK97d, McK97e, McK97b, McK98b, MR08, MVM09a, MVM09b]. Unicode [Mit21b, Moo97, Moo98a, Mor98, Mul08, Mun19, Mur98, Ned04, Nee00, Nes94, N^{+}05a, N^{+}05b, Nie96, Nie97a, Nie97b, Nie98, NK07, Noj01, Oet12, Pan98e, Pan98c, Pan98d, PCM^{+}97, Pet93a, Pet93b, Pet93c, PWC12, Pot05, Pou02, PHE17, Pra96a, Pra96b, Pra97b, Pra97a, Pra98a, Pra98b, Pra98c, WA99, Cos03, McG04, Cri05, Ran97a, Ran97b, Ran97c, Ran96, Ree17, Res98, Rob98b, Rob98a, Rob10, Rob15, RSLO04, RM16, Rub96, Saa08, Saj19, Sar96, Sar97, Sar98b, Sar06, SD02, SD06, See13, SSB^{+}05, Shi98a, Sid06, Sui98, Sun00a, Sun98c, Swi09, Tan13, Tan96, Tan97, Tau16, Ten97a, Ten97b, Ten98, Tex98a, Tex98b, Ngô05, Tiv97, The90, Uni94c, Ull21, UY98, The96, The00, Uni07, The03, The06a, The06b, The10, The16a, The21b, The21c, The21d, The21e, The21f, The21g, The21h]. Unicode [The21k, The21l, The21m, Uns96, Vac91, Van94, Wan97, Wel97b, Wel97a, WFD08, WA98a, WD98b, WF04, Whi05a, Whi05b, Whi11, WG05, WK01, Wis01, Wol96, Wol97, WWW^{+}97, W^{+}00, Woo99, Wri20, Wri22, Wri23, Wu00, Yun00, YY97, Yan98, Yas05, Yer96, Yer98b, Yon97, Zha97a, Zha97b, Zif01, dC00, Mil03, Pog03]. Unicode-Anpassungen [Ano07a]. Unicode-Based [FDW06, Gav96, RSLO04]. Unicode-Character [Yas05]. UNICODE-compliant [Swo05]. Unicode-enabled [JBC98]. Unicode-land [Wri22]. Unicode-like [Bie23]. Unicode-Tastaturtreiber [See13]. Unicode/ISO [Ada94, Bri96, Bri97a, Uni96a].

Unicode/ISO10646 [Uni98a].

unicodefonttable [Mit21a]. Unicodeland [Wri15]. Unicode(R) [Ano15, The16b, The16c, The16d]. Unicode^{T M} [Gra99]. Unified [HNTT96, Sun96, Shi98b]. uniform [Leg91]. Unihan [Jen97b, JC08, Kid98, The21g]. Unikodu [J^{+}09]. UNIMARC [Wei07, Har07a, MfFH07, Wei07]. Uniprep [Dür94]. Unique [Leg91]. Uniscribe [Bro98]. UniSpaCh [PWC12]. Universal [Ano89, Atk96, Atk97b, BB93, Bre98, Coy06, DO00, IEE90, Pol91, Ran97b, Ano07, Kin12, Kwa95b, Leg91, Mik98]. UNIverse [CMD^{+}97, PCM^{+}97]. University [DO00]. Unix [Czy98, Hiu96, Hiu97c, Lei96a, Lei96b, Li98, Ran96, Ric98, Sun96, Kuh99b, Lei96e, PB21]. Unix/Linux [Kuh99b]. unterschatzen [Ano07a]. Unveils [Ano05a]. Update [Bos97a, DF06, DW06, Gwy93]. upTEX [Tan13]. urls [Pot05]. USA [Uni02a, Uni03a, Uni04b, BJ93, IEE90, Uni96c, Uni97b, Uni99b, USE93, Uni04a, Uni05a, Uni06]. Use [Ano89, Cob98, Hon98a, McK97e, McK98b, Ano07, BEE95, Dic22, ES97, Fre06, Gra99]. User [Bec98, Dem98, Pan98c, Pan98d, Yan98, dN96]. Users [Yon97]. Using [Atk97b, CS01, GD94a, LBD08, Per09, Nas96, NK07, Noj01, PWC12, SSB^{+}05, STS03]. Utah [SC98]. UTF [Ano03b, Ano06c, Dav99d, Dav01e, Due97b, Dür97b, GD94b, GD97, Hag07, Jun97, KL21, Kuh99b, Phi01, Pik03, Pik12, Yer03, Cri05, RGN16, Sch04, Uma02, WKH97, Yer96, Yer98a, Zif01]. UTF-16 [Ano06c, Phi01, Sch04]. UTF-18 [Cri05]. UTF-32 [Dav99d, Dav01e]. UTF-7 [GD94b, GD97]. UTF-8 [Ano03b, Due97b, Dür97b, Jun97, KL21, Kuh99b, Pik03, Pik12, Yer03, RGN16, WKH97, Yer96, Yer98a, Zif01]. UTF-9 [Cri05]. UTF-EBCDIC [Uma02]. Utilities [Ric98]. UX [Kli98a, Kli98b]. v3 [WKH97]. VA [BJ93]. Validating [KL21]. Variants [Dür98c]. Variation [HM06]. Variations [Saa08]. vector [MLJ13, MLJ17]. Vendors [McK97e, McK98b]. verbatim [Bee19]. verknupfen [Ano07b]. verschiedenen [Ano07b]. Version [Ano05b, Ano19a, Ano20a, Atk96, The21i, The21j, Fre96b, Fre97f, Moo98a, Uni94c, Uni91, The91, Uni95b, The96, The00, The03, The06b, The10, Uni11, Uni12, The16b, The16c, The16d, Uni20, Tan13]. Versus [LaB00]. Vertical [Ete05, Ish13]. via [Hag16]. Victor [Ano03a]. Vietnam [DDG96]. Vietnamese [N^{+}05a, Che97, Lun99, N^{+}05b, Ngô05]. View [Rob98a]. Viewing [Pet93c, RSLO04]. Visual [DD11, Hip94, Kap01c]. Volume [Cha97, The21i, The21j]. Volumes [The91]. voranzutreiben [Ano06a]. Voraussetzung [Ano07b]. vs [Dür98b, Yas05]. VSCII [N^{+}05a, Ngô05, N^{+}05b].

W3C [Bos97a]. WAIT [Top01]. Warp [Atk96]. Was [Ano06b]. Washington [Uni02a, Uni04a]. Way [Ams97, CPJ^{+}98]. Weaving [Ano96d, Ano96e, Ano97d, Ano97e, Ano97f, DNYA96a, DYN97a, DNYA96b, DYN97c, DYN97b, Dür98d]. Web [Bos97b, DYI^{+}04a, DYI^{+}04b, Lil98a, Uni99b, Uni99a, Uni00b, Uni00a, Uni01a, Uni01b, Uni02b, Uni02a, Uni02c, Uni03b, Uni03a, Swi09, Ams97, Ano96d, Ano96e, Ano97c, Ano97g, Ano97d, Ano97e, Ano97f, Bar98, Ben97, Bre98, DNYA96a, DYN97a, DNYA96b, DYN97c, DYN97b, Dür98c, Dür98d, DYI^{+}02, Jor02, Lil97, Moo97, RSLO04, RC98, Woo99, Yon97, van98]. Web-and [RSLO04]. Webg [FDW06]. Week [Uni03b]. Welcome [Kob98, Moo98b]. Well [Dav97b]. Well-mannered [Dav97b]. Welt [BP11]. weltweit [Ano07b]. Wer [Ano07b, Ano07a]. which [Che97, CMD^{+}97, PCM^{+}97]. who [BEE95]. whose [Bee00]. Wide [Bos97b, DYI^{+}02, DYI^{+}04a, DYI^{+}04b, Jor02, Pet93b, RC98]. Width [Fre88b, Fre00a]. Will [Kin12, Ano07b, CMD^{+}97, Dic22, PCM^{+}97]. Williamsburg [BJ93]. Win [Rev93]. Win32 [Hel93, Bro98, Bro97, Hal96a, Hal96b, Hal97a, Hal97b, Hal98b, Hal98c, O’C98, Ric94, Ric95]. WinCALIS [Kun94]. Window [PZ92, Hiu96, Hiu97c, Mat96]. Windows [Bis97b, Bis97a, Hel92, MH93, AKW07, Agr93, Ams97, BBM98, Box95, Hal94a, Hal94b, Hal98a, H^{+}95, Hud00, Kan95, Li98, MP93, Nie96, Nie97b, Nie98, Pet93c, Ric94, Ric95, WK01, Wis01]. Windows95 [Nie98]. Windoze [CPJ^{+}98]. Winter [USE93]. Without [Oet12]. Word [Bec84, Ran97a, Ran97c, Bec87]. Words [Ano89, PM17, Ano07]. Work [Nic98b]. Workshop [Ano94c, Ano02b, SD95, Uni94a, Uni94b, UIW92, Ada94, BM98c, Ano94b]. workstation [ZV97]. World [Aki80, Ano03a, Ano19b, Bec97, Bec98, Bro97, Bun97, CPJ^{+}98, FCJ04, Hod98, Hos24, Keb97, KT97, May97, McK97e, McK97b, McK98b, O’D94, OL98, PT93, PT00, Uni04b, Ano22, BP11, DB96, MAD^{+}06, Ull21, Bos97b, DYI^{+}02, DYI^{+}04a, DYI^{+}04b, Jor02, RC98]. World-Class [McK97e, McK97b, McK98b]. World-Ready [Bec97]. WORLDLY [AKW07]. Worldwide [Atk97a, Che94, Sui98, TS96, The91, Uni95b, Cle88]. Worried [Fow97]. Wrapping [Pra96a, Pra96b]. Write [Ber97, Jen70]. Writing [Age91, Aki80, Ano19b, Bro97, Cor00, Ish96a, Ish97d, Ish97b, Ish98d, Ish98e, Ish98c, DB96, Dic22]. written [Box95]. WWW [BG97, CMD^{+}97, Dür98a, Gil97, Ho96, Ho97, JBGC97, JBC98, Nic97b, PCM^{+}97].

X [Hiu96, Hiu97c, IF97, Mat96, PZ92, See13, Sui98]. X-4051-1995 [Sui98]. X-Window [Hiu96, Hiu97c, Mat96]. X11 [TK01]. xelatex [Bee20]. XLL [BMN98a, BMN98b, BN98, Nic98b]. XML [Bos98, BMN98a, BMN98b, BN98, DDHF99, DF02, HM04, Lev98, Mul08, Nic98b, Noj01, Swo05, YLM^{+}05]. XSL [BMN98a, BMN98b, BN98, Lil98b].

Years [Oet12, Pik12]. yesterday [Pik12]. York [Pog03]. Yunikodo [Yun00]. Yves [Ali10].

Z39.50 [CMD^{+}97, PCM^{+}97]. zbornik [J^{+}09]. Zeichensatze [Ano07b]. zeros [Ram22]. Zhukov [Pog03]. zunachst [Ano07b].

References

Anderson:2004:U

[ABP04] Deborah Anderson, Johannes Bergerhausen, and Ingo Preuß. Unicode. Ed. Waechterpappel, Grimma, Germany, 2004. ISBN 3-933629-12-8. 47 pp. LCCN ????

Abramson:1998:OIB

[Abr98] Dean Abramson. Optimized implementations of bidirectional text layout and bidirectional caret movement. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Abulhab:2004:MTS

[Abu04] Saad Abulhab. The Mutamathil type style: towards free, technologically-friendly ‘Arabetic’ types. Visible Language, 38(3):306–335, September 2004. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V38N3_

2004_E.pdf. See comments [?] and reply [?]. See also [?].

Anan:2008:RJT

[ACE^{+}08] Yasuhiro Anan, Hiroyuki Chiba, Junsaburo Edamoto, Richard Ishida, Keiichiro Ishino, Tatsuo Kobayashi, Toshi Kobayashi, Kenzou Onozawa, and Felix Sasaki. Requirements of Japanese text layout. Working draft, W3C, October 15, 2008. URL

http://www.w3.org/TR/2008/
WD-jlreq-20081015/.

Adams:1994:IUU

[Ada94] Glenn Adams. Introduction to Unicode: Unicode/ISO 10646 implementors’ workshop 5. tutorial. The Unicode Consortium, Cambridge, MA, USA, 1994. ISBN ?? v + 207 pp. LCCN ?? Unicode/ISO 10646 implementors’ workshop tutorial, Tokyo, Japan, April 11–12, 1994.

Adobe:1998:UGN

[Ado98] Adobe Systems, Inc. Unicode and glyph names. World-Wide Web document, Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, December 17, 1998. URL http:

//partners.adobe.com/asn/

developer/type/unicodegn.

html. Document version 1.1. Last updated 17 December 1998.

Agenbroad:1991:NPI

[Age91] James Edward Agenbroad.

Nonromanization: Prospects for improving automated cataloging of items in other writing systems. opinion papers no. 3. In Meeting of the Library of Congress Cataloging Forum (Washington, DC, July 22, 1991), page 20. Cataloging Forum Steering Committee, Library of Congress, Washington, DC 20402, USA, 1991. ISBN ???? LCCN ????

Andre:1995:ECC

[AGH95] Jacques André, Michel Goossens, and Yannis Haralambous. Editorial: 1 codage des caractères, d’Ascii à Unicode: Journée Omega, Genève, 16 mars 1995. Cahiers GUTenberg, 20:??, 1995. ISSN 1140-9304.

Agrusa:1993:IWN

[Agr93] R. L. Agrusa. Is Windows NT the PC platform for the future? Control Engineering, 40(5):55, 57, 59, April 1993. CODEN CENGAX. ISSN 0010-8049.

Andre:2003:UED

[AH03] Jacques André and Henri Hudrisier, editors. Unicode, écriture du monde. Hermès, Paris, France, 2003. ISBN 2-7462-0594-7. 364 pp. LCCN ????

Asmus:2005:LBP

[AH05] F. Asmus and A. Heninger. Line breaking properties: Unicode Standard Annex 14, line breaking properties. Technical report, Unicode Consortium, 2005. URL http://www.

                                        unicode.org/reports/tr14/.

Asthana:2012:SCD

[AH12] Puneet Asthana and David Haslam. SAP certified development associate — ABAP with SAP NetWeaver 7.02. SAP certification success guide. Galileo Press, Bonn, Germany, second edition, 2012. ISBN 1-59229-435-9. 634 pp. LCCN TK5105.8885.S24 A88 2012.

Akira:1980:WSW

[Aki80] Nakanishi Akira. Writing Systems of the World: Alphabets, Syllabaries, Pictograms. C. E. Tuttle Co., Rutland, VT, USA; Tokyo, Japan, 1980. ISBN 0-8048-1293-4 (hardcover), 0-8048-1654-9 (paperback). 122 pp. LCCN Z40.N2613 1980. US$19.50. Translation of original Japanese edition, Sekai no moji, 1975.

Allen:2007:WWE

[AKW07] Julie D. Allen, Michael S. Kaplan, and Cathy Wissink. WORLDLY WINDOWS — extend the global reach of your applications with Unicode 5.0. MSDN magazine, 95:??, January 2007. ISSN 1528-4859.

Aliprand:1994:SMA

[Ali94] J. M. Aliprand. Unicode and ISO/IEC 10646: An overview. In McCallum and Ertel [ME94], pages 87–102. ISBN 3-598-21797-8. ISSN 0344-6891. LCCN Z699.5.F67 A872 1994.

Aliprand:2005:SCM

[Ali05] Joan M. Aliprand. The structure and content of MARC 21 records in the Unicode environment. Information Technology and Libraries, 24(4):170, ???? 2005. CODEN ITLBDC. ISSN 0730-9295.

Alix:2010:YDB

[Ali10] Yves Alix. Yves Desrichard, bibliothèques et Écritures, d’ASCII à Unicode. Bulletin des bibliothéques de France, 55(1): 82, ???? 2010. ISSN 0006-2006.

Alvestrand:1998:RIP

[Alv98] H. Alvestrand. RFC 2277:

IETF policy on character sets and languages, January 1998.

URL ftp://ftp.internic.
net/rfc/bcp18.txt;   ftp:/
/ftp.internic.net/rfc/rfc2277.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/bcp18.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/
rfc2277.txt.

Amstein:1997:URW

[Ams97] Peter Amstein. Unicode: the right way to develop Web applications for Windows. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Annobile:1997:LU

[Ann97] Michael Annobile. LISA and Unicode. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Anonymous:1889:UUT

[Ano89] Anonymous. UNICODE: The Universal Telegraphic Phrase-Book: a Code of Cypher Words for Commercial, Domestic, and Familiar Phrases in Ordinary Use in Inland and foreign telegrams. Cassell and Company, Limited, London, Paris, New York, and Melbourne, sixth edition, 1889. x + 110 pp.

Anonymous:1900:UUT

[Ano07] Anonymous. Unicode, the universal telegraphic phrase-book, a code of cypher words for commercial, domestic, and familiar phrases in ordinary use in inland and foreign telegrams. Cassell, London, UK, 1900, 1907. xxiii + 358 pp.

Anonymous:1991:HCU

[Ano91] Anonymous. Here comes Unicode. New York Times, ??(??): F2, February 24, 1991. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.
proquest.com/docview/108788404.

This appears to be second mention of Unicode in this newspaper, and the first in a headline. See [Pol91] for the first report of Unicode.

Anonymous:1994:IAS

[Ano94a] Anonymous. Increased Asian support for The Unicode Standard. Information Standards Quarterly, 6(2):11–??, April 1, 1994. CODEN ISQUEK. ISSN 1041-0031.

Anonymous:1994:SUI

[Ano94b] Anonymous. Sixth Unicode Implementors’ Workshop held in Santa Clara. Information Standards Quarterly, 6(4):15– ??, October 1, 1994. CODEN ISQUEK. ISSN 1041-0031.

Anonymous:1994:UWS

[Ano94c] Anonymous. UNICODE Workshop: September 8–9. Information Standards Quarterly, 6 (3):10–??, July 1994. CODEN ISQUEK. ISSN 1041-0031.

Anonymous:1995:LAL

[Ano95a] Anonymous. LAN answers — lending a hand with Unicode, MIGRATE, Ethernet-II, and directory entries. LAN times, 12 (25):128–??, 1995. ISSN 1040-5917.

Anonymous:1995:TCP

[Ano95b] Anonymous, editor. Translating and the Computer 17. Papers from Aslib Conference. London, UK, 9–10 November, 1995. Aslib, London, UK, ???? 1995.

ISBN 0-85142-351-5. LCCN P308 .T724 1995.

Anonymous:1996:ALL

[Ano96a] Anonymous, editor. Association for Literary and Linguistic Computing and Association for Computers and the Humanities: Joint international conference — June 1996, Bergen, Norway, number 70 in Report Series — Norwegian Computing Centre for the Humanities 1996. Norwegian Computing Centre for the Humanities, Bergen, Norway, 1996. ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Anonymous:1996:JIA

[Ano96b]  Anonymous.  Java interna-

tionalization architecture — continued. In Unicode Consortium [Uni96c], pages A4– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Anonymous:1996:TSA

[Ano96c] Anonymous. Text searching across multiple character sets in Unicode. In Unicode Consortium [Uni96c], pages A5– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Anonymous:1996:WMWa

[Ano96d] Anonymous. Weaving the multilingual Web: An open discussion forum (BOF). In Unicode Consortium [Uni96c], pages A7–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Anonymous:1996:WMWb

[Ano96e] Anonymous. Weaving the multilingual Web: An open discussion forum (BOF) — continued. In Unicode Consortium [Uni96c], pages A8– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Anonymous:1997:CST

[Ano97a] Anonymous. Case study of two multiscript locales in Solaris. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Anonymous:1997:UFP

[Ano97b]  Anonymous.  Unicode and

fonts. In Unicode Consortium [Uni97b], pages B4–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Anonymous:1997:UMW

[Ano97c] Anonymous. Unicode meets the Web. Language International, 9 (3):22–23, ???? 1997. CODEN LAINE2. ISSN 0923-182X.

Anonymous:1997:WMWa

[Ano97d] Anonymous. Weaving the multilingual Web: An open discussion forum (BOF). In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Anonymous:1997:WMWb

[Ano97e] Anonymous. Weaving the multilingual Web: An open discussion forum (BOF) — continued. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Anonymous:1997:WMWc

[Ano97f] Anonymous. Weaving the multilingual Web: An open discussion forum (BOF) — continued. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Anonymous:1997:WIO

[Ano97g] Anonymous. Web internationalization — an open forum (BOF). In Unicode Consortium [Uni97b], pages TA4–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Anonymous:19xx:SIA

[Anoxxa] Anonymous. Support for interlinear annotations. Unicode Technical Report 12, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 19xx.  URL http:/
/www.unicode.org/unicode/
reports/techreports.html;
 http://www.unicode.org/unicode/
reports/tr12/. In progress.

Anonymous:19xx:URE

[Anoxxb] Anonymous. Unicode regular expression guidelines. Unicode Technical Report 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 19xx.  URL http:/
/www.unicode.org/unicode/
reports/techreports.html.

In progress.

Anonymous:2001:PDU

[Ano01] Anonymous. Proposed draft Unicode 3.2. Unicode Technical Report 28, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, December 21, 2001.

                                        URL http://www.unicode.
                                        org/unicode/reports/tr28/.

Anonymous:2002:LU

[Ano02a] Anonymous. LaserHebrew in Unicode. Technical report, Linguist’s Software, Edwards, WA, USA, 2002.

Anonymous:2002:LWS

[Ano02b] Anonymous. LRC Workshop Standards in localisation — an industry takes shape: in cooperation with the 21st International Unicode Conference The Burlington Hotel — Dublin (Ireland), 13 May 2002. Technical report, Localisation Research Centre, University of Limerick, Limerick, Ireland, 2002. 15 pp.

Anonymous:2003:GPT

[Ano03a] Anonymous. Gentium: Providing type to the world. interview with Victor Gaultney, creator of Gentium. The LISA Newsletter, 12(4.3):??, November 4, 2003.

                                        URL http://www.lisa.org/
                                        archive_domain/newsletters/
                                        2003/4.3/gaultney.html.

Anonymous:2003:U

[Ano03b] Anonymous. UTF-8. Wikipedia article., December 17, 2003.

                                        URL https://en.wikipedia.
                                        org/wiki/UTF-8. Under fre-

quent revision, with numerous authors/contributors.

Anonymous:2005:AUU

[Ano05a] Anonymous. AquaBrowser unveils Unicode options. Advanced technology libraries, 34 (10):5, ???? 2005. ISSN 0044-636X.

Anonymous:2005:TWV

[Ano05b] Anonymous. Teach-in — what is version control, what is Unicode?, the Unicode Standard. Software world, 36(2):7, ???? 2005. ISSN 0038-0652.

Anonymous:2006:IUA

[Ano06a]  Anonymous.  International-

isierung mit Unicode – Der Armaturenhersteller Hansgrohe migrierte von R/3 auf Mysap, um die Internationalisierung voranzutreiben. Computerwoche, 33(35):32, ???? 2006. ISSN 0170-5121.

Anonymous:2006:PTC

[Ano06b]  Anonymous.  PRODUKTE

& TECHNOLOGIEN — CW-Serie SAP-Tuning — Teil 6: Unicode beseitigt Sprachbarrieren. Was im Fall einer Umstellung zu beachten ist. Computerwoche, 33(15):26, ???? 2006. ISSN 0170-5121.

Anonymous:2006:UBU

[Ano06c] Anonymous. UTF-16 (16-bit Unicode Transformation Format), 2006. URL http://en.

wikipedia.org/wiki/UTF-16.

Anonymous:2007:SUW

[Ano07a] Anonymous. SAP und Unicode — Wer von R/3 auf ERP 6.0 umsteigt, sollte den Aufwand fur Unicode-Anpassungen nicht unterschatzen. Computerwoche, 34(44):20, ???? 2007. ISSN 0170-5121.

Anonymous:2007:SEK

[Ano07b] Anonymous. SCHWERPUNKT: ERP-KONSOLIDIERUNG — Unicode uberwindet Sprachbarrieren — Wer seine Geschafte global betreibt, muss sein ERP-System an den verschiedenen Standorten weltweit verknupfen. Voraussetzung dafur ist jedoch eine einheitliche Sprache, um Informationen austauschen zu konnen. Unicode sorgt für kompatible Zeichensatze, will aber zunachst sorgfaltig implementiert sein. Computerwoche, 34(4):22, ???? 2007. ISSN 0170-5121.

Anonymous:2008:KDH

[Ano08] Anonymous. Kangxi dictionary: hytung Unicode database revision. ????, ????, 2008. ISBN 7-80230-904-2. 58 + 2058 pp. LCCN ???? Includes CD-ROM.

Anonymous:2009:DOT

[Ano09] Anonymous. DB2 Optimization Techniques for Sap Database Migration and Unicode Conversion. IBM, International Technical Support Organization, ????, ??, USA, 2009. ISBN 0-7384-3308-X. xiv + 278 pp. LCCN ????

Anonymous:2015:U

[Ano15] Anonymous. Unicode(R) 8.0.0. Web site, June 17, 2015.

URL http://www.unicode.
org/versions/Unicode8.0.0/

. The version 8 standard is only available in electronic form.

Anonymous:2019:AUS

[Ano19a]  Anonymous.  Announcing

The Unicode Standard, version 12.0. Web site, March 5, 2019. URL http://

blog.unicode.org/2019/03/
announcing-unicode-standard-
version-120.html.

Anonymous:2019:WWS

[Ano19b] Anonymous. The world’s writing systems. Web site, January

15, 2019. URL http://www.
worldswritingsystems.org/;
http://www.youtube.com/watch?
v=CHh2Ww_bdyQ;   https://
shop.designinmainz.de/produkt/
the-worlds-writing-systems-
poster/.

Anonymous:2020:AUS

[Ano20a] Anonymous. Announcing The Unicode Standard, version 13.0. Web site, March 11, 2020.

URL http://blog.unicode.
org/2020/03/.

Anonymous:2020:TCU

[Ano20b] Anonymous. Tableaux des caractères Unicode 13.0. Unicode Web site, August 2020.

URL https://www.unicode.
org/charts/fr/.

Anonymous:2022:SSD

[Ano22] Anonymous. Special syllabics developed in Nunavut mean Nattilingmiutut can be read anywhere in the world. CBC Radio Web site., February 27,

2022.  URL https://www.
cbc.ca/news/canada/north/
nattilingmiutut-hadlari-typesetting-
nilaulaq-aglukkaq-unicode-
font-1.6363538.

Asthana:2015:SCD

[Ast15] Puneet Asthana. SAP certified development associate — ABAP with SAP NetWeaver 7.4. Rheinwerk Pub., Bonn, Germany, third edition, 2015. ISBN 1-4932-1212-5, 1-4932-1213-3 (e-book), 1-4932-1214-1 (print and e-book). ???? pp. LCCN TK5105.8885.S24 A88 2015.

Atkin:1996:ULS

[Atk96] S. Atkin. Universal language support in OS/2 Warp Version 4. In Unicode Consortium [Uni96c], pages C2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Atkin:1997:IWL

[Atk97a] S. Atkin. IBM worldwide language support. In Unicode Consortium [Uni97b], pages C1–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Atkin:1997:DAU

[Atk97b] Steve Atkin. Developing applications using universal language support. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Atkin:1998:IGLa

[Atk98a] Steven Atkin. IBM graphical locale builder. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Atkin:1998:IGLb

[Atk98b] Steven Atkin. IBM graphical locale builder. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Atkin:2001:IBR

[Atk01] Steven Atkin. Implementations of bidirectional reordering algorithms. In Unicode Consortium [Uni01a], page ?? ISBN ????

LCCN ???? URL http://www.
cs.fit.edu/~satkin/.

Anderson:2005:RCN

[AwRMW05] Deborah Anderson, with Rick Mc-Gowan, and Ken Whistler. Recommendations for creating new orthographies. Unicode Technical Note 19, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, January 5, 2005.

URL http://www.unicode.
org/notes/tn19/;   http://

www.unicode.org/notes/tn19/
tn19-1.html.

Bright:2020:ODP

[BAP20] Walter Bright, Andrei Alexandrescu, and Michael Parker. Origins of the D programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):73:1–73:38, June 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3386323.

Barry:1997:RCS

[Bar97] Randall K. Barry. The role of character sets in library automation: The development of 8-bit sets and Unicode. International cataloguing and bibliographic control: Quarterly bulletin of the IFLA UBCIM Programme, 26(1):14–16, October 1997. CODEN ICBCEH. ISSN 1011-8829b.

Barrett:1998:BMW

[Bar98] Alan Barrett. Building multilingual Web applications. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Bettels:1993:UUC

[BB93] Jörgen Bettels and F. Avery Bishop. Unicode: a universal character code. Digital Technical Journal, 5(3): 21–31, Summer 1993. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/info/

DTJ/v5n3/UNICODE_A_UNIVERSAL_
CHARACTE_01jan1994DTJB02P8.
ps;   http://www.digital.
com:80/info/DTJB02/DTJB02SC.
TXT.

Bishop:1997:SML

[BB97] F. Avery Bishop and Lori Brownell. Supporting multiple languages and character sets with Internet Explorer. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Berry:2002:LCT

[BBB^{+}02] John D. Berry, Mark Batty, Robert Bringhurst, Maxim Zhukov, Matthew Carter, John Hudson, Gayanch Bagdasaryan, Vincent Figgins, and Roger Robertson, editors. Language culture type: international type design in the age of Unicode. ATypI, New York, NY, USA, 2002. ISBN 1-932026-01-0. ix + 373 pp. LCCN Z250.A2 L36 2002.

Bishop:1998:SMT

[BBM98] F. Avery Bishop, David C.

Brown, and David M. Meltzer. Supporting multilanguage text layout and complex scripts with Windows NT 5.0. Microsoft Systems Journal, 13(11):57– ??, November 1998. CODEN MSJOED. ISSN 0889-9932.

Becker:1992:STM

[BDA^{+}92] Joe Becker, Andy Daniels, Lloyd Anderson, Ken Whistler, Peter Lofting, and Rick Mc-Gowan. Sinhala, Tibetan, Mongolian. Unicode Technical Report 2, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 1992.  URL http:/
/www.unicode.org/unicode/
reports/tr2.html.

Becker:1984:MWP

[Bec84] Joseph D. Becker. Multilingual word processing. Scientific American, 251(1):96–107 (Intl. ed. 82–93), July 1984. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
nature.com/scientificamerican/
journal/v251/n1/pdf/scientificamerican0784-
96.pdf.

Becker:1987:AWP

[Bec87] Joseph D. Becker. Arabic word processing. Communications of the ACM, 30(7): 600–610, July 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
28570.html.

Beck:1997:WJN

[Bec97] B. Beck. World-ready Java: The next steps. In Unicode Consortium [Uni97b], pages A16–?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Beck:1998:WRJ

[Bec98] Brian Beck. World ready Java — the next step — part 1: General I18N, text layout and user interface. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Bauwens:1995:SPA

[BEE95] B. Bauwens, F. Evenepoel, and J. J. Engelen. Standardization as a prerequisite for accessibility of electronic text information for persons who cannot use printed material. IEEE Transactions on Rehabilitation Engineering, 3(1):84–89, March 1995. CODEN IEEREN. ISSN 1063-6528.

Beeton:2000:UMC

[Bee00]  Barbara Beeton.  Unicode

and math, a combination whose time has come — Finally! TUGboat, 21(3):176– 185, September 2000. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb21-3/
tb68beet.pdf.

Beeton:2007:APU

[Bee07] Barbara Beeton. The STIX Project — From Unicode to fonts. TUGboat, 28 (3):299–304, 2007. ISSN 0896-3207. URL http://

www.tug.org/TUGboat/tb28-
3/tb90beet.pdf.

Beeton:2015:EC

[Bee15]  Barbara Beeton.  Editorial

comments. TUGboat, 36(3): 180–181, 2015. ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb36-3/tb114beet.
pdf.

Beebe:2019:TMV

[Bee19] Nelson H. F. Beebe. Typesetting multilingual verbatim text with pdflatex. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 18, 2019. 6 pp. URL

https://www.math.utah.edu/
pub/tex/doc/mlverb.pdf.

Beebe:2020:TMT

[Bee20] Nelson H. F. Beebe. Typesetting multilingual text with

lualatex and xelatex. Re-

port, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 08, 2020. 29

pp.  URL https://www.
math.utah.edu/pub/tex/doc/
mlverb.pdf.

Benitez:1996:MPH

[Ben96] M. T. C. Benitez. Multilingual parallel hypertext. In Unicode Consortium [Uni96c], pages A6–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Benitez:1997:WIM

[Ben97] M. T. Carrasco Benitez. Web internationalization & multilingualism symposium — A report. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Berg:1997:JQH

[Ber97] Cliff Berg. Java Q and A: How do I write an international application? Dr. Dobbs Journal, 22(7):105–106, 116, July 1997. CODEN DDJSDM. ISSN 1044-789X.

Berndt:1998:MSS

[Ber98] Matthias Berndt. Microsoft SQL server. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Berry:2009:UBL

[Ber09] D. Berry. UNICODE: Breaking the language barrier. Supply Chain Management Review, 13 (2):8–9, ???? 2009. ISSN 1521-9747.

Berry:2015:PN

[Ber15] Karl Berry. Production notes.

TUGboat, 36(3):268, 2015. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb36-3/
tb114prod.pdf.

Berry:2023:PNa

[Ber23] Karl Berry. Production notes.

TUGboat, 44(1):115, 2023. CO-DEN ???? ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb44-1/tb136prod.
pdf.

Bethers:1997:GDP

[Bet97a]  Dale Bethers.  Global dis-

tributed printing with Unicode. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Bettels:1997:IUL

[Bet97b] Jürgen Bettels. Internationalization and Unicode in LDAP directory services. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc10/program.
html.

Beeton:2002:USM

[BFS02] Barbara Beeton, Asmus Freytag, and Murray Sargent III. Unicode support for mathematics. Proposed Draft Unicode Technical Report 25, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, January 3, 2002.

URL http://www.unicode.
org/unicode/reports/tr25/.

Bunz:1997:UAL

[BG97] Carl-Martin Bunz and Jost Gippert. Unicode, ancient languages and the WWW. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Becker:2006:UG

[BGD06] Joe Becker, Rich Gillam, and Mark Davis. Unicode Guide. BarCharts, Inc., Boca Raton, FL, USA, 2006. ISBN 1-4232-0180-9. 6 pp. LCCN ????

US$5.95. URL http://www.
barcharts.com/default.asp?
page=chart_details&pid=0180-
9;   http://www.unicode.org/
consortium/guideform.html.

Baird:1994:FEP

[BGI94] H. S. Baird, D. Gilbert, and D. J. Ittner. A family of European page readers. In IEEE [IEE94], pages 540–543. ISBN 0-8186-6270-0. LCCN TK 7882 P3 I57 1994. Three volumes. IEEE catalog number 94CH3440-5.

Bigelow:1993:DUF

[BH93] Charles Bigelow and Kris Holmes. The design of a Unicode font. Electronic Publishing—Origination, Dissemination, and Design, 6(3):289– 305, September 1993. CODEN EPODEU. ISSN 0894-3982.

Bigelow:1998:GC

[BH98] Charles Bigelow and Kris Holmes. Glyphs for characters. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.

                                        unicode.org/unicode/iuc13/
                                        program.html.

Brown:1993:EAC

[BHT93] James A. Brown, Brent Hawks, and Ray Trimble. Extending the APL character set. APL Quote Quad, 24(1):41–46, August 1993. CODEN APLQD9. ISBN 0-89791-612-3. ISSN 0163-6006.

Baird:1995:DSP

[BI95] H. S. Baird and D. J. Ittner. Data structures for page readers. In Spitz and Dengel [SD95], pages 3–15. ISBN 981-02-2122-3, 981-02-2122-3. LCCN TK7882.P3I55 1994.

Bien:2022:RPA

[Bie22] Janusz S. Bień. Representing Parkosz’s alphabet in the Junicode font. TUGboat, 43(3): 247–251, 2022. CODEN ???? ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb43-3/
                                        tb135bien-parkosz.pdf.

Bien:2023:CAU

[Bie23] Janusz S. Bień. Creating annotated Unicode-like font charts. TUGboat, 44(1):110–115, 2023. CODEN ???? ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb44-1/tb136bien-
                                        unichart.pdf.

Bingler:1997:DJU

[Bin97] R. Bingler. Developing Javabased Unicode applications: a case study. In Unicode Consortium [Uni97b], pages A12–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Bishop:1997:CSSb

[Bis97a] F. A. Bishop. Complex script support in Microsoft Windows NT. In Unicode Consortium [Uni97b], pages B12–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Bishop:1997:CSSa

[Bis97b] F. Avery Bishop. Complex script support in Microsoft Windows NT. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Borchardt:1993:CAS

[BJ93] F. L. Borchardt and E. M. T.

Johnson, editors. CALICO ’93 Annual Symposium. Computer Assisted Learning and Instruction Consortium 1993 Annual Symposium on ‘Assessment’ (Williamsburg, VA, USA, 8–13 March 1993). Duke University, Durham, NC, USA, 1993. ISBN ???? LCCN ????

Borchardt:1994:CAS

[BJ94] F. L. Borchardt and E. M. T.

Johnson, editors. CALICO ’94: Annual symposium entitled “Human factors” — March 1994, Flagstaff, AZ. ????, ????,

Baelde:1997:RMA

[BM97a] R. Baelde and T. Milo. Rendering multi-level Arabic script. In Unicode Consortium [Uni97b], pages B9–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Buchta:1997:LSUb

[BM97b] Stefan Buchta and Mike McKenna. Language support and Unicode features in the SQL standards. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Buchta:1997:LSUa

[BM97c] Stefan Buchta and Mike McKenna. Language support and Unicode in relational databases and SQL. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Buchta:1998:UPCa

[BM98a] Stefan Buchta and Michael McKenna. Unicode and proposed changes to the SQL standard. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Buchta:1998:UPCb

[BM98b] Stefan Buchta and Michael G. McKenna. Unicode and proposed changes to the SQL standard. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Byrum:1998:MSM

[BM98c] John D. Byrum, Jr. and Olivia Madison, editors. Multi-script, multilingual, multi-character issues for the online environment: proceedings of a workshop, Istanbul, Turkey, August 24, 1995, volume 85 of IFLA publications. K. G. Saur, München, Germany, 1998. ISBN 3-598-21814-1. LCCN Z699.5.F67 M85 1998.

Bodomo:2006:UKA

[BMCM06] A. Bodomo, C. Marfo, A. Cunningham, and S. Y. K. Mok. A Unicode keyboard for African languages: The case of Dagaare and Twi. International Journal of Technology and Human Interaction, 2(1):1–20, ???? 2006. ISSN 1548-3908.

Bosak:1998:UXXa

[BMN98a] Jon Bosak, Makoto Murata, and Gavin Thomas Nicol. Unicode, XML, XLL and XSL. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Bosak:1998:UXXb

[BMN98b] Jon Bosak, Makoto Murata, and Gavin Thomas Nicol. Unicode, XML, XLL and XSL — continued. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc12/program.html.

Bosak:1998:XXX

[BN98] Jon Bosak and Gavin Thomas Nicol. XML, XLL, and XSL: Syntax and semantics. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Bos:1997:UWA

[Bos97a] B. Bos. Update on W3C’s activities on internationalization. In Unicode Consortium [Uni97b], pages A1–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/program.html.

Bos:1997:WWW

[Bos97b] Bert Bos. World Wide Web Consortium activities in internationalization. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ????

                                        URL http://www.unicode.
                                        org/unicode/iuc10/program.
                                        html.

Bosak:1998:KAX

[Bos98] Jon Bosak. Keynote address: XML and Unicode: Implications for global publishing. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Botchkarev:1997:MC

[Bot97a]  O. Botchkarev.  Mail con-

verter. In Unicode Consortium [Uni97b], pages C13–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Botchkarev:1997:MCP

[Bot97b]  O. Botchkarev.  Mail con-

verter. In Unicode Consortium [Uni97b], pages C13–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Box:1995:PKP

[Box95] Don Box. Porting to Klingon — porting source code written to process English language text to other languages is not easily done. Don explains why Unicode is important for Windows developers. C++ report: the international newsletter for C++ programmer, 7(4):15–19, May 1995. ISSN 1040-6042.

Bergerhausen:2011:DSW

[BP11] Johannes Bergerhausen and Siri Poarangan. Decodeunicode: Die Schriftzeichen der Welt. (German) [Decode Unicode: the characters of the world]. H. Schmidt, Mainz am Rhein, Germany, 2011. ISBN 3-87439-813-7. 656 pp. LCCN ????

URL http://www.slanted.
de/eintrag/decodeunicode-
die-schriftzeichen-der-welt;
http://www.typografie.de/
product_info.php?products_
id=1421&language=en.

Beeton:2016:CMT

[BP16] Barbara Beeton and Richard Palais. Communication of mathematics with TEX. Visible Language, 50(2):40–51, August 2016. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://

                                        visiblelanguage.herokuapp.
                                        com/issue/202/article/1362;
                                         https://s3-us-west-2.amazonaws.
                                        com/visiblelanguage/pdf/V50N2_
                                        2016_E.pdf.

Breen:1993:JED

[Bre93] J. W. Breen. A Japanese electronic dictionary project (part 1: The dictionary files). Technical Report 93-13, Monash University, Clayton, Victoria, Australia, November 30 1993. 25 pp. Available via anonymous

                                        ftp in pub/techrports/RDT/
                                        93-13.ps.Z from ftp.rdt.
                                        monash.edu.au.

Brewer:1998:KAA

[Bre98] Judy Brewer. Keynote address: Accessibility and universal design of the Web. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Brickell:1996:ELC

[Bri96] A. Brickell. The European library community and Unicode/ ISO 10646. In Unicode Consortium [Uni96c], pages C3–

??  ISBN ????  LCCN

QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Brickell:1997:UIC

[Bri97a] A. Brickell. Unicode/ISO 10646 and the Chase Project. In Segbert et al. [SSB97], pages 119–126. ISBN 92-828-0305-8. LCCN ????

Brickell:1997:CPE

[Bri97b] Anthony Brickell. The CHASE Project — experiments with Unicode in European libraries. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Brownell:1997:WWR

[Bro97]  Lori Brownell.  Writing

world ready applications for Win32. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Brown:1998:UCU

[Bro98] David C. Brown. Uniscribe — A common Unicode API for scripts on Win32. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Buchta:1996:NLUa

[Buc96a]  S. Buchta.  National lan-

guage and Unicode support in relational databases and SQL2/3. In Unicode Consortium [Uni96c], pages B1– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Buchta:1996:NLUb

[Buc96b]  S. Buchta.  National lan-

guage and Unicode support in relational databases and SQL2/3. In Unicode Consortium [Uni96c], pages B2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Buchta:1996:NLUc

[Buc96c]  S. Buchta.  National lan-

guage and Unicode support in relational databases and SQL2/3. In Unicode Consortium [Uni96b], pages B1–?? (or B2–??). ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Buchta:1997:NLS

[Buc97] S. Buchta. National language support and Unicode in relational databases and SQL. In Unicode Consortium [Uni97b], pages TC2–??, B2–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Buchta:1998:UD

[Buc98] Stefan Buchta. Unicode in databases. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Bulley:2005:UB

[Bul05] Michael Bulley. Unicode for beginners. English Today, 21(2): 63–64, ???? 2005. ISSN 0266-0784.

Bunz:1997:BMW

[Bun97] C.-M. Bunz. Browsing the memory of the world. In Unicode Consortium [Uni97b], pages A7–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Bunz:2000:ESP

[Bun00] Carl-Martin Bunz. Encoding scripts from the past: Conceptual and practical problems and solutions. Unicode Technical Note 3, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org,  Septem-

ber 2000. URL http://

www.unicode.org/notes/tn3/
bunz-iuc17pap.pdf;   http:
//www.unicode.org/notes/tn3/
tn3-1.html.

Burke:2001:UC

[Bur01] Sean M. Burke. Unidecode. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):P54, P56–P60, December 2001. CODEN SYADE7. ISSN 1061-2688. Perl Journal supplement.

Bik:1997:JPJ

[BVG97] A. J. C. Bik, J. E. Villacis, and D. B. Gannon. javar: a prototype Java restructuring compiler. Concurrency, Pract. Exp. (UK), Concurrency: Practice and Experience, 9(11):1181– 1191, November 1997. CODEN CPEXEI. ISSN 1040-3108.

Christian:2009:DOT

[C^{+}09] Andreas Christian et al. DB2 Optimization Techniques for SAP Database Migration and Unicode Conversion. IBM, International Technical Support Organization, Poughkeepsie, NY, USA, 2009. ISBN 0-7384-3308-X. xiv + 278 pp. LCCN QA76.9.D3 D43 2009.

Cain:1994:PAU

[Cai94] J. Cain. Practical applications of Unicode. In McCallum and Ertel [ME94], pages 103– 114. ISBN 3-598-21797-8. ISSN 0344-6891. LCCN Z699.5.F67 A872 1994.

Caldwell:1991:USI

[Cal91] J. T. Caldwell. Unicode: a standard international character code for multilingual information processing. In Mair and Liu [ML91], pages 180–191. ISBN 90-5199-061-8. LCCN PL1074.5 .C372 1991.

Caplan:1998:ILT

[Cap98] Barry Caplan. Internationalization and localization through the product lifecycle — an architectural approach. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Carr:1993:PBC

[Car93] Robert Carr. Pen-based computing and its impact. In USENIX Association [USE93], pages x + 530. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993. Keynote address.

Cavanaugh:1997:ULC

[Cav97a] K. Cavanaugh. Unicode and the localization challenges of corporate developers. In Unicode Consortium [Uni97b], pages C2–?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Cavanaugh:1997:LDN

[Cav97b]  Kevin Cavanaugh.  Lotus,

Domino, Notes and Java. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Chang:1997:PSF

[Cha97] P. Chang. Personalized searching and filtering for large volume of Chinese documents. In Unicode Consortium [Uni97b], pages B13–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Chandrakar:2002:MSB

[Cha02] R. Chandrakar. Multi-script bibliographic database: an Indian perspective. Online Information Review, 26(4):246–251, August 21, 2002. CODEN ???? ISSN 1468-4527.

Chabillion:2007:DLL

[Cha07] Christian Chabillion. Dossier — langues et langages – petit precis de codage des caracteres — Unicode dans le Sudoc. Bulletin des bibliothéques de France, 52 (3):28, ???? 2007. ISSN 0006-2006.

ChandraSekar:2012:CS

[Cha12] V. Chandra Sekar. Communication systems. Oxford University Press, New Delhi, India, 2012. ISBN 0-19-807805-6 (paperback). xxiii + 923 pp. LCCN TK5105.5 .C4587 2012.

Cherlin:1994:WIU

[Che94] Edward Cherlin. The Worldwide Impact of the Unicode Character Set. Character Type, 6611 Linville Drive, Weed, CA 96094, 1994. vi + 181 pp. US$995.00.

Cherlin:1997:BLB

[Che97] Edward Cherlin. Breaking the language barrier — Java, which lacks legacy code, is accelerating the pace of adoption for Unicode, the global character set that’s as fluent in Vietnamese as it is in English. Unix developer: strategies for the open systems architect, 1(1):56–59, January and February 1997. CO-DEN UNDEF5. ISSN 1090-2279. Publication of this journal ceased after the first issue.

Chen:1998:UBC

[Che98] Jack Chen. Unicode based Chinese text retrieval. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Cheng:2001:IIS

[Che01] Yan Chee Cheng. The importance of international standards on Hong Kong’s e-business infrastructure. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/ka.doc.

Chellappan:2004:TUCa

[Che04a] P. Chellappan. TAB to Unicode conversion. Unicode Technical Note 17, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, September 23, 2004.

URL http://www.unicode.
org/notes/tn17/;   http://
www.unicode.org/notes/tn17/
tab_to_unicode.pdf;   http:
//www.unicode.org/notes/tn17/
tn17-1.html.

Chellappan:2004:TUCb

[Che04b] P. Chellappan. TAM to Unicode conversion. Unicode Technical Note 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, September 23, 2004.

URL http://www.unicode.
org/notes/tn18/;   http://
www.unicode.org/notes/tn18/
tam_to_unicode.pdf;   http:
//www.unicode.org/notes/tn18/
tn18-1.html.

Chivers:2017:OUR

[Chi17]  Howard Chivers.  Optimis-

ing Unicode regular expression evaluation with previews. Software—Practice and Experience, 47(5):669–688, May 2017. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clausecker:2023:TUC

[CL23] Robert Clausecker and Daniel Lemire. Transcoding Unicode characters with AVX-512 instructions. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53(12):2430– 2462, December 2023. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Clark:1997:OSA

[Cla97] Jonathan Clark. Overcoming search ambiguities in Unicode 2.0. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Clews:1988:LAW

[Cle88] John Clews. Language automation worldwide: the development of character set standards. Number 5962 in British Library R & D reports. SESAME Computer Projects, Harrogate, North Yorkshire, UK, 1988. ISBN 1-870095-01-4. 104 pp. LCCN QA76.5 .C54 1988.

Clews:1994:IHF

[Cle94] J. Clews. Information handling in foreign languages and scripts. Journal of Document and Text Management, 2(3):

186–200, ???? 1994. CODEN JDTMEJ. ISSN 0969-9325.

Cao:1998:SHM

[CLLL98] Ling Cao, Mun-Kew Leong, Ying Lu, and Hwee-Boon Low. Searching heterogeneous multilingual bibliographic sources. Technical report, Kent Ridge Digital Labs, 21 Heng Mui Keng Terrace, Singapore 119597, Singapore, 1998. URL http:

                                        //www7.conf.au/programme/
                                        posters/1874/com1874.htm.

Clissman:1997:UPS

[CMD^{+}97] C. Clissman, R. Murray, E. Davidson, J. Hands, O. Sijtsma, A. Noordzij, R. Moulton, S. Shanawa, and J. Darzentas. The UNIverse Project: Stateof-the-art of the standards, softwares and systems which will underpin the development. part 1: Z39.50: WWW integration with Z39.50 and Unicode. New library world, 98(1138):267–274, ???? 1997. ISSN 0307-4803.

Coatsworth:1993:IIP

[Coa93] D. Coatsworth. Issues in internationalizing PHIGS applications. In First Annual PHIGS User’s Group Conference. Orlando, FL, USA, 21–24 March, 1993, pages 284–289. Rensselaer Polytechnic Institute, Troy, NY, USA, 1993.

Cobb:1998:UUJ

[Cob98] Joan Cobb. Use of Unicode and Java in a bibliographic application. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Constable:2002:TML

[Con02] Peter Constable. Toward a model for language identification and An analysis of ISO 639. Unicode Technical Note 8, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 2002. URL http://www.
unicode.org/notes/tn8/.

Correll:2000:GER

[Cor00] Sharon Correll. Graphite: An extensible rendering engine for complex writing systems. In Unicode Consortium [Uni00a], page 21. ISBN ???? LCCN ????

URL http://graphite.sil.
org/pdf/IUC17_paper.pdf.

RFC3492

[Cos03] A. Costello. RFC 3492: Punycode: a bootstring encoding of Unicode for Internationalized Domain Names in Applications (IDNA), March

2003.  URL ftp://ftp.
internic.net/rfc/rfc3492.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc3492.txt. Sta-

tus: PROPOSED STAN-DARD.

Coyle:2006:UUC

[Coy06] K. Coyle. Unicode: The universal character set part 2: Unicode in library systems. Journal of Academic Librarianship, 32(1):101–103, ???? 2006. ISSN 0099-1333.

Cozens:2017:SNT

[Coz17] Simon Cozens. SILE: A new typesetting system. TUGboat, 38(1):23–27, ???? 2017. CODEN ???? ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb38-1/tb118cozens.
pdf.

Clinton:1998:LBM

[CPJ^{+}98] Terry Clinton, Tom Parsons, Capers Jones, William Adams, Garth Klatt, Eric Haines, Ted Lewis, Philip Machanik, Stig Nilsson, Karl Reed, Howard R. Stearns, Neville Holmes, and John Brownie. Letters: The benefits of model-based integration; documentation is not green; picking on the overdog; buggy, slow windoze; there’s no such thing as free software; Linus’ law of open source development; bug-free development? No way; governmental IT planning and the Computer Society; text encoding questions; encoding the world’s languages. Computer, 31(11):4, 5–7, 11, November 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/ry004.

pdf. Two letters discuss Unicode and Multicode [Mud97].

Craft:1998:SPS

[Cra98] David J. Craft. Simple preprocessors significantly improve LZ 1 compression. In Storer and Cohn [SC98], pages 538– ?? CODEN DDCCF9. ISBN 0-8186-8406-2, 0-8186-8408-9 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1998. IEEE Computer Society Press order number PR08406. IEEE Order Plan catalog number 98TB100225.

RFC4042

[Cri05] M. Crispin. RFC 4042: UTF-9 and UTF-18 efficient transformation formats of Unicode, April 2005. URL ftp://ftp.

internic.net/rfc/rfc4042.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc4042.txt. Sta-

tus: INFORMATIONAL.

Chan:2001:MEU

[CS01] Hon-Wah Chan and Murray Sargent III. Multilingual editing using RichEdit 4+. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/b9.ppt.

Chikamatsu:1998:DJK

[CY98] Nobuko Chikamatsu and Shoichi Yokoyama. Development of a Japanese Kanji character frequency list. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Czyborra:1998:GUI

[Czy98] Roman Czyborra. Der Globalzeichensatz Unicode im Betriebssystem Unix [English: The global character set Unicode in the Unix operating system]. Diplomarbeit, Fachgebiet Informatik in Natur- und Ingenieurwissenschaften, Technische Universität Berlin, Berlin, Germany, November 30, 1998.

                                        URL http://czyborra.com/
                                        diplom/abgabe.html.  Ab-

stract in German, thesis in English.

Daniels:1992:BKE

[DAA^{+}92] Andy Daniels, Lloyd Anderson, Glenn Adams, Lee Collins, and Joe Becker. Burmese, Khmer, Ethiopian. Unicode Technical Report 1, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

                                        inc@unicode.org, November

3, 1992. URL http:/

                                        /www.unicode.org/unicode/
                                        reports/tr1.html.

Daniels:1993:UCU

[Dan93] P. T. Daniels. The Unicode Consortium: The Unicode standard. Language: journal of the Linguistic Society of America, 69(1):225, March 1993. ISSN 0097-8507.

Danish:2008:USA

[Dan08] Magda Danish. Unicode standard annex #41: Common references for Unicode standard annexes. Technical Report 41, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, March 26, 2008.

URL http://www.unicode.
org/reports/tr41/;   http:
//www.unicode.org/reports/
tr41/tr41-3.html.

Davis:1997:CGA

[Dav97a] M. Davis. Creating global applications with JDK 1.1. In Unicode Consortium [Uni97b], pages A11–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Davis:1997:WMO

[Dav97b] M. Davis. Well-mannered objects: tips on porting C++ to Java. C++ Report, 9(4):14–21, April 1997. CODEN CRPTE7. ISSN 1040-6042.

Davis:1998:BAR

[Dav98a]  Mark Davis.  The bidirec-

tional algorithm reference implementation. Unicode Technical Report 9, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org,  October

30, 1998. URL http:/

/www.unicode.org/unicode/
reports/techreports.html.

In progress.

Davis:1998:UNG

[Dav98b] Mark Davis. Unicode newline guidelines. Unicode Technical Report 13, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, April 1, 1998. URL http:

                                        //www.unicode.org/unicode/
                                        reports/techreports.html;
                                         http://www.unicode.org/unicode/
                                        reports/tr13/.

Davis:1998:UNF

[Dav98c] Mark Davis. Unicode normalization forms. Unicode Technical Report 15, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, August 20, 1998.

                                        URL http://www.unicode.
                                        org/unicode/reports/tr15/.

In progress.

Davis:1999:CM

[Dav99a] Mark Davis. Case mappings. Unicode Technical Report 21, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

                                        inc@unicode.org, November

03, 1999. URL http:/

                                        /www.unicode.org/unicode/
                                        reports/tr21/.

Davis:1999:CMT

[Dav99b] Mark Davis. Character mapping tables. Unicode Technical Report 22, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, November 23, 1999.

URL http://www.unicode.
org/unicode/reports/tr22/.

Davis:1999:URE

[Dav99c] Mark Davis. Unicode regular expression guidelines. Unicode Technical Report 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, November 23, 1999.

URL http://www.unicode.
org/unicode/reports/tr18/.

Davis:1999:U

[Dav99d] Mark Davis. UTF-32. Unicode Technical Report 19, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, November 23, 1999.

URL http://www.unicode.
org/unicode/reports/tr19/.

Davis:2000:BA

[Dav00a]  Mark Davis.  The bidirec-

tional algorithm. Unicode Standard Annex 9, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, August 31, 2000.

                                        URL http://www.unicode.
                                        org/unicode/reports/tr9/.

Davis:2000:UNG

[Dav00b] Mark Davis. Unicode newline guidelines. Unicode Standard Annex 13, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, August 31, 2000.

                                        URL http://www.unicode.
                                        org/unicode/reports/tr13/.

Davis:2001:BU

[Dav01a] Mark Davis. Bits of Unicode. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ????

                                        URL http://www.macchiato.
                                        com/.

Davis:2001:GC

[Dav01b] Mark Davis. The guts of collation. In Unicode Consortium [Uni01a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        macchiato.com/.

Davis:2001:SN

[Dav01c] Mark Davis. Script names. Unicode Technical Report 24, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, February 6, 2001.

URL http://www.unicode.
org/unicode/reports/tr24/.

Davis:2001:TB

[Dav01d] Mark Davis. Text boundaries. Proposed Draft Unicode Technical Report 29, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, March 11, 2001.

URL http://www.unicode.
org/unicode/reports/tr29/.

Davis:2001:Ua

[Dav01e] Mark Davis. UTF-32. Unicode Standard Annex 19, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, February 6, 2001.

URL http://www.unicode.
org/unicode/reports/tr19/.

Davis:2003:DS

[Dav03a] Mark Davis. Deterministic sorting. Unicode Technical Note 9, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, July 28, 2003.

URL http://www.unicode.
org/notes/tn9/.

Davis:2003:TB

[Dav03b] Mark Davis. Text boundaries.

Unicode Standard Annex 29, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, March 25,
2003.  URL http://www.
unicode.org/reports/tr29/
tr29-6.html;   http://www.
unicode.org/reports/tr29/
tr29-9.html;   http://www.
unicode.org/unicode/reports/
tr29/.

Davis:2003:URE

[Dav03c] Mark Davis. Unicode regular expressions. Proposed Update Unicode Technical Standard 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, August 13,
2003.  URL http://www.
unicode.org/reports/tr18/
tr18-8.html.

Davis:2004:GL

[Dav04a] Mark Davis. GDP by language. Unicode Technical Note 13, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, January 22,
2004.  URL http://www.
unicode.org/notes/tn13/;
http://www.unicode.org/notes/
tn13/tn13-1.html.

Davis:2004:LDM

[Dav04b]  Mark Davis.  Locale Data

Markup Language (LDML). Draft Unicode Technical Standard 35, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, April 19, 2004. URL http:

//www.unicode.org/press/press_
release-cldr.html;   http:
//www.unicode.org/reports/
tr35/.

Davis:2004:URE

[Dav04c] Mark Davis. Unicode regular expressions. Unicode Technical Standard 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, January 9, 2004.

URL http://www.unicode.
org/unicode/reports/tr18/.

Davis:2005:IPS

[Dav05a] Mark Davis. Identifier and pattern syntax. Unicode Standard Annex 31, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, March 25,
2005.  URL http://www.
unicode.org/reports/tr31/
tr31-4.html;   http://www.
unicode.org/reports/tr31/
tr31-5.html;   http://www.
unicode.org/unicode/reports/
tr31/.

Davis:2005:LDM

[Dav05b]  Mark Davis.  Locale Data

Markup Language (LDML). Unicode Technical Standard 35, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, June 2, 2005.

                                        URL http://unicode.org/
                                        reports/tr35/tr35-4.html;
                                         http://unicode.org/reports/
                                        tr35/tr35-5.html;   http:
                                        //www.unicode.org/unicode/
                                        reports/tr35/.

Davis:2015:UE

[Dav15] Mark Davis. Unicode emoji. Technical Report 51, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, June 9, 2015. URL http:

                                        //www.unicode.org/reports/
                                        tr51/.

Davis:2021:URE

[Dav21] Mark Davis. Unicode regular expressions. Unicode Technical Report 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

                                        inc@unicode.org, February

16, 2021. URL http:/

                                        /www.unicode.org/unicode/
                                        reports/tr18/tr18-22.html.

Davis:2022:URE

[Dav22] Mark Davis. Unicode regular expressions. Unicode Technical Report 18, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, February 8, 2022.

URL https://www.unicode.
org/reports/tr18/tr18-23.
html.

Daniels:1996:WWS

[DB96] Peter T. Daniels and William Bright, editors. The world’s writing systems. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-507993-0. xlv + 920 pp. LCCN P211 .W714 1996.

Davis:1990:U

[DC90] Mark Davis and Lee Collins.

Unicode. In IEEE [IEE90], pages 499–504. CODEN PICYE3. ISBN 0-87942-597-0. ISSN 0884-3627. LCCN TA 168 I19 1990. IEEE catalog number 90CH2930-6.

daCruz:2000:STG

[dC00] Frank da Cruz. Supplemental terminal graphics for Unicode. Technical report, The Kermit Project, Columbia University, New York City, NY, USA, March 31, 2000. URL

ftp://kermit.columbia.edu/
kermit/ucsterminal/ucsterminal.

txt. UTC Document L2/00-159 ISO WG2 Document N2265. Approved by ISO WG2 for ISO/IEC ISO 10646-1:2000/Amd.1 at Athens, and by Resolution M39.20 (Terminal Graphic Symbols).

Deitsch:2001:JI

[DC01] Andy Deitsch and David Czarnecki. Java Internationalization. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00019-7. xvi + 444 pp. LCCN QA76.73.J38 D45 2001. URL http:/

                                        /www.oreilly.com/catalog/
                                        javaint/.

Davis:2000:UNF

[DD00] Mark Davis and Martin Dürst. Unicode normalization forms. Unicode Standard Annex 15, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, August 31, 2000.

                                        URL http://www.unicode.
                                        org/unicode/reports/tr15/.

Deitel:2003:JHP

[DD03] Harvey M. Deitel and Paul J. Deitel. Java: how to program. How to program series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 2003. ISBN 0-13-101621-0, 0-13-101623-7 (CD-ROM), 0-13-120236-7. li + 1447 pp. LCCN QA76.73.J38 D45 2003.

Deitel:2007:JHP

[DD07] Paul J. Deitel and Paul J. Deitel. Java: how to program. How to program series. Pearson Prentice Hall, Upper Saddle River, NJ, USA, seventh edition, 2007. ISBN 0-13-222220-5. xliv + 1596 pp. LCCN QA76.73.J38 D45 2007.

Deitel:2011:VCH

[DD11] Paul J. Deitel and Harvey M.

Deitel. Visual C# 2010: how to program. How to program. Pearson Prentice Hall, Upper Saddle River, NJ, USA, fourth edition, 2011. ISBN 0-13-215142-1 (paperback). xxx + 959 pp. LCCN QA76.73.C154 D46 2011.

Do:1996:VIT

[DDG96] James Do, Phan Dinh Dieu, and Seymour E. Goodman. Vietnam: Information technology for the transition. Computer, 29(3):88–94, March 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Mentions ISO/IEC 10646 and the Unicode support for Vietnamese.

Durst:1999:UXO

[DDHF99] Martin Dürst, Mark Davis, Hideki Hiura, and Asmus Freytag. Unicode in XML and other markup languages. Unicode Technical Report 20, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, September 28, 1999.

URL http://pages.infinit.
net/hapax/RTU-20-Unicode-
et-XML.html;   http://www.
unicode.org/unicode/reports/
tr20/.

Davis:2016:PDU

[DE16] Mark Davis and Peter Edberg. Proposed draft: Unicode emoji mechanisms. Technical Report 52, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, May 25, 2016. ???? pp. This draft was later withdrawn, and never reached official report status.

Davis:2001:Ub

[DEFM01] Mark Davis, Michael Everson, Asmus Freytag, and Members of the Editorial Committee. Unicode 3.1. Draft Unicode Technical Report 27, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, February

23, 2001. URL http:/

/www.unicode.org/unicode/
reports/tr27/.

Demchenko:1998:TMS

[Dem98] Yuri Demchenko. Testing multilingual support in mail user agents. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Desgraupes:2005:PPU

[Des05] Bernard Desgraupes. Passeport pour Unicode. Vuibert informatique, Paris, France, 2005. ISBN 2-7117-4827-8. xix + 264 pp. LCCN ????

Desrichard:2008:PAU

[Des08] Yves Desrichard. Patrick Andries, Unicode 5.0 en pratique. Bulletin des bibliothéques de France, 53(6):108, ???? 2008. ISSN 0006-2006.

Durst:2002:UXO

[DF02] Martin J. Dürst and Asmus Freytag. Unicode in XML and other markup languages. Unicode Technical Report 20, World Wide Web Consortium (W3C), Laboratory for Computer Science at MIT, Cambridge, MA, USA, February 18, 2002.

URL http://www.unicode.
org/unicode/reports/tr20/
;   http://www.w3.org/TR/

unicode-xml/. Also available as W3C Note.

Davis:2005:SN

[DF05] Mark Davis and Asmus Freytag. Script names. Unicode Standard Annex 24, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, March 28,

                                        2005.  URL http://www.
                                        unicode.org/reports/tr24/
                                        tr24-5.html;   http://www.
                                        unicode.org/reports/tr24/
                                        tr24-7.html;   http://www.
                                        unicode.org/unicode/reports/
                                        tr24/.

Davis:2006:PUUa

[DF06] Mark Davis and Asmus Freytag. Proposed update Unicode Standard annex #24: Script names. Technical report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

                                        inc@unicode.org, January 18,
                                        2006.  URL http://www.
                                        unicode.org/reports/tr24/
                                        tr24-8.html.

Davis:1995:ISA

[DG95] M. Davis and J. Grimes. International support in applications and systems software. Objects in Europe, 2(3):10, 12, 14, May– June 1995. CODEN OBEUE9.

Davis:1996:CGS

[DGK96] M. E. Davis, J. D. Grimes, and D. J. Knoles. Creating global software: text handling and localization in Taligent’s CommonPoint application system. IBM Systems Journal, 35(2): 227–243, ???? 1996. CODEN IBMSA7. ISSN 0018-8670.

Dickson:2022:DLG

[Dic22]  Courtney Dickson.  Dakelh

language to get standardized writing system for keyboard use: Designer hopes accurate syllabics for use on phones, tablets and computers will help revitalize language. CBC Radio Web site., March 23, 2022. URL https:

//www.cbc.ca/news/canada/
british-columbia/dakelh-indigenous-
language-standard-syllabics-
1.6392552.

Downey:2019:TJ

[DL19] Allen B. Downey and Ben Lauwens. Think Julia. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4920-4500-4 (e-book), 1-4920-4503-9. LCCN QA76.73.J85 L38 2019.

Davis:2022:ASC

[DLCS22] Mark Davis, Robin Leroy, Peter Constable, and Markus Scherer. Avoiding source code spoofing. Report L2/22-007R2, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org,  January

25, 2022. URL https://

www.unicode.org/L2/L2022/
22007r2-avoiding-spoof.pdf.

Driscoll:1995:IDS

[DLLT95] Kevin Driscoll, Tuoc V. Luong, James S. H. Lok, and David J. Taylor. Internationalization: Developing Software For Global Markets, Paperback. Wiley, New York, NY, USA, June 1995. ISBN 0-471-07661-9. vii + 293 pp. LCCN QA76.76.D47 I599 1995.

Do:1997:LEU

[DM97] James Do and Muhammed Mudawwar. Letters to the editor: Unicode misunderstood. Computer, 30(6):6, 9, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response and rebuttal to [Mud97].

delGaldo:1996:IUI

[dN96] Elisa M. del Galdo and Jakob Nielsen, editors. International User Interfaces. Wiley, New York, NY, USA, June 1996. ISBN 0-471-14965-9. x + 276 pp. LCCN QA76.9.U83I57 1996. Includes articles on Arabic Langbox and Chinese Word Splitting.

Duerst:1996:WMW

[DNYA96a] M. Duerst, G. Nicol, F. Yeargeau, and G. Adams. Weaving the multilingual Web: Standards and their implementation. In Unicode Consortium [Uni96b], pages A2–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Durst:1996:WMW

[DNYA96b] M. Dürst, G. Nicol, F. Yeargeau, and G. Adams. Weaving the multilingual Web: Standards and their implementation. In Unicode Consortium [Uni96c], pages A2–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Davis:2000:TUT

[DO00] M. W. Davis and W. C. Ogden. Towards universal text retrieval: Tipster text retrieval research at New Mexico State University. Information Retrieval, 3(4):339–356, December 2000. CODEN IFRTFY. ISSN 1386-4564 (print), 1573-7659 (electronic).

Doman:1996:TIS

[Dom96] J. Doman. Testing international software. In Unicode Consortium [Uni96c], pages C7– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Diringer:1968:AKH

[DR68] David Diringer and Reinhold Regensburger. The alphabet: a key to the history of mankind.

Hutchinson, London, UK, 1968. ISBN 0-09-067640-8. xxi + 473 (vol. 1), 6 + 452 (vol. 2) pp. LCCN P211 .D53 1968. Two volumes.

DrInternational:2003:DIS

[Dr.03]  Dr. International.  Devel-

oping International Software. Microsoft Press, Bellevue, WA, USA, second edition, 2003. ISBN 0-7356-1583-7. xxxviii + 1060 pp. LCCN QA76.76.D47D479 2003. Includes CD-ROM.

Davis:2001:BBO

[DS01] Mark Davis and Markus Scherer. BOCU: Binaryordered compression for Unicode. Technical report, IBM Corporation, May 30,

2001.  URL http://oss.
software.ibm.com/icu/docs/
papers/binary_ordered_compression_
for_unicode.html.

Davis:2005:CMM

[DS05a] Mark Davis and Markus Scherer. Character mapping markup language (CharMapML). Unicode Technical Standard 22, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, June 4,
2005.  URL http://www.
unicode.org/reports/tr22/
tr22-3.html;   http://www.
unicode.org/reports/tr22/
tr22-5.html;   http://www.

unicode.org/unicode/reports/
tr22/.

Davis:2005:USC

[DS05b] Mark Davis and Michel Suignard. Unicode security considerations. Unicode Technical Report 36, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, July 7, 2005. URL http:

//www.unicode.org/unicode/
reports/tr36/.

Davis:2005:USMa

[DS05c] Mark Davis and Michel Suignard. Unicode security mechanisms. Unicode Technical Report 39, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, July 8, 2005. URL http:

//www.unicode.org/unicode/
reports/tr39/.

Davis:2005:USMb

[DS05d] Mark Davis and Michel Suignard. Unicode security mechanisms. Proposed Draft Unicode Technical Standard 39, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, July 8, 2005. URL http:

//www.unicode.org/unicode/
reports/tr39/.

Davis:2012:UIC

[DS12] Mark Davis and Michel Suignard. Unicode IDNA compatibility processing. Technical Standard 46, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, September 12,

2012.  URL http://www.
unicode.org/reports/tr46/.

Duddi:1997:NDF

[Dud97] S. Duddi. The Netscape dynamic font architecture. In Unicode Consortium [Uni97b], pages B7–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Duerst:1996:SID

[Due96] M. Duerst. Simple implementation of difficult scripts. In Unicode Consortium [Uni96b], pages B6–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Duerst:1997:III

[Due97a] M. J. Duerst. Internationalizing Internet identifiers. In Unicode Consortium [Uni97b], pages A4–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Duerst:1997:PPU

[Due97b] M. J. Duerst. The properties and promises of UTF-8. In Unicode Consortium [Uni97b], pages C14–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Durst:1994:UPC

[Dür94] Martin J. Dürst. Uniprep:

preparing a C/C++ compiler for Unicode. ACM SIGPLAN Notices, 29(1):53, January 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Durst:1996:SID

[Dür96] M. Dürst. Simple implementation of difficult scripts. In Unicode Consortium [Uni96c], pages B6–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Durst:1997:III

[Dür97a] M. J. Dürst. Internationalizing Internet identifiers. In Unicode Consortium [Uni97b], pages A4–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Durst:1997:PPU

[Dür97b] M. J. Dürst. The properties and promises of UTF-8. In Unicode Consortium [Uni97b], pages C14–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Durst:1997:UIS

[Dür97c] Martin J. Dürst. Unicode and the Internet: Status and developments. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc10/program.html.

Durst:1998:CMW

[Dür98a] Martin J. Dürst. A character model for the WWW: Purpose and status. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc13/program.html.

Durst:1998:EVN

[Dür98b] Martin J. Dürst. Equivalence vs. normalization: The case of identifiers. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Durst:1998:EPW

[Dür98c] Martin J. Dürst. Exploring the potentials of Web technologies for the handling of rare ideographs and ideograph variants. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Durst:1998:WMW

[Dür98d] Martin J. Dürst. Weaving the multilingual Web: Standards and their implementations. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Davis:1988:UCA

[DW88] Mark Davis and Ken Whistler.

Unicode collation algorithm. Unicode Technical Standard 10, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, March 30, 1988.

URL http://www.unicode.
org/unicode/reports/techreports.
html;   http://www.unicode.
org/unicode/reports/tr10/

index.html. In progress.

Davis:2002:MSS

[DW02] Mark Davis and Vladimir Weinstein. Migrating software to supplementary characters. Unicode Technical Note 7, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 2002. URL http://www.
unicode.org/notes/tn7/.

Davis:2006:PUUb

[DW06] Mark Davis and Ken Whistler. Proposed update Unicode technical standard #10: Unicode collation algorithm. Technical report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, May 7, 2006. URL http:

//www.unicode.org/reports/
tr10/;   http://www.unicode.
org/reports/tr10/tr10-15.

html. Version 5.0, revision 15.

Davis:2008:USA

[DW08] Mark Davis and Ken Whistler. Unicode standard annex #44: Unicode character database. Technical Report 44, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, August 27,
2008.  URL http://www.
unicode.org/reports/tn44/

tr44-3.html;   http://www.
unicode.org/reports/tr44/.

Davis:2012:UCD

[DW12] Mark Davis and Ken Whistler.

Unicode character database. Unicode Standard Annex 44, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, September
11, 2012. URL http://www.
unicode.org/reports/tr44/.

Durst:2002:CMW

[DYI^{+}02] Martin J. Dürst, François Yergeau, Richard Ishida, Misha Wolf, Asmus Freytag, and Tex Texin. Character model for the World Wide Web 1.0. W3c working draft, World Wide Web Consortium (W3C), Laboratory for Computer Science at MIT, Cambridge, MA, USA, February 20, 2002. URL

http://www.w3.org/TR/2002/
WD-charmod-20020220/.

Durst:2004:CMWa

[DYI^{+}04a] Martin J. Dürst, François Yergeau, Richard Ishida, Misha Wolf, Tex Texin, and Addison Phillips. Character model for the World Wide Web 1.0: Fundamentals. World-Wide Web document., February 25, 2004.

URL http://www.w3.org/TR/
2004/WD-charmod-20040225/.

W3C Working Draft.

Durst:2004:CMWb

[DYI^{+}04b] Martin J. Dürst, François Yergeau, Richard Ishida, Misha Wolf, Tex Texin, and Addison Phillips. Character model for the World Wide Web 1.0: Normalization. World-Wide Web document., February 25,

                                        2004.  URL http://www.
                                        w3.org/TR/2004/WD-charmod-
                                        norm-20040225/.

Duerst:1997:WMW

[DYN97a] M. J. Duerst, F. Yergeau, and G. T. Nicol. Weaving the multilingual Web: Standards and their implementation. In Unicode Consortium [Uni97b], pages TA3–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Durst:1997:WMWb

[DYN97b] M. J. Dürst, F. Yergeau, and G. T. Nicol. Weaving the multilingual Web: Standards and their implementation. In Unicode Consortium [Uni97b], pages TA3–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Durst:1997:WMWa

[DYN97c] Martin J. Dürst, François Yergeau, and Gavin Thomas Nicol. Weaving the multilingual Web: Standards and their implementation. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Gavin:1998:UHS

[DYN98] Martin J. Dürst, François Yergeau, and Gavin Thomas Nicol. Unicode and HTML: Status and developments. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Eckel:1997:ISJ

[Eck97] B. Eckel. IO streams in Java 1.1. Web Techniques, 2(7):30, 32–36, July 1997. CODEN WETEFA. ISSN 1086-556X.

Edberg:1997:SCE

[Edb97a] P. Edberg. Survey of character encodings. In Unicode Consortium [Uni97b], pages TC4–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Edberg:1997:SMC

[Edb97b] Peter Edberg. Supporting multiple character sets on the Internet. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Edberg:1998:SCE

[Edb98] Peter K. Edberg. Solutions for complex encoding conversion problems. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc13/program.html.

Eevee:2015:DCU

[Eev15] Eevee. Dark corners of Unicode. Web site, September 12, 2015. URL https:

                                        //eev.ee/blog/2015/09/12/
                                        dark-corners-of-unicode/.

Edberg:1998:CCS

[EHJ98] Peter K. Edberg, Edwin Hart, and John H. Jenkins. Characters, character sets and glyphs. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Eijkhout:2008:CE

[Eij08]  Victor Eijkhout.  Charac-

ter encoding. TUGboat, 29 (2):270–277, 2008. ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb29-2/
                                        tb92eijkhout.pdf.

Emerson:2001:HGC

[Eme01a] Thomas Emerson. The hitchhiker’s guide to Chinese encodings. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

                                        www.unicode.org/iuc/iuc18/
                                        papers/ta1.pdf.

Emerson:2001:ISP

[Eme01b] Thomas Emerson. Issues and solution in pan-China information retrieval. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/b15.pdf.

Emly:2006:MED

[Eml06] Michael Emly. Michael Emly describes Unicode for nonroman scripts. Catalogue & index: periodical of the Library Association Cataloguing and Indexing Group, 154:7, ???? 2006. ISSN 0008-7629.

UC:1993:E

[Enc93] Encoding, 1993. ISSN 1076-1918. The Unicode Consortium, Mountain View, CA, USA. Quarterly.

Erickson:1997:OPM

[ES97] Janet C. Erickson and Abazar Sepehri. Options for presentation of multilingual text: use of the Unicode standard. Library Hi Tech, 15(3–4):172–187 (or 172–188??), March 1, 1997. CODEN LIHTD2. ISSN 0737-8831.

Esselink:1998:PGS

[Ess98] Bert Esselink. A practical guide to software localization: for translators, engineers and project managers, volume 3 of The language international world dictionary. John Benjamins Pub. Co., Amsterdam, The Netherlands and Philadelphia, PA, USA, 1998. ISBN 90-272-1953-2 (hardback), 90-272-1954-0 (paperback), 1-55619-742-X (hardback), 1-55619-743-8 (paperback). 309 pp. LCCN QA76.76.D63 E87 1998.

Esselink:2000:PGL

[Ess00] Bert Esselink. A practical guide to localization, volume 4 of Language international world directory. John Benjamins Pub. Co., Amsterdam, The Netherlands and Philadelphia, PA, USA, 2000. ISBN 1-58811-006-0. 488 pp. LCCN QA76.76.D63 E87 2000. US$32.95.

Etemad:2005:RVT

[Ete05] Elika J. Etemad. Robust vertical text layout. Unicode Technical Note 22, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 25, 2005.

URL http://www.unicode.
org/notes/tn22/;   http://
www.unicode.org/notes/tn22/
RobustVerticalLayout.pdf;
 http://www.unicode.org/notes/
tn22/tn22-1.html.

Everson:2002:LUP

[Eve02] Michael Everson. Leaks in the Unicode pipeline: Script, script, script, . . .. Unicode Technical Note 4, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, May 2002.
URL http://www.unicode.
org/notes/tn4/;   http://
www.unicode.org/notes/tn4/
everson-iuc21pap.pdf. Pre-

sented at the 21st International Unicode Conference in May 2002.

Ewell:2004:SUCa

[Ewe04a] Doug Ewell. A survey of Unicode compression. Technical report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, January 1, 2004. URL

http://users.adelphia.net/
~dewell/compression.html;
 http://users.adelphia.net/
~dewell/UnicodeCompression.
pdf.

Ewell:2004:SUCb

[Ewe04b] Doug Ewell. A survey of Unicode compression. Unicode Technical Note 14, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, January 30,
2004.  URL http://www.
unicode.org/notes/tn14/;
http://www.unicode.org/notes/
tn14/tn14-1.html;   http:/
/www.unicode.org/notes/tn14/
UnicodeCompression.pdf.

Fairbairns:1995:OWB

[Fai95]  Robin Fairbairns.  Omega

— why bother with Unicode? TUGboat, 16(3):325– 328, September 1995. ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb16-3/
                                        tb48fair.pdf.

Fenwick:1998:CUF

[FB98] Peter Fenwick and Simon Brierley. Compression of Unicode files. In Storer and Cohn [SC98], pages 547–?? CODEN DD-CCF9. ISBN 0-8186-8406-2, 0-8186-8408-9 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1998. IEEE Computer Society Press order number PR08406. IEEE Order Plan catalog number 98TB100225.

Fieschi:2004:MPW

[FCJ04] Marius Fieschi, Enrico Coiera, and Yu-Chan Jack, editors. Medinfo 2004: Proceedings of the 11th World Congress on Medical Informatics, San Francisco, September 7–11, 2004, volume 107 of Studies in health technology and informatics. IOS Press, Amsterdam, The Netherlands, 2004. ISBN 1-58603-444-8. Two volumes.

Fu:2006:RTU

[FDW06] A. Y. Fu, X. Deng, and L. Wenyin. REGAP: a tool for Unicode-based Webg identity fraud detection. Journal of Digital Forensic Practice, 1(2): 83–97, ???? 2006. ISSN 1556-7281. Inaugural research oriented conference; Special issue: phishing and online fraud, part 1.

Feibus:1994:SPL

[Fei94] Andy Feibus. Standards in progress: a look at the 16-bit Unicode character set. Open systems today: the newspaper of Unix and interoperable computing, 25(4):28–31, April 1994. ISSN 1061-0839.

Fennessy:1996:PBM

[Fen96] Quentin Fennessy. Plan 9: Bad movie, good operating system. Computer, 29(5):117–119, 121– 123, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Mentions Plan 9’s support for Unicode.

Firmage:1993:AAS

[Fir93]  Richard A. Firmage.  The

Alphabet Abecedarium: Some Notes on Letters. David R. Godine, Publisher, Boston, MA, USA, 1993. ISBN 0-87923-987-5. xi + 307 pp. LCCN P211 .F6 1993.

Flynn:1998:UDB

[Fly98] Randall E. Flynn. Unicode in data bases. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Freytag:2006:KAU

[FMW06] Asmus Freytag, Rick Mc-Gowan, and Ken Whistler. Known anomalies in Unicode character names. Unicode Technical Note 27, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

                                        unicode.org, May 8, 2006.
                                        URL http://www.unicode.
                                        org/notes/tn27/;   http://
                                        www.unicode.org/notes/tn27/
                                        tn27-1.html.

Forsander:1996:UTE

[For96] C. Forsander. Unicode text editor applet. In Unicode Consortium [Uni96c], pages B5– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Fowles:1996:IUY

[Fow96] K. Fowles. Implementing Unicode in your Internet applications. In Unicode Consortium [Uni96c], pages A2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Fowles:1997:GSU

[Fow97] Ken Fowles. Globalization software: Unicode evolves: Worried about localization problems? Content-enabled software, through Unicode, may be what really matters. Byte Magazine, 22(3):105, 106, 108, 110, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Freed:1998:RIC

[FP98] N. Freed and J. Postel. RFC 2278: IANA charset registration procedures, January 1998.

URL ftp://ftp.internic.
net/rfc/bcp19.txt;   ftp:/
/ftp.internic.net/rfc/rfc2278.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/bcp19.txt;   ftp:
//ftp.math.utah.edu/pub/rfc/
rfc2278.txt.

Freytag:1988:LBP

[Fre88a] Asmus Freytag. Line breaking properties. Unicode Technical Report 14, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, May 19, 1988.

URL http://www.unicode.
org/unicode/reports/dtr14.
html;   http://www.unicode.
org/unicode/reports/techreports.

html. In progress.

Freytag:1988:UCP

[Fre88b] Asmus Freytag. Unicode character property “East Asian width”. Unicode Technical Report 11, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, May 19, 1988.

URL http://www.unicode.
org/unicode/reports/dtr11.
html;   http://www.unicode.
org/unicode/reports/techreports.

html. In progress.

Freytag:1996:JIA

[Fre96a]  A. Freytag.  Java interna-

tionalization architecture. In Unicode Consortium [Uni96c], pages A3–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Freytag:1996:USV

[Fre96b] A. Freytag. The Unicode standard: Version 2. In Unicode Consortium [Uni96c], pages B2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Freytag:1997:RTM

[Fre97a] A. Freytag. Report from the trenches: Microsoft Publisher goes Unicode. In Unicode Consortium [Uni97b], pages C7–

??  ISBN ????  LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Freytag:1997:SCS

[Fre97b] A. Freytag. A standard compression scheme for Unicode. In Unicode Consortium [Uni97b], pages C15–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Freytag:1997:UST

[Fre97c] A. Freytag. The Unicode standard: a tutorial. In Unicode Consortium [Uni97b], pages TA1–?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Freytag:1997:JIAa

[Fre97d] Asmus Freytag. Java internationalization architecture. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Freytag:1997:JIAb

[Fre97e] Asmus Freytag. Java internationalization architecture — continued. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Freytag:1997:USV

[Fre97f] Asmus Freytag. The Unicode Standard: Version 2. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Freytag:1998:BL

[Fre98a]  Asmus Freytag.  Breaking

lines. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Freytag:1998:IU

[Fre98b] Asmus Freytag. Introduction to Unicode 2.1. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Freytag:2000:EAW

[Fre00a] Asmus Freytag. East Asian width. Unicode Standard Annex 11, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, August 31, 2000. URL

http://www.unicode.org/unicode/
reports/tr11/.

Freytag:2000:LBP

[Fre00b] Asmus Freytag. Line breaking properties. Unicode Standard Annex 14, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, August 31, 2000.

URL http://www.unicode.
org/unicode/reports/tr14/.

Freytag:2002:CF

[Fre02a]  Asmus Freytag.  Charac-

ter foldings. Unicode Technical Report 30, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, May 8, 2002. URL http:

//www.unicode.org/reports/
tr30/.

Freytag:2002:UCP

[Fre02b] Asmus Freytag. The Unicode character property model. Unicode Technical Report 23, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 23, 2002.

URL http://www.unicode.
org/reports/tr23/;   http:
//www.unicode.org/reports/
tr23/tr23-4.html;   http:
//www.unicode.org/reports/
tr23/tr23-5.html.

Freytag:2006:UPO

[Fre06] Asmus Freytag. The use of phonetic and other symbols in dictionaries: A brief survey. Unicode Technical Note 29, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 12, 2006.

URL http://www.unicode.
org/notes/tn29/;   http://
www.unicode.org/notes/tn29/
tn29-1.html.

Fitzpatrick:1997:PBS

[FV97] Michael E. Fitzpatrick and Laurence Vanhelsuwé. Programmer’s bookshelf: The software project and Unicode. Dr. Dobbs Journal, 22(8):125, 127, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Gosling:19xx:JLS

[G^{+}xx] James Gosling et al. The Java Language Specification. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905, Metro Plaza Tower 2, No. 223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? URL http:

                                        //java.sun.com/people/jag/
                                        ;   http://www.gotop.com.

tw. Chinese translation by Thi Shiang Workshop.

Garfinkel:2012:PU

[Gar12] Simson L. Garfinkel. Programming Unicode. ;login: the USENIX Association newsletter, 37(2):25– 37, April 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2012/programming-unicode.

Gavrishin:1996:IMI

[Gav96] I. Gavrishin. Input of multilingual information in Unicodebased applications. In Unicode Consortium [Uni96c], pages C4– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Gavrishin:1997:MBN

[Gav97] Ilya Gavrishin. Multilingual bulletin of new receipts of the Russian Library for Foreign Literature now on the Net. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Gerovac:1996:SSH

[GC96] Branko J. Gerovac and David C.

Carver. Standards: Standardizing headers in the quest for interoperability. IEEE MultiMedia, 3(1):84–88, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1996/pdf/u1084.

pdf. Comments of the use of Unicode for multimedia data storage.

Goldsmith:1994:RUU

[GD94a] D. Goldsmith and M. Davis.

RFC 1641: Using Unicode with MIME, July 1994. URL ftp://

ftp.internic.net/rfc/rfc1641.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc1641.txt. Sta-

tus: EXPERIMENTAL.

Goldsmith:1994:RUM

[GD94b] D. Goldsmith and M. Davis.

RFC 1642: UTF-7 — A mail-safe transformation format of Unicode, July 1994.

URL ftp://ftp.internic.
net/rfc/rfc1642.txt;   ftp:
//ftp.internic.net/rfc/rfc2152.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc1642.txt;
ftp://ftp.math.utah.edu/pub/
rfc/rfc2152.txt. Obsoleted

by RFC2152 [GD97]. Status: EXPERIMENTAL.

Goldsmith:1997:RUM

[GD97] D. Goldsmith and M. Davis. RFC 2152: UTF-7: a mail-safe transformation format of Unicode, May 1997. URL ftp://

                                        ftp.internic.net/rfc/rfc1642.
                                        txt;   ftp://ftp.internic.
                                        net/rfc/rfc2152.txt;   ftp:
                                        //ftp.math.utah.edu/pub/rfc/
                                        rfc1642.txt;   ftp://ftp.
                                        math.utah.edu/pub/rfc/rfc2152.

txt. Obsoletes RFC1642 [GD94b]. Status: INFORMA-TIONAL.

Gidali:2001:ALB

[Gid01] Israel Ervin Gidali. Alphabets of languages with bidirectional scripts and aspects of these scripts support. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ????

                                        URL http://www.unicode.
                                        org/iuc/iuc18/papers/c10.
                                        ppt.

Gillig:1994:ESP

[Gil94] James R. Gillig. Endian-neutral software, part 2. Dr. Dobbs Journal, 19(13):44–51, November 1994. CODEN DDJOEB. ISSN 1044-789X.

Gilbert:1997:MTP

[Gil97] Judith Gilbert. Multilingual template pages for the WWW based on Unicode. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ????

                                        URL http://www.unicode.

org/unicode/iuc10/program.
html.

Gillam:1998:TBA

[Gil98a] Richard Gillam. Text boundary analysis in JDK 1.2. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Gillam:1998:RBA

[Gil98b] Richard T. Gillam. A rulebased approach to number spellout. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Gillam:2003:UDP

[Gil03] Richard Gillam. Unicode demystified: a practical programmer’s guide to the encoding standard. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-70052-2. xxxiii + 853 pp. LCCN QA76.6 .G5535 2002. UK£37.99.

Gosling:1996:JLS

[GJS96] James Gosling, Bill Joy, and Guy L. Steele. The Java Language Specification. The Java Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63451-1. xxv + 825 pp. LCCN QA76.73.J38G68 1996. US$36.75. URL http://www.

aw.com/cp/javaseries.html.

Gosling:2000:JLS

[GJSB00] James Gosling, Bill Joy, Guy L. Steele, and Gilad Bracha. The Java Language Specification. Java series. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-31008-2. xxv + 505 pp. LCCN QA76.73.J38 G68 2000.

                                        URL http://java.sun.com/
                                        people/jag/.

Girard:1998:UBCa

[GN98a] Joann Girard and R. Douglas Nordeman. Unicode on a budget: a case study for embedded products. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc12/program.html.

Girard:1998:UBCb

[GN98b] Joann K. Girard and R. Douglas Nordeman. Unicode on a budget: a case study for embedded products. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc13/program.html.

Goldsmith:1997:SMC

[Gol97a] D. Goldsmith. Supporting multiple character sets on the Internet. In Unicode Consortium [Uni97b], pages C12–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/program.html.

Goldstein:1997:NTG

[Gol97b] D. Goldstein. New tools for globalizing software. In Unicode Consortium [Uni97b], pages C3–?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Goldstein:1997:SGS

[Gol97c] David Goldstein. Strategies for globalizing software. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Goldsmith:1998:HCU

[Gol98a] David Goldsmith. Handling complex Unicode rendering problems. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Goldsmith:1998:SMC

[Gol98b] David Goldsmith. Supporting multiple character sets on the Internet. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Goldsmith:1998:SCE

[Gol98c]  David Goldsmith.  Survey

of character encodings. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Gong:1998:UD

[Gon98] Yu Gong. Unicode in databases. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Gray:1996:EDM

[Gra96] Jim Gray. Evolution of data management. Computer, 29 (10):38–46, October 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Laments lack of Unicode support in SQL for databases.

Graham:1999:UWI

[Gra99] T. Graham. Unicode^{T M}: What is it and how do I use it? Markup languages: theory & practice, 1(4):75–102, Fall 1999. CODEN MLTPFG. ISSN 1099-6621.

Graham:2000:UP

[Gra00] Tony Graham. Unicode: a primer. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2000. ISBN 0-7645-4625-2. xxvi + 475 pp. LCCN QA268.G72 2000.

Greenwood:1998:MICa

[Gre98a]  Tim Greenwood.  Multilin-

gual Internet commerce: a case study. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Greenwood:1998:MICb

[Gre98b]  Tim Greenwood.  Multilin-

gual Internet commerce: a case study. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Griffith:1998:NDF

[Gri98a] Tague Griffith. The Netscape dynamic font architecture. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Griffith:1998:UISa

[Gri98b] Tague Griffith. Unicode and internationalization support on Netscape ONE. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc12/program.
html.

Griffith:1998:UISb

[Gri98c] Tague Griffith. Unicode and internationalization support on Netscape ONE. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc13/program.
html.

Goossens:1992:STP

[GV92] M. Goossens and E. Van Herwijnen. Scientific text processing. International Journal of Modern Physics C [Physics and Computers], 3(3):479–546, June 1992. CODEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic). Describes progress in standardization efforts, including that of Unicode.

Gwyn:1993:CSU

[Gwy93] Douglas A. Gwyn. C Standards update. The Journal of C Language Translation, 5(1): ??, September 1993. ISSN 1042-5721.

Houlette:1995:IGW

[H^{+}95] Forrest Houlette et al. Insider’s guide to Windows 95 programming. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56529-679-6. 678 pp. LCCN QA76.76.O63 I566 1995. US$34.99, CDN$46.99.

Hadley:2001:CSB

[Had01] Joshua Hadley. Case studies in building megafonts. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        agfamonotype.com/presentations/

.

Hagen:2007:CMI

[Hag07] Hans Hagen. ConTEXt MkIV: Going UTF. TUGboat, 28(3): 314–316, 2007. ISSN 0896-

                                        3207. URL https://tug.org/

TUGboat/tb28-3/tb90hagen.
pdf.

Hagen:2011:HMR

[Hag11] Hans Hagen. Handling math:

a retrospective. TUGboat, 32 (1):43–46, 2011. ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb32-1/tb100hagen-
math.pdf.

Hagen:2013:PBD

[Hag13]  Nathan Hagen.  Bibulous

— a drop-in BibTEX replacement based on style templates. TUGboat, 34(3):332–339, 2013. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb34-3/
tb108hagen.pdf.

Hagen:2016:CEU

[Hag16] Hans Hagen. Colorful emojis via Unicode and OpenType. TUGboat, 37(3):305, 2016. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb37-3/
tb117hagen-emoji.pdf.

Hahn:1997:EAC

[Hah97] Thomas H. Hahn. East Asian character coding and new software implementations. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Hall:1994:IWNa

[Hal94a] William S. Hall. Internationalization in Windows NT, part I: Programming with Unicode. Microsoft Systems Journal, 9 (6):57–72, June 1994. CODEN MSJOED. ISSN 0889-9932.

Hall:1994:IWNb

[Hal94b] William S. Hall. Internationalization in Windows NT, part II: Locales, languages, and resources. Microsoft Systems Journal, 9(7):55–76, July 1994. CODEN MSJOED. ISSN 0889-9932.

Hall:1996:UW

[Hal96a] B. Hall. Unicode and Win32. In Unicode Consortium [Uni96b], pages C1–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Hall:1996:UWP

[Hal96b] B. Hall. Unicode and Win32. In Unicode Consortium [Uni96c], pages C1–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Hall:1997:IUC

[Hal97a] B. Hall. An introduction to Unicode and code pages in Win32. In Unicode Consortium [Uni97b], pages TC3–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Hall:1997:WU

[Hal97b] William S. Hall. Win32 and Unicode. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Hall:1998:UWN

[Hal98a] Bill Hall. Unicode, Windows NT, and Windows 95. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Hall:1998:UWIa

[Hal98b] William S. Hall. Unicode and Win32 internationalization. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Hall:1998:UWIb

[Hal98c] William S. Hall. Unicode and Win32 internationalization — continued. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Hamilton:1996:JSN

[Ham96] Marc A. Hamilton. Java and the shift to net-centric computing. Computer, 29(8):31– 39, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Mentions Java’s use of Unicode characters.

Haralambous:1995:TAT

[Har95] Yannis Haralambous. The traditional Arabic typecase extended to the Unicode set of glyphs. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 111–123, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Haralambous:1997:TAT

[Har97a] Yannis Haralambous. The traditional Arabic typecase, Unicode, TEX and . TUGboat, 18(1):17–29, March 1997. ISSN 0896-3207.

Hart:1997:WYN

[Har97b] E. Hart. What you need to know about processing and rendering multilingual text. In Unicode Consortium [Uni97b], pages B10–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Haralambous:1999:UTC

[Har99] Yannis Haralambous. From Unicode to typography, a case study: the Greek script. In Unicode Consortium, editor, Proceedings of the 11th Unicode Conference, Boston, 1999, page ???? The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, 1999. ISBN

????  LCCN ????  URL

http://genepi.louis-jean.
com/omega/boston99.pdf.

Haralambous:2004:FCG

[Har04] Yannis Haralambous. Fontes & codages: Glyphes et caractères à l’ère du numérique. O’Reilly, Paris, France, 2004. ISBN 2-84177-273-X. xx + 990 pp. LCCN ???? URL http:

//www.oreilly.com/catalog/
9782841772735.

Harai:2007:UUJ

[Har07a] Naoko Harai. UNIMARC and UNICODE — Japanese scripts and UNIMARC. International cataloguing and bibliographic control: quarterly bulletin of the IFLA UBCIM Programme, 36(3):55, 2007. ISSN 1011-8829.

Hardie:2007:LEU

[Har07b] A. Hardie. From legacy encodings to Unicode: the graphical and logical principles in the scripts of South Asia. Language Resources and Evaluation, 41 (1):1–25, ???? 2007. ISSN 1574-020X.

Haralambous:2014:SAT

[Har14] Yannis Haralambous. A simple Arabic typesetting system for mixed Latin/Arabic documents: d.ād. TUGboat, 35(3):277–283, 2014. ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb35-3/
                                        tb111haralambous.pdf.

Heller:1992:AWP

[Hel92] Martin Heller. Advanced Windows Programming. Wiley, New York, NY, USA, 1992. ISBN 0-471-54711-5 (book), 0-471-55172-4 (book/disk set). xi + 353 pp. LCCN QA76.76.W56 H45 1992. US$32.95.

Heller:1993:AWP

[Hel93] Martin Heller. Advanced Win32 programming. Wiley, New York, NY, USA, 1993. ISBN 0-471-59245-5. xiii + 463 pp. LCCN QA76.76.A65 H45 1993. US$44.95.

Hetherington:1998:CSU

[Het98]  David Hetherington.  Case

study: Unicode in large network systems management. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Huang:1989:ICJ

[HH89] Jack K. T. Huang and Timothy D. Huang. Introduction to Chinese, Japanese and Korean Computing, volume 12 of World Scientific Series in Computer Science. World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, June 1989. ISBN 9971-5-0664-5. xxi + 437 pp. LCCN QA76.H7795 1989.

Haralambous:2007:FE

[HH07] Yannis Haralambous and P. Scott Horne. Fonts and encodings. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-10242-9. xx + 1016 pp. LCCN Z250.

URL http://www.oreilly.
com/catalog/9780596102425.

Hipson:1994:WEV

[Hip94] Peter D. Hipson. What every Visual C++ 2 programmer should know. Howard W. Sams, Indianapolis, IN 46268, USA, 1994. ISBN 0-672-30493-7. xxviii + 626 pp. LCCN QA76.73.C153 H57 1994. US$29.99, CDN$39.99.

Hiura:1996:UIU

[Hiu96] H. Hiura. Unicode and internationalization with UNIX and X-Window’s system. In Unicode Consortium [Uni96c], pages C3–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Hiura:1997:PIIb

[Hiu97a] H. Hiura. Platform independent input method protocol for Internet, NC, and Java. In Unicode Consortium [Uni97b], pages A14–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Hiura:1997:PIIa

[Hiu97b] Hideki Hiura. Platform independent input method protocol for the Internet, network computer (NC) and Java. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Hiura:1997:UIU

[Hiu97c] Hideki Hiura. Unicode and internationalization with UNIX and the X-Window System. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Hiura:1998:IDI

[Hiu98] Hideki Hiura. IIIMF — distributed input method architecture for the Internet, network computer, and Java. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Hiura:1998:UST

[HK98] Hideki Hiura and Tatsuo Kobayashi. The Unicode standard: a tutorial. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc12/program.
html.

Hadley:2001:SSF

[HM01] Joshua Hadley and Dirk Meyer.

Script-specific font features. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN

????  URL http://www.
agfamonotype.com/presentations/

.

Harold:2004:XND

[HM04] Elliotte Rusty Harold and W. Scott Means. XML in a Nutshell: [a Desktop Quick Reference]. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00764-7. xix + 689 pp. LCCN QA76.76.H94; Internet; QA76.76.H94 H36 2004eb; QA76.76.H94 H36 2004.

Hiura:2006:IVD

[HM06] Hideki Hiura and Eric Muller.

Ideographic variation database. Unicode Technical Standard 37, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, January 13,
2006.  URL http://www.
unicode.org/draft/reports/
tr37/tr37.html;   http://
www.unicode.org/unicode/reports/
tr37/.

Handa:1996:UEM

[HNTT96] Kenichi Handa, Mikiko Nishikimi, Satoru Tomura, and Naoto Takahashi. Unified and extensible mechanism for multilingual text processing, 1996.

                                        URL ftp://etlport.etl.go.
                                        jp/pub/mule/papers/PRICAI96.
                                        tex.gz.

Ho:1996:CCU

[Ho96] Y.-F. Ho. Challenge of creating a Unicode WWW homepage. In Unicode Consortium [Uni96c], pages A6–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Ho:1997:PPR

[Ho97]  Yean-Fee Ho.  A platform

to publish and retrieve multilingual information on the WWW. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Hodgson:1998:SWT

[Hod98] Andrew Hodgson. Searching the world’s text — Unicode is only the first step. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc13/program.html.

Hoffnagle:1996:Pb

[Hof96]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 35(2):122–123, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-2.html#one.

Hohberger:1997:UBS

[Hoh97] Clive P. Hohberger. Ultracode:

a barcode symbology encoding Unicode 2.0. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Holmes:2003:PPU

[Hol03] Neville Holmes. The profession:

The problem with Unicode. Computer, 36(6):116, 114–115, June 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/06/r6116.
htm;   http://csdl.computer.
org/dl/mags/co/2003/06/r6116.
pdf.

Honomichl:1996:MBU

[Hon96] L. Honomichl. Mapping between Unicode and other character sets: Methods and issues. In Unicode Consortium [Uni96c], pages B4–?? ISBN ???? LCCN QA268 .I56 1996. URL http:/

/www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Honomichl:1997:CMM

[Hon97]  Lloyd Honomichl.  Charac-

ter mapping: Methods and issues. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Honomichl:1998:ESU

[Hon98a] Lloyd Honomichl. Efficient storage and use of Unicode property data. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Honomichl:1998:MBU

[Hon98b] Lloyd Honomichl. Mapping between Unicode and other character sets: Methods and issues. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Horstmann:2012:JEC

[Hor12] Cay S. Horstmann. Java for everyone: compatible with Java 5, 6, and 7. Wiley, New York, NY, USA, second edition, 2012. ISBN 1-118-06331-7 (paperback). xxxiii + 589 pp. LCCN QA76.73.J38 H675445 2012.

Hossain:2024:TSR

[Hos24] Anushah Hossain. Text standards for the “rest of world”: The making of the Unicode standard and the OpenType format. IEEE Annals of the History of Computing, 46(1): 20–33, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hosken:2004:RMU

[HT04] Martin Hosken and Maung Tuntunlwin. Representing Myanmar in Unicode: Details and examples. Unicode Technical Note 11, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, January 9, 2004.

URL http://www.unicode.
org/notes/tn11/;   http://
www.unicode.org/notes/tn11/
myanmar_uni.pdf.

Hudson:2000:WGP

[Hud00] John Hudson. Windows glyph processing. World-Wide Web document., November 7, 2000.

URL http://www.microsoft.
com/typography/developers/
opentype/default.htm.

Hufflen:2017:MNH

[Huf17] Jean-Michel Hufflen. MlBibTEX now handles Unicode. TUGboat, 38(2):245–248, ???? 2017. CODEN ???? ISSN 0896-3207.

URL https://www.tug.org/
TUGboat/tb38-2/tb119hufflen-
mlbibtex.pdf.

IEEE:1990:IIC

[IEE90] IEEE, editor. 1990 IEEE International Conference on Systems, Man, and Cybernetics: conference proceedings, November 4–7, 1990, Sheraton Universal Hotel, Los Angeles, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN PICYE3. ISBN 0-87942-597-0. ISSN 0884-3627. LCCN TA 168 I19 1990. IEEE catalog number 90CH2930-6.

IEEE:1994:PII

[IEE94] IEEE, editor. Proceedings of the 12th IAPR International Conference on Pattern Recognition. Jerusalem, Israel, 9–13 October, 1994, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6270-0. LCCN TK 7882 P3 I57 1994. Three volumes. IEEE catalog number 94CH3440-5.

Irvine:1997:LEL

[IF97] Aaron Irvine and Ken Fowles. Letter to the editor: Latin-X. Byte Magazine, 22(7):16, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Comments and response on ISO 8859 Roman code pages, and Unicode.

Ishida:1996:NWS

[Ish96a] R. Ishida. Non-Latin writing systems: Characteristics and impacts on multinational product design. In Unicode Consortium [Uni96c], pages C1–?? (or C2–??). ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Ishida:1996:WDL

[Ish96b] R. Ishida. What does localisation involve? In Unicode Consortium [Uni96c], pages C5– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Ishida:1997:CID

[Ish97a] R. Ishida. Cultural issues in developing on-screen information. In Unicode Consortium [Uni97b], pages TB3–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Ishida:1997:NWS

[Ish97b] R. Ishida. Non-Latin writing systems: Characteristics and impact on multinational product design. In Unicode Consortium [Uni97b], pages TB2–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Ishida:1997:IMT

[Ish97c] Richard Ishida. Internationalisation: More than just character sets. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Ishida:1997:NLW

[Ish97d]  Richard Ishida.  Non-Latin

writing systems: Characteristics and impacts on multinational product design. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Ishida:1998:CIS

[Ish98a] Richard Ishida. Cultural issues with on-screen information. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Ishida:1998:MYP

[Ish98b] Richard Ishida. Making your products translatable. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Ishida:1998:NLWc

[Ish98c] Richard Ishida. Non-Latin writing systems: Characteristics and impact on product design. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.

unicode.org/unicode/iuc13/
program.html.

Ishida:1998:NLWa

[Ish98d]  Richard Ishida.  Non-Latin

writing systems: Characteristics and impacts on multinational product design. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Ishida:1998:NLWb

[Ish98e]  Richard Ishida.  Non-Latin

writing systems: Characteristics and impacts on multinational product design — continued. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Ishimoto:2001:POB

[Ish01] Kenya Ishimoto. Preference object-based internationalization for distributed application framework in Java. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/iuc/iuc18/papers/a16.
ppt.

Ishida:2003:IIS

[Ish03] Richard Ishida. An introduction to Indic scripts. Unicode Technical Note 10, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, August 15, 2003.

URL http://www.unicode.
org/notes/tn10/;   http://
www.unicode.org/notes/tn10/
tn10-1.html.

Ishii:2013:UVT

[Ish13] Koji Ishii. Unicode vertical text layout. Technical Report 50, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, August 13, 2013.

URL http://www.unicode.
org/reports/tr50/.

Jovanovic:2009:SSC

[J^{+}09] Gordana Jovanović et al., editors. Standardizacija staroslovenskog c irilic kog pisma i njegova registracija u Unikodu: zbornik radova sa me unarodnog nauc nog skupa odrz anog od 15. do 17. oktobra 2007. godine, volume 125 of Nauvcni skupovi (Srpska akademija nauka i umetnosti). Srpska akademija nauka i umetnosti, Beograd, Serbia, 2009. ISBN 86-7025-494-8. LCCN PG637 .S68 2007.

Juneau:2012:JRP

[J^{+}12] Josh Juneau et al. Java 7 recipes: a problem-solution approach. Apress, Berkeley, CA, USA, 2012. ISBN 1-4302-4056-3 (paperback). xxii + 848 pp. LCCN QA76.73.J38 J35946 2012.

Jerman-Blazic:1998:UEM

[JBC98] Borka Jerman-Blazic and Konstantin Chuguev. Unicodeenabled multilingual WWW server. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Jerman-Blazic:1997:SSE

[JBCG97] B. Jerman-Blazic, K. Chuguev, and A. Gogala. Standard specification extensions for provision of language and character enabled server. Computer Communications, 20(11): 927–936, October 15, 1997. CO-DEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

Jerman-Blazic:1997:MFS

[JBGC97] Borka Jerman-Blazic, Andrej Gogala, and Konstantin V. Chuguev. Migration to a fully supported Unicode WWW server. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Jenkins:2008:USA

[JC08] John Jenkins and Richard Cook. Unicode standard annex #38: Unicode Han database (Unihan). Technical Report 38, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, March 28, 2008.

                                        URL http://www.unicode.
                                        org/reports/tr38/;   http:
                                        //www.unicode.org/reports/
                                        tr38/tr38-5.html.

Jensen:1970:SSS

[Jen70] Hans Jensen. Sign, symbol, and script: an account of man’s efforts to write. Allen & Unwin, London, UK, third edition, 1970. ISBN 0-04-400021-9. 613 pp. LCCN P211 .J413 1970. Translated from the German “Geschichte der Schrift” by George Unwin.

Jenkins:1997:HCUb

[Jen97a] J. H. Jenkins. Handling complex Unicode rendering problems. In Unicode Consortium [Uni97b], pages B11–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/program.html.

Jenkins:1997:UDM

[Jen97b] J. H. Jenkins. Unihan: Dispelling the myths. In Unicode Consortium [Uni97b], pages B14–?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Jenkins:1997:HCUa

[Jen97c] John H. Jenkins. Handling complex Unicode rendering problems. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.

unicode.org/unicode/iuc10/
program.html.

Jenkins:1998:IUF

[Jen98] John H. Jenkins. Ideographs in Unicode’s future. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc13/program.
html.

Jenkins:2008:DUT

[Jen08]  John Jenkins.  Draft Uni-

code technical report #45: U-source ideographs. Technical Report 45, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, June 4, 2008.

URL http://www.unicode.
org/reports/tr45/;   http:
//www.unicode.org/reports/
tr45/tr45-1.html.

Jenkins:2012:USI

[Jen12]  John Jenkins.  U-source

ideographs. Unicode Standard Annex 45, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, September 12,

2012.  URL http://www.
unicode.org/reports/tr45/.

Jordan:2002:LLB

[Jor02] D. K. Jordan. Languages left behind: Keeping Taiwanese off the World Wide Web. Language Problems & Language Planning, 26(2):111–127, August 2002. CODEN ???? ISSN 0272-2690.

Jung:1997:USS

[Jun97] Chang Hoon Jung. Unicode support in Solaris 2.6: The Korean UTF-8 multiscript locale. In Unicode Consortium [Uni97b], pages C5–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/program.html.

Kano:1995:DIS

[Kan95] Nadine Kano. Developing International Software for Windows 95 and Windows NT. Microsoft Press, Bellevue, WA, USA, 1995. ISBN 1-55615-840-8. xvi + 743 pp. LCCN QA76.76.D47 K36 1995. URL

                                        http://msdn.microsoft.com/
                                        library/books/devintl/S24AB.
                                        HTM;   http://www.microsoft.
                                        com/globaldev/wrguide/wrguide.
                                        asp.

Kaplan:2001:SSM

[Kap01a] Michael Kaplan. Surrogate support in Microsoft products. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

                                        www.unicode.org/iuc/iuc18/
                                        papers/a8.ppt.

Kaplan:2001:UCS

[Kap01b] Michael Kaplan. Unicode and collation support in Microsoft SQL Server. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/b3.ppt.

Kaplan:2001:UVB

[Kap01c] Michael Kaplan. Unicode and Visual Basic: a case study. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/iuc/iuc18/papers/a18.
ppt.

Kaplan:2005:TN

[Kap05]  Michael Kaplan.  Tamil

numbers. Unicode Technical Note 21, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, March 11, 2005.

URL http://www.unicode.
org/notes/tn21/;   http://
www.unicode.org/notes/tn21/
tamil_numbers.pdf;   http:
//www.unicode.org/notes/tn21/
tn21-1.html.

Kelly:2016:NRH

[KBKHF16] Jerry Kelly, Misha Beletsky, Jerry Kelly, and Amelia Hugill-Fontanel. The noblest roman: a history of the Centaur types of Bruce Rogers, volume 236 of Book Club of California publication. The Book Club of California, San Francisco, CA, USA, 2016. ISBN 0-9819597-9-2. 120 + 16 + 2 pp. LCCN Z232.R67 K45 2016.

Keber:1997:UGEa

[Keb97] Wolfgang Keber. Unicode — the glue in the European coding world — some case studies. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Kew:2005:TML

[Kew05a] Jonathan Kew. The multilingual lion: TEX learns to speak Unicode. In Unicode Consortium [Uni05b], pages n+1– n+17. ISBN ???? LCCN ????

Kew:2005:XML

[Kew05b] Jonathan Kew. XETEX, the Multilingual Lion: TEX meets Unicode and smart font technologies. TUGboat, 26(2): 115–124, 2005. ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb26-2/kew.pdf.

Kew:2006:UMT

[Kew06]  Jonathan Kew.  Unicode

and multilingual typesetting with XETEX. TUGboat, 27 (2):228–229, 2006. ISSN 0896-3207. URL http://

                                        www.tug.org/TUGboat/tb27-
                                        2/tb87kew.pdf.

Kobayashi:1998:ACT

[KH98] Tatsuo Kobayashi and Hideki Hiura. Advanced CJK topics. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Kida:1998:DMT

[Kid98] Yasuo Kida. Dispelling the myths: The truth about Unihan. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Kim:1992:CAD

[Kim92a] Kyongsok Kim. A common approach to designing the Hangul code and keyboard. Computer Standards and Interfaces, 14(4): 297–325, August 1992. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic). Discusses the Unicode encoding of Hangul.

Kim:1992:FDS

[Kim92b] Kyongsok Kim. A future direction in standardizing international character codes — with a special reference to ISO/IEC 10646 and Unicode. Computer Standards and Interfaces, 14 (3):209–221, May 1992. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Kinch:1996:ETU

[Kin96] Richard J. Kinch. Extending TEX for Unicode. TUGboat, 17 (2):147–160, June 1996. ISSN 0896-3207.

King:2012:WUS

[Kin12] R. S. King. Will Unicode soon be the universal code? [the data]. IEEE Spectrum, 49(7):60, July 2012. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keiser:2021:VUL

[KL21] John Keiser and Daniel Lemire. Validating UTF-8 in less than one instruction per byte. Software—Practice and Experience, 51(5):950–964, May 2021. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kline:1998:UIEa

[Kli98a] Susan L. Kline. Unicode: An interoperability enabler in HP-UX. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Kline:1998:UIEb

[Kli98b] Susan L. Kline. Unicode: An interoperability enabler in HP-UX. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Kobayashi:1998:W

[Kob98] Tatsuo Kobayashi. Welcome. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Korpela:2006:UE

[Kor06] Jukka Korpela. Unicode Explained. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006.

ISBN 0-596-10121-X. xvii + 658 pp. LCCN QA268 .K67 2006. US$59.99. URL http:

//www.oreilly.com/catalog/
unicode/index.html.

Kamei:1998:GDE

[KRE98] Takashi Kamei, Kono Rokuro, and Chino Eiichi, editors. Gengogaku daijiten [English: The Sanseido encyclopaedia of linguistics]. Shohan, Tokyo, Japan, 1998. ISBN 4-385-15215-2 (vol. 1), 4-385-15213-6 (vol. 1), 4-385-15216-0 (vol. 2), 4-385-15217-9 (vol. 3), 4-385-15212-8 (vol. 4), 4-385-15214-4 (vol. 5), 4-385-15218-7 (vol. 6). LCCN P29 .G4 1988. 48,000 Yen. In Japanese.

Kristian:1997:IUL

[Kri97] J. Kristian. Internationalization and Unicode in LDAP directory services. In Unicode Consortium [Uni97b], pages A3–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Ksar:1996:IUC

[Ksa96] M. Ksar. ISO 10646 — Unicode:

Continued convergence. In Unicode Consortium [Uni96c], pages C1–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Ksar:1997:IIU

[Ksa97a] M. Ksar. ISO/IEC 10646 and Unicode — continued convergence. In Unicode Consortium [Uni97b], pages C11–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/program.html.

Ksar:1997:IUC

[Ksa97b] Michael Ksar. ISO 10646– Unicode: Continued convergence. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Ksar:1998:UIAa

[Ksa98a] Mike Ksar. Unicode and ISO 10646: Achievements and directions. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Ksar:1998:UIAb

[Ksa98b] Mike Ksar. Unicode and ISO 10646: Achievements and directions. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Keber:1997:UGEb

[KT97] W. Keber and L. Tossounidis. Unicode — the glue in the European coding world — some case studies. In Unicode Consortium [Uni97b], pages A8–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Kuepper:2005:UFS

[Kue05] Tilman Kuepper. Unicode and filtering stream buffers. C/C++ Users Journal, 28(3):28, 30–33, March 2005. CODEN CCU-JEX. ISSN 1075-2838.

Kuhn:1999:AUQ

[Kuh99a] Markus Kuhn. ASCII and Unicode quotation marks. World-Wide Web document., 12 19, 1999. URL http:/

/www.cl.cam.ac.uk/~mgk25/
ucs/quotes.html.

Kuhn:1999:UUF

[Kuh99b] Markus Kuhn. UTF-8 and Unicode FAQ for Unix/Linux. Web site, June 4, 1999. URL https:

//www.cl.cam.ac.uk/~mgk25/
unicode.html#linux. Last re-

vision: 2009-05-11.

Kumhyr:1998:DIC

[Kum98] David B. Kumhyr. Developing an international calendar in Java. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Kunst:1994:UIJ

[Kun94] R. Kunst. A Unicode implementation of Japanese, Korean, Thai, and Chinese in Win-CALIS 2.0. In Borchardt and Johnson [BJ94], pages 142–??

Kuster:2005:NU

[Küs05] Johannes Küster. Newmath and Unicode. TUGboat, 27 (0):165, March 2005. CO-DEN ???? ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb27-0/kuster.pdf.

Kwan:1995:UIT

[Kwa95a] P. Kwan. UNICODE: implications for the translator. In Anonymous [Ano95b], pages 1– 15. ISBN 0-85142-351-5. LCCN P308 .T724 1995.

Kwan:1995:UUC

[Kwa95b] Peter Kwan. Unicode: a universal character set. Language International, 7(4):5–??, 1995. CODEN LAINE2. ISSN 0923-182X.

LaBonte:1993:MOU

[La 93] A. La Bonté. Multiscript ordering for UNICODE. In SHARE Europe ’93 [SHA93], pages 377– 399. ISBN ???? LCCN ????

LaBonte:2000:FII

[LaB00] Alain LaBonté. FAQ: ISO/ IEC 106461 versus Unicode. ;login: the USENIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        standards/38.labonte.html.

Lu:2008:ULT

[LBD08] C. J. Lu, A. C. Browne, and G. Divita. Using lexical tools to convert Unicode characters to ASCII. In Annual Symposium proceedings / AMIA Symposium. AMIA Symposium, 2008, volume 1031, page ?? ????, ????, 2008. US$.

Leggott:1991:UUU

[Leg91] Mark Leggott. Unique, universal and uniform character encoding (Unicode). Canadian Library Journal, 48(5):345–346, October 1991. CODEN CLI-JBX. ISSN 0008-4352.

Leisher:1996:AIUa

[Lei96a] M. Leisher. An adventure in implementing Unicode support on UNIX platforms. In Unicode Consortium [Uni96b], pages C6–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Leisher:1996:AIUc

[Lei96b] M. Leisher. An adventure in implementing Unicode support on UNIX platforms. In Unicode Consortium [Uni96c], pages C6– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:

                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Leisher:1996:IMDa

[Lei96c] M. Leisher. Input method design. In Unicode Consortium [Uni96b], pages A1–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Leisher:1996:IMDc

[Lei96d] M. Leisher. Input method design. In Unicode Consortium [Uni96c], pages A1– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Leisher:1996:AIUb

[Lei96e] Mark Leisher. An adventure in implementing Unicode support on Unix platforms. In Unicode Consortium [Uni96c], page 8. ISBN ???? LCCN QA268 .I56

                                        1996. URL ftp://crl.nmsu.
                                        edu/CLR/multiling/unicode/
                                        paper.ps.gz.

Leisher:1996:IMDb

[Lei96f] Mark Leisher. Input method design. tutorial. In Unicode Consortium [Uni96c], page 18. ISBN ???? LCCN QA268 .I56

1996. URL ftp://crl.nmsu.
edu/CLR/multiling/unicode/
tutorial.ps.gz.

Lemberg:1997:CPL

[Lem97] Werner Lemberg. The CJK package for LaTEX2e — multilingual support beyond babel. TUGboat, 18(3):214–224, September 1997. ISSN 0896-3207.

Leong:1998:UJI

[Leo98] Kok Yong Leong. Unicode, Java and input methods. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Lerner:2003:FU

[Ler03] Reuven M. Lerner. At the forge:

Unicode. Linux Journal, 107:??, March 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

noframes.linuxjournal.com/
lj-issues/issue107/article.
php?sid=6504.

Lerner:2013:FU

[Ler13] Reuven M. Lerner. At the forge:

Unicode. Linux Journal, 2013 (230):6:1–6:??, June 2013. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Leventhal:1998:XCD

[Lev98]  M. Leventhal.  XML. can

the desperate Perl hacker do it? Web Techniques, 3(3): 45–49, March 1998. CODEN WETEFA. ISSN 1086-556X.

Li:1998:DUG

[Li98] Yan Li. Developing a Unicode GIS application for Microsoft Windows and UNIX. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Lilley:1997:DIW

[Lil97] C. Lilley. The design of an international Web font extension for Cascading Style Sheets. In Unicode Consortium [Uni97b], pages B6–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Lilley:1998:DIW

[Lil98a] Chris Lilley. The design of an International Web Font Extension for Cascading Style Sheets. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Lilley:1998:RMD

[Lil98b] Chris Lilley. Rendering multilingual documents — CSS and XSL. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Lindenberg:1998:JIM

[Lin98] Norbert Lindenberg. Java input method framework. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Liu:1997:CCT

[Liu97] L. Y. Liu. Checking Chinese text errors in the Unicode environment. In Unicode Consortium [Uni97b], pages B15–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Liu:1994:PMA

[LL94a] R. Liu and J. Lions. A platform for multilingual application development with Unicode standard. In Tan [Tan94], pages 228–234. ISBN 981-00-5648-6. LCCN ????

Liu:1994:DRU

[LL94b] Raymond Liu and John Lions. Data representation with Unicode standard in presentation layer of OSI. In National Conference Publication — Institution of Engineers, Australia Proceedings of the International Symposium on Information Theory & Its Applications 1994. Part 1 (of 2) Nov 20–24 1994, pages 589– 592. University of New South Wales, Kensington, NSW, Australia, November 1994. CODEN NPIEDX. ISBN ???? ISSN 0313-6922. LCCN ????

Liu:1998:DMA

[LL98] Raymond Liu and John Lions. Distributed multilingual applications of the OSI architecture. Computer Standards and Interfaces, 19(1):9–29, January 1998. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Lemire:2022:TBU

[LM22] Daniel Lemire and Wojciech Muła. Transcoding billions of Unicode characters per second with SIMD instructions. Software—Practice and Experience, 52(2):555–575, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Longley:2000:RUN

[Lon00] Ian Longley. The role of Unicode in natural language classification. Technical report, School of Computer Studies, University of Leeds, Leeds, UK, 2000. Final year project.

Lu:2001:ICS

[Lu01] Qin Lu. The ideographic composition scheme and its applications in Chinese text processing. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

                                        www.unicode.org/iuc/iuc18/
                                        papers/b16.ppt.

Lua:1990:PMC

[Lua90] K. T. Lua. A proposal for multilingual computing code standardization. Computer Standards and Interfaces, 10(2): 117–124, ???? 1990. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Lunde:1993:UJI

[Lun93a] Ken Lunde. Understanding Japanese Information Processing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-043-0. xxxii + 435 pp. LCCN PL524.5.L86 1993. US$29.95.

Lunde:1993:HJC

[Lun93b] Ken R. Lunde. The history of the Japanese character set and its encoding. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 7(1):85–94, June 1993. CODEN CPCLE6. ISSN 0715-9048.

Lunde:1996:DEJ

[Lun96] K. Lunde. Designing an extended Japanese character set: From conception to implementation. In Unicode Consortium [Uni96c], pages B7– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Lunde:1998:AUG

[Lun98a]  Ken Lunde.  Accessibility

of unencoded glyphs. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL ftp://ftp.
                                        oreilly.com/pub/examples/
                                        nutshell/ujip/unicode/iuc13-
                                        a10-paper.pdf; http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Lunde:1998:ICL

[Lun98b]  Ken Lunde.  Implementing

cross-locale CJKV code conversion. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL ftp://ftp.

                                        oreilly.com/pub/examples/
                                        nutshell/ujip/unicode/iuc13-
                                        c2-paper.pdf; http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Lunde:1999:CIP

[Lun99] Ken Lunde. CJKV Information Processing: Chinese, Japanese, Korean & Vietnamese Computing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-224-7. 1174 pp. LCCN PL1074.5 .L85 1999. US$64.95. URL http:

                                        //www.oreilly.com/catalog/
                                        cjkvinfo/.

MacKay:1986:TPS

[Mac86] Pierre A. MacKay. Typesetting problem scripts. computer typesetting provides a solution for Arabic and similar scripts.

Byte Magazine, 11(2):201–218, February 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moseley:2006:AWL

[MAD^{+}06] Christopher Moseley, R. E.

Asher, Giles Darkes, et al., editors. Atlas of the world’s languages. Routledge, London, UK, second edition, 2006. ISBN 0-415-31074-1. ???? pp. LCCN G1046.E3 A8

2007.  US$570.00.  URL

http://www.loc.gov/catdir/
enhancements/fy0711/2007060125-
d.html.

Mahajan:2009:IUO

[Mah09] Aditya Mahajan. Integrating Unicode and OpenType math in ConTEXt. TUGboat, 30(2):243–246, 2009. ISSN 0896-3207. URL http://

www.tug.org/TUGboat/tb30-
2/tb95mahajan-cmath.pdf.

Mansour:1997:UFb

[Man97a] K. Mansour. Unicode and fonts. In Unicode Consortium [Uni97b], pages B3–?? (or B4–??). ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Mansour:1997:UFa

[Man97b] Kamal Mansour. Unicode & fonts. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Mansour:1998:UFa

[Man98a] Kamal Mansour. Unicode & fonts. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Mansour:1998:UFb

[Man98b] Kamal Mansour. Unicode and fonts. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Mansour:1998:UFC

[Man98c]  Kamal Mansour.  Unicode

and fonts — continued. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Mansour:2001:NSU

[Man01] Kamal Mansour. New scripts in Unicode 3.0. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        agfamonotype.com/presentations/

.

Maruyama:1998:UBI

[Mar98] Takashi Maruyama. Unicode based input method and Hancharacter search. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ????

                                        URL http://www.unicode.
                                        org/unicode/iuc12/program.
                                        html.

Matsuzuka:1996:MIM

[Mat96] S. Matsuzuka. Multilanguage input model on X-Window System. In Unicode Consortium [Uni96c], pages C4– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Mayeur:1997:NGD

[May97] Thierry Mayeur. Notes Global Designer — developing multilingual Notes applications for the world. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

McFarland:1997:DIS

[MBH97] T. McFarland, B. Beck, and H. Hiura. Developing internationalized software with JDK 1.1. In Unicode Consortium [Uni97b], pages TA2–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

McCarthy:1972:RAC

[McC72] John McCarthy. RFC 373:

Arbitrary character sets, July 14, 1972. URL ftp://ftp.

internic.net/rfc/rfc373.txt;
ftp://ftp.math.utah.edu/
pub/rfc/rfc373.txt. Status:

UNKNOWN.

McConnell:1996:UIM

[McC96] J. McConnell. Unicode and internationalization on Mac OS 8. In Unicode Consortium [Uni96c], pages B3– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

McFarland:1997:EDI

[McF97] T. McFarland. Experiences developing internationalized applets with Java. In Unicode Consortium [Uni97b], pages A15–?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

McFarland:1998:DIA

[McF98a]  Tom McFarland.  Develop-

ing internationalized applications/applets with Java. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

McFarland:1998:DIS

[McF98b] Tom McFarland. Developing internationalized software with Java. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

McGowan:1996:DNT

[McG96] R. McGowan. Designing the next text system. In Unicode Consortium [Uni96c], pages B4– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

McGowan:1998:RP

[McG98] Rick McGowan. The Rhapsody platform. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

McGowan:2002:GMR

[McG02]  Rick McGowan.  A gen-

eral method for rendering combining marks. Unicode Technical Note 2, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, October 28,
2002.  URL http://www.
unicode.org/notes/tn2/.

RFC3718

[McG04] R. McGowan. RFC 3718: a summary of Unicode Consortium procedures, policies, stability, and public access, February 2004. URL ftp://ftp.

internic.net/rfc/rfc3718.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc3718.txt. Sta-

tus: INFORMATIONAL.

McKenna:1996:CMS

[McK96a] M. McKenna. Conversion of a multinational software developer to Unicode: a case history to-date. In Unicode Consortium [Uni96c], pages C2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

McKenna:1996:UDC

[McK96b] M. McKenna. Unicode in distributed computing environments. In Unicode Consortium [Uni96c], pages B1– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

McKenna:1997:CMSb

[McK97a] M. McKenna. Converting a multinational software developer to Unicode: a case history to-date. In Unicode Consortium [Uni97b], pages A6–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/program.html.

McKenna:1997:UWD

[McK97b] M. McKenna. Unicode in worldclass database systems. In Unicode Consortium [Uni97b], pages B1–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

McKenna:1997:CMSa

[McK97c] Mike McKenna. Conversion of a multinational software developer to Unicode — A casehistory-to-date. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc10/program.
html.

McKenna:1997:UDC

[McK97d] Mike McKenna. Unicode in distributed computing environments. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

McKenna:1997:UUW

[McK97e] Mike McKenna. Use of Unicode by world-class database vendors. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

McKenna:1998:GIS

[McK98a] Michael G. McKenna. Global information systems. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

McKenna:1998:UUW

[McK98b] Mike McKenna. Use of Unicode by world-class database vendors. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

McCallum:1994:ASA

[ME94] Sally McCallum and Monica Ertel, editors. Automated systems for access to multilingual and multiscript library materials: proceedings of the Second IFLA Satellite Meeting, Madrid, August 18–19, 1993, volume 70 of IFLA publications. K. G. Saur, München, Germany; New Providence, NJ, USA, 1994. ISBN 3-598-21797-8. ISSN 0344-6891. LCCN Z699.5.F67 A872 1994.

Meersseman:1997:DBD

[Mee97] Yann Meersseman. Double the bytes, double the trouble. Language International, 9(6):18–19, ???? 1997. CODEN LAINE2. ISSN 0923-182X.

Meyer:2001:TNC

[Mey01] Dirk Meyer. Two new coded character standards from China and their implementation: HK SCS & GB 18030. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN

????  URL ftp://ftp.
oreilly.com/pub/examples/
nutshell/cjkv/pdf/IUC18B17.
pdf.

Mao:2007:UUC

[MfFH07] Ching-Chen Anthony Mao and Ching fen Frances Hsu. UNI-MARC and UNICODE — Chinese MARC (Taiwan) and its bibliographic database. International cataloguing and bibliographic control: quarterly bulletin of the IFLA UBCIM Programme, 36(3):58, ???? 2007. ISSN 011-8829.

Myers:1993:MWN

[MH93] Brian Myers and Eric Hamer.

Mastering Windows NT programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1993. ISBN 0-7821-1264-1. xxxv + 1236 pp. LCCN QA76.76.O63 M94 1993. US$44.99.

McClure:1995:CGA

[MH95] Wanda L. McClure and Stan A.

Hannah. Communicating globally: The advent of Unicode. Computers in Libraries, 15(5): 19–24, May 1995. CODEN CPLIE8. ISSN 1041-7915.

Microsoft:1999:CDS

[Mic99] Microsoft Corporation. Character design standards. Technical report, Microsoft Corporation, Redmond, WA, USA, July 8, 1999. URL http://www.

microsoft.com/typography/
developers/fdsspec/default.
htm.

Mikiko:1998:MUE

[Mik98]  Nishikimi Mikiko.  MULE-

Universal enCoding System: a flexible codepoint translation mechanism. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Milo:1996:ESN

[Mil96]  T. Milo.  Encoding sense

and nonsense: The Arabic ligature. In Unicode Consortium [Uni96c], pages B8– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Milo:1997:TAC

[Mil97a] T. Milo. Towards Arabic computer typography. In Unicode Consortium [Uni97b], pages B8–?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Milo:1997:ESN

[Mil97b] Thomas Milo. Encoding sense and nonsense — the Arabic ligature. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Milo:1998:AAA

[Mil98a] Thomas Milo. Authentic Arabic — aesthetic and technical challenges. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Milo:1998:AAB

[Mil98b] Thomas Milo. Authentic Arabic — backgrounds. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc13/program.
html.

Milo:2003:AAB

[Mil03] Thomas Milo. ALI-BABA and the 4.0 Unicode characters — New input and output concepts under Unicode. TUGboat, 24 (3):502–511, 2003. ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb24-3/milo.pdf.

Mittelbach:2021:PUP

[Mit21a] Frank Mittelbach. The uni-

codefonttable package. TUG-

boat, 42(3):287–304, ?? 2021. CODEN ???? ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb42-3/tb132mitt-
unicodefonttable.pdf.

Mittelbach:2021:PTP

[Mit21b] Frank Mittelbach. Preventing tofu with and Unicode engines. TUGboat, 42(3):284– 286, ?? 2021. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb42-3/
tb132mitt-tofu.pdf.

Mabry:2007:USE

[MJF07] Frank J. Mabry, John R. James, and Aaron J. Ferguson. Unicode steganographic exploits: Maintaining enterprise border security. IEEE Security & Privacy, 5(5):32–39, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Matsuoka:1997:KDI

[MKG97] Eiji Matsuoka, Tatsuo Kobayashi, and Koji Gotoda. Kanji database and input method of ‘Itaiji’. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Mair:1991:CC

[ML91] Victor H. Mair and Yongquan Liu, editors. Characters and computers. IOS Press, Amsterdam, The Netherlands, 1991. ISBN 90-5199-061-8. LCCN PL1074.5 .C372 1991.

Mainland:2013:EVI

[MLJ13] Geoffrey Mainland, Roman Leshchinskiy, and Simon Peyton Jones. Exploiting vector instructions with generalized stream fusion. ACM SIGPLAN Notices, 48(9):37–48, September 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mainland:2017:EVI

[MLJ17] Geoffrey Mainland, Roman Leshchinskiy, and Simon Peyton Jones. Exploiting vector instructions with generalized stream fusion. Communications of the ACM, 60(5):83–91, May 2017. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//cacm.acm.org/magazines/
2017/5/216312/fulltext.

Mudur:1999:ASI

[MNSJ99a] S. P. Mudur, N. Nayak, S. Shanbhag, and R. K. Joshi. An architecture for the shaping of Indic texts. Computers and Graphics, 23(1):7–24, February 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mudur:1999:CGI

[MNSJ99b] S. P. Mudur, Niranjan Nayak, Shrinath Shanbhag, and R. K. Joshi. Computer graphics in India — an architecture for the shaping of Indic texts. Computers and Graphics, 23 (1):7–24, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://

www.elsevier.com/cas/tree/
store/cag/sub/1999/23/1/647.
pdf.

Moorcraft:1997:URC

[Moo97] Colin Moorcraft. A Unicode revolution comes to the Web. Language International, 9(3): 22–23, 36, ???? 1997. CODEN LAINE2. ISSN 0923-182X.

Moore:1998:USV

[Moo98a] Lisa Moore. The Unicode standard, version 2.1. Unicode Technical Report 8, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, February

18, 1998. URL http:/

                                        /www.unicode.org/unicode/
                                        reports/tr8.html.

Moore:1998:W

[Moo98b] Lisa Moore. Welcome. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Mori:1998:ULB

[Mor98]  Nobuyoshi Mori.  Unicode

and largest business application software. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc12/program.html.

Murray:1993:APW

[MP93] William H. Murray III and Chris H. Pappas. Application programming for Windows NT. Osborne/McGraw-Hill, Berkeley, CA, USA, 1993. ISBN 0-07-881933-4. xxiv + 680 pp. LCCN QA76.76.A65M88 1993.

Miklavec:2008:PCB

[MR08] Mojca Miklavec and Arthur Reutenauer. Putting the Cork back in the bottle — Improving Unicode support in TEX. TUGboat, 29(3):454–457, 2008. ISSN 0896-3207. URL http:/

                                        /www.tug.org/TUGboat/tb29-
                                        3/tb93miklavec.pdf.

Miklavec:2016:HTE

[MR16] Mojca Miklavec and Arthur Reutenauer. Hyphenation in TEX and elsewhere, past and future. TUGboat, 37 (2):209–213, 2016. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb37-2/
tb116miklavec.pdf.

Metcalf:2011:MFE

[MRC11] Michael Metcalf, John Ker Reid, and Malcolm Cohen. Modern Fortran explained. Numerical mathematics and scientific computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-960142-9 (paperback), 0-19-960141-

0. xx + 488 pp. LCCN QA76.73.F235 M48 2011. URL

http://www.loc.gov/catdir/
enhancements/fy1312/2012360954-
b.html;   http://www.loc.
gov/catdir/enhancements/fy1312/
2012360954-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy1312/2012360954-t.html.

Mudawwar:1997:MTM

[Mud97] Muhammed F. Mudawwar.

Multicode: a truly multilingual approach to text encoding. Computer, 30(4):37–43, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also response and rebuttal [DM97], and letters [CPJ^{+}98].

Muller:2005:BB

[Mul05] Eric Muller. To the BMP and beyond. Unicode Technical Note 23, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, May 24, 2005.

                                        URL http://www.unicode.
                                        org/notes/tn23/;   http://
                                        www.unicode.org/notes/tn23/
                                        Muller-Slides+Narr.pdf;
                                        http://www.unicode.org/notes/
                                        tn23/Muller-Slides.pdf;
                                        http://www.unicode.org/notes/
                                        tn23/tn23-1.html.

Muller:2008:USA

[Mul08] Eric Muller. Unicode standard annex #42: Unicode character database in XML. Technical Report 42, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, April 1, 2008.

                                        URL http://www.unicode.
                                        org/reports/tr42/;   http:
                                        //www.unicode.org/reports/
                                        tr42/tr42-3.html.

Munroe:2019:CHU

[Mun19] Randall Munroe. Comic: The history of Unicode. TUGboat, 40(3):278, ???? 2019. CODEN ???? ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb40-3/tb126chest.
                                        pdf.

Murray:1996:SML

[Mur96] R. Murray. Supporting multiple languages and character sets on the Internet. In Unicode Consortium [Uni96c], pages A1– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Murty:1998:UIL

[Mur98] Uppuluri Murty. Unicode issues and localization of software in Indian languages. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Miller:2009:AAA

[MVM09a] Frederic P. Miller, Agnes F.

Vandome, and John McBrewster, editors. Arabic Alphabet: Arabic alphabet, Abjad numerals, ‘i’rab, Arabic diacritics, Arabic numerals, History of the Arabic alphabet, Arabic Unicode, Islamic calligraphy, … Romanization of Arabic, Arabic chat alphabet. Alphascript Publishing, 17 Rue Meldrum, Beau Bassin, 1713-01 Mauritius, 2009. ISBN 613-0-04650-2. 100 (est.) pp. LCCN ???? US$50.

Miller:2009:RNR

[MVM09b] Frederic P. Miller, Agnes F.

Vandome, and John McBrewster, editors. Roman Numerals: Roman numeral analysis, Roman arithmetic, Roman abacus, Kharosthi, Unicode numerals, Etruscan numerals, Positional notation, Arabic numerals, Districts of Turku. Alphascript Publishing, 17 Rue Meldrum, Beau Bassin, 1713-01 Mauritius, 2009. ISBN 613-0-06480-2. 88 (est.) pp. LCCN ???? US$50.

Ngo:2005:MNC

[N^{+}05a] Thanh Nhàn Ngô et al., editors. Manual, the Nôm Na coded character set: with Nôm Na Tˆ óng light true type font and Hán Nôm character knowledge-base: in accordance with Unicode 4.0, ISO/IEC 10646-1:2000 and Vietnamese national standards, TCVN 5712:1993 [VSCII], TCVN 5773:1993 [NSCII], TCVN 6056:1995 [NSCII]. Nôm Na Group, Hanoi, Vietnam, 2005. xiii + 570 pp. LCCN PL4372 .M36 2005.

Nhan:2005:MNCb

[N^{+}05b] Ngo Thanh Nhan et al., editors. Manual, the Nom Na coded character set: with Nom Na Tong light True Type font and Han Nom character knowledgebase: in accordance with Unicode 4.0, ISO/IEC 10646-1:2000 & Vietnamese national standards, TCVN 5712:1993 [VSCII], TCVN 5773:1993 [NSCII], TCVN 6056:1995 [NSCII]. Nom Na Group, Hanoi, Vietnam, 2005. xiii + 570 pp. LCCN PL4372 .M36 2005.

Nakanishi:1998:KAB

[Nak98] Hidehiko Nakanishi. Keynote address: a brief history of the Japanese printing industry and Kanji typesetting. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Nassaney:1996:LMC

[Nas96] B. Nassaney. Localizing multilingual cross-platform software using ForeignDesk. In Unicode Consortium [Uni96c], pages C8– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Nedumaran:2004:TCT

[Ned04]  Muthu Nedumaran.  Text

conversion from TSCII 1.7 to Unicode. Unicode Technical Note 15, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, May 1, 2004.
URL http://www.unicode.
org/notes/tn15/;   http://
www.unicode.org/notes/tn15/
tn15-1.html;   http://www.
unicode.org/notes/tn15/Tscii2Unicode.
pdf.

Needleman:2000:US

[Nee00] M. Needleman. The Unicode Standard. Serials review, 26(2):

51–54, August 2000. CODEN ???? ISSN 0098-7913.

Nestved:1994:MCC

[Nes94] Timothy D. Nestved. Migrating C code to Unicode. Dr. Dobbs Journal, 19(8):28, 30, 32, 91–93, August 1994. CODEN DDJOEB. ISSN 1044-789X.

Thanh:2005:MNC

[Ngô05] Thanh Nhàn Ngô, editor. Manual, the Nôm Na coded character set: with Nôm Na Tˆ óng light true type font & Hán Nôm character knowledgebase: in accordance with Unicode 4.0, ISO/IEC 10646-1:2000 & Vietnamese national standards, TCVN 5712:1993 [VSCII], TCVN 5773:1993 [NSCII], TCVN 6056:1995 [NSCII]. Nôm Na Group, Hanoi, Vietnam, 2005. ISBN ???? xiii + 570 pp. LCCN ????

Nicol:1996:DCG

[Nic96] G. Nicol. Dynaweb: Content generation based on client preferences. In Unicode Consortium [Uni96c], pages A5– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Nicol:1997:MNL

[Nic97a] G. T. Nicol. Making natural language markup natural. In Unicode Consortium [Uni97b], pages A5–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Nicol:1997:SGM

[Nic97b]  G. T. Nicol.  A simple

generic markup language for the WWW. In Unicode Consortium [Uni97b], pages A2–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Nicol:1997:DCG

[Nic97c] Gavin Thomas Nicol. Dynaweb: Content generation based on client preferences. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Nicol:1998:DOM

[Nic98a] Gavin Thomas Nicol. The Document Object Model: What it is, and what it means. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Nicol:1998:MPA

[Nic98b] Gavin Thomas Nicol. Multilingual, and parallel aligned texts: Putting XML and XLL to work. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Nicholas:2006:BMN

[Nic06]  Nick Nicholas.  Byzantine

musical notation. Unicode Technical Note 20, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, March 1, 2006.

URL http://www.unicode.
org/notes/tn20/;   http://
www.unicode.org/notes/tn20/
byznotation.1.1.pdf;   http:
//www.unicode.org/notes/tn20/
tn20-1.html;   http://www.
unicode.org/notes/tn20/tn20-
2.html.

Nielsen:1996:DUD

[Nie96] T. Nielsen. Developing a Unicode DTP application for Windows NT/95. In Unicode Consortium [Uni96c], pages B6– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Nielsen:1997:DUB

[Nie97a] T. Nielsen. Developing a Unicode based DTP application. In Unicode Consortium [Uni97b], pages C6–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Nielsen:1997:DUD

[Nie97b] Thomas Nielsen. Developing a Unicode desktop publishing application based on Windows NT and Windows 95. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Nielsen:1998:DUD

[Nie98] Thomas Nielsen. Developing a Unicode desktop publishing application based on Windows NT and Windows95. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Nieminen:2007:EIA

[NK07] Janne Nieminen and Pekka Kilpeläinen. Efficient implementation of Aho–Corasick pattern matching automata using Unicode. Software—Practice and Experience, 37(6):669–690, May 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Noerr:1994:MCS

[Noe94] Peter Noerr. Multiple character sets for libraries and information services. Proceedings — National Online Meeting, pages 345–348, 1994. CODEN PNOMDR. ISBN 0-938734-84-9. ISSN 0739-1471.

Noji:2001:OSP

[Noj01] Kentaro Noji. Open source project for Unicode locales for Linux using Unicode databases, collation keys and a XML based locale data. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

                                        www.unicode.org/iuc/iuc18/
                                        papers/a2-paper.pdf.

North:1996:UOS

[Nor96]  K. North.  Understanding

ODBC 3.0 standards and OLE DB. DBMS, 9(4):S15–S18, April 1996. CODEN DBMSEO. ISSN 1041-5173.

OConner:1998:CLS

[O’C98] John O’Conner. Comparing locale support in the Java and Win32 platforms. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

ODonnell:1994:PW

[O’D94] Sandra Martin O’Donnell. Programming for the World. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-722190-8. xv + 440 pp. LCCN QA 76.6 O37 1993.

Oetiker:2012:HPA

[Oet12]  Tobi Oetiker.  How Perl

added Unicode support 10 years ago without you noticing it. ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2012/how-perl-added-unicode-

support-10-years-ago-without-
you-noticing-it.

Okutsu:1998:WRJ

[OL98] Masayoshi Okutsu and Norbert Lindenberg. World ready Java — the next step — part 2: Input method framework. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Osborn:2010:ALD

[Osb10] Don Osborn. African languages in a digital age: challenges and opportunities for indigenous language computing. HSRC Press, Cape Town, South Africa, 2010. ISBN 0-7969-2249-7 (paperback). xvi + 150 pp. LCCN PL8005 .O83 2010.

URL http://www.hsrcpress.
ac.za/product.php?cat=1\%
26newtitles=1\%26freedownload=
1\%26productid=2269.

Oudet:1997:MI

[Oud97] Bruno Oudet. Multilingualism on the Internet. Scientific American, 276(3):77–78, March 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Discusses the problems of multilingualism, with brief mention of ISO10646 and Unicode character sets, and of automated language translation systems in current use.

Panel:1998:II

[Pan98a] Panel of speakers. Internationalization of the Internet. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Panel:1998:IIC

[Pan98b] Panel of speakers. Internationalization of the Internet — continued. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Panel:1998:UUPa

[Pan98c] Panel of speakers. Unicode from the users’ perspective. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Panel:1998:UUPb

[Pan98d] Panel of speakers. Unicode from the users’ perspective — continued. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Panel:1998:UJ

[Pan98e] Panel of speakers. Unicode in Japan. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Park:2016:KCC

[Par16] Dongoh Park. The Korean character code: A national controversy, 1987–1995. IEEE Annals of the History of Computing, 38(2):40–53, April/June 2016. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Pigeon:1998:AA

[PB98] Steven Pigeon and Yoshua Bengio. Algorithm alley: Memoryefficient adaptive Huffman coding. Dr. Dobbs Journal, 23 (10):131–132, 134–135, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_
10/Shas/Shas.htm;   http:/
/www.ddj.com/ftp/1998/1998_
10/aa108.txt;   http://www.
ddj.com/ftp/1998/1998_10/
aa108.zip.

Paloque-Berges:2021:RPM

[PB21] Camille Paloque-Bergès. The real “poor man’s Arpanet”? A conversation about Unix networks with Kilnam Chon, godfather of the Asian Internet. Internet Histories, 5(3–4):376– 388, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/
10.1080/24701475.2021.1982167.

Peruginelli:1992:CST

[PBA92] S. Peruginelli, G. Bergamin, and P. Ammendola. Character sets: towards a standard solution. Program, 26(3):215–223, July 1992. CODEN PRGMBD. ISSN 0033-0337.

Pettman:1997:UPS

[PCM^{+}97] I. Pettman, C. Clissman, R. Murray, E. Davidson, J. Hands, O. Sijtsma, A. Noordzij, R. Moulton, S. Shanawa, and J. Darzentas. The UNIverse Project: state-of-the-art of the standards, softwares and systems which will underpin the development. part 1: Z39.50; WWW integration with Z39.50; and Unicode. New Library World, 98(7):267–274, July 1, 1997. CODEN ???? ISSN 0307-4803.

Pohl:2017:BJT

[PDR17] Henning Pohl, Christian Domin, and Michael Rohs. Beyond just text: Semantic emoji similarity modeling to support expressive communication [emoji string]. ACM Transactions on Computer-Human Interaction, 24(1):6:1–6:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Perry:2009:ATU

[Per09] David Perry. Ancient TEX: Using XETEX to support classical and medieval studies. TUGboat, 30(1):12, 2009. ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb30-1/tb94perry.
                                        pdf.

Petzold:1993:MOA

[Pet93a] Charles Petzold. Move over, ASCII! Unicode is here. PC Magazine, 12(18):374–??, October 26, 1993. CODEN PCMGEP. ISSN 0888-8507.

Petzold:1993:UWC

[Pet93b] Charles Petzold. Unicode, wide characters, and C. PC Magazine, 12(19):369–??, November 9, 1993. CODEN PCMGEP. ISSN 0888-8507.

Petzold:1993:VUT

[Pet93c] Charles Petzold. Viewing a Unicode TrueType font under Windows NT. PC Magazine, 12(20): 379–??, November 23, 1993. CODEN PCMGEP. ISSN 0888-8507.

Preston:1997:GSG

[PF97] Holly Hubbard Preston and Udo Flohr. Globalizing software: Global from day one. Byte Magazine, 22(3):97, 98, 100, 102, 104, March 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Unicode.

Pournader:2017:PDU

[PHE17] Roozbeh Pournader, Bob Hallissy, and Lorna Evans. Proposed draft: Unicode Arabic mark ordering algorithm. Technical Report 53, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, September 14, 2017. 13 pp. URL http://www.

unicode.org/reports/tr53.

Phipps:2001:CES

[Phi01] Toby Phipps. Compatibility encoding scheme for UTF-16: 8-bit (CESU-8). Draft Unicode Technical Report 26, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 2001.  URL http:/
/www.unicode.org/unicode/
reports/tr26/tr26-1.

Pike:2003:HU

[Pik03] Rob Pike. The history of UTF-8. Archived e-mail message., April 30, 2003. URL http:

//doc.cat-v.org/bell_labs/
utf-8_history.

Pike:2012:UTB

[Pik12] Rob Pike. UTF-8 turned 20 years old yesterday. Web document., September 6, 2012. URL

http://googleblog.blogspot.
com/2012/02/unicode-over-
60-percent-of-web.html; http:
//www.cl.cam.ac.uk/~mgk25/
ucs/utf-8-history.txt;   https:
//plus.google.com/u/0/101960720994009339267/
posts/Rz1udTvtiMg.  The

author remembers how, two decades ago, he and Ken Thompson worked out the UTF-8 encoding scheme at a diner, eliminating serious deficiencies of earlier compact encodings of Unicode characters.

Plaice:1993:LDL

[Pla93]  John Plaice.  Language-

dependent ligatures. TUGboat, 14(3):271–274, October 1993. ISSN 0896-3207. Describes extensions to TEX to support 16bit character set input, with a particular view to Unicode support.

Plaice:1994:POP

[Pla94] John Plaice. Progress in the Omega project. TUGboat, 15 (3):320–324, September 1994. ISSN 0896-3207. Omega is an extension of TEX for 16-bit character sets.

Plauger:1999:SCCd

[Pla99]  P. J. Plauger.  Standard

C/C++: Unicode files. C/C++ Users Journal, 17(4):??, April 1999. CODEN CCUJEX. ISSN 1075-2838.

Pong:1995:HCF

[PLYZ95] Man-Chi Pong, Fung Fung Lee, R. Yeung, and Yongguang Zhang. Han character font sharing across incompatible bitmap file formats. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 9(2): 145–153, December 1995. CO-DEN CPCLE6. ISSN 0715-9048.

Punchimudiyanse:2017:AFW

[PM17] Malinda Punchimudiyanse and Ravinda Gayan Narendra Meegama. Animation of fingerspelled words and number signs of the Sinhala sign language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):24:1–24:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Poggenpohl:2003:BRL

[Pog03] Sharon Helmer Poggenpohl. Book review: Language Culture Type: International Type Design in the Age of Unicode, John D. Berry, editor. New York: Association Typographique Internationale (ATypI), 2002. Hardbound, 374 pages, extensively illustrated, some signatures in color, $60 Designed by Maxim Zhukov. ISBN 1-932026-01-0. Visible Language, 37(2):233–236, May 2003. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.
com/visiblelanguage/pdf/V37N2_
2003_E.pdf.

Pollack:1991:UCC

[Pol91] Andrew Pollack. Universal computer code due: Rivals join forces to design standard. New York Times, ??(??):D1– D1, D5–D5, February 20, 1991. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.
proquest.com/docview/108790329.

This appears to be first mention of Unicode in this newspaper.

Potter:2005:DUU

[Pot05] B. Potter. Dangerous urls: Unicode & IDN. Network Security, 2005(3):5–6, ???? 2005. ISSN 1353-4858.

Pournader:2002:CU

[Pou02] Roozbeh Pournader. Catching up to Unicode. TUGboat, 23(1):80–85, 2002. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb23-1/
roozbeh.pdf.

Pratley:1996:SUP

[Pra96a] C. Pratley. Shrink-wrapping Unicode. In Unicode Consortium [Uni96c], pages B1– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Pratley:1996:SWU

[Pra96b] C. Pratley. Shrink-wrapping Unicode. In Unicode Consortium [Uni96b], pages B1– ?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Pratley:1997:MOUb

[Pra97a] C. Pratley. Microsoft Office97 and Unicode. In Unicode Consortium [Uni97b], pages C9–

??  ISBN ????  LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Pratley:1997:MOUa

[Pra97b] Chris Pratley. Microsoft Office97 and Unicode. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ????

                                        URL http://www.unicode.
                                        org/unicode/iuc10/program.
                                        html.

Pratley:1998:MOUa

[Pra98a] Christopher Pratley. Microsoft Office2000 and Unicode. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Pratley:1998:MOUb

[Pra98b] Christopher Pratley. Microsoft Office2000 and Unicode — continued. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Pratley:1998:MOUc

[Pra98c] Christopher Pratley. Microsoft Office97 and Unicode. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.

unicode.org/unicode/iuc12/
program.html.

Team:2015:LNI

[Pro15] LaTEX Project Team. LaTEX news, issue 23, October 2015. TUGboat, 36(3):212–213, 2015. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb36-3/
tb114ltnews23.pdf.

Team:2016:LNI

[Pro16] LaTEX Project Team. LaTEX news, issue 24, January 2016. TUGboat, 37(1):13–14, 2016. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb37-1/
tb115ltnews24.pdf.

Pike:1993:HW

[PT93] Rob Pike and Ken Thompson. Hello world. In USENIX Association [USE93], pages 43–50. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84

1993. URL http://plan9.
bell-labs.com/sys/doc/utf.
pdf.

Pike:2000:HWL

[PT00] Rob Pike and Ken Thompson.

Hello world or Kαλημ ´ ϵρα κ oσμϵ or . . .. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA,

2000. URL http://plan9.
bell-labs.com/sys/doc/utf.
html;   http://plan9.bell-
labs.com/sys/doc/utf.pdf;
http://plan9.bell-labs.com/
sys/doc/utf.ps. Originally

appeared, in a slightly different form, in Proc. of the Winter 1993 USENIX Conf., pp. 43–50, San Diego.

Por:2012:UTB

[PWC12] Lip Yee Por, KokSheik Wong, and Kok Onn Chee. UniSpaCh: a text-based data hiding method using Unicode space characters. The Journal of Systems and Software, 85(5): 1075–1082, May 2012. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0164121211003177.

Pong:1992:CCT

[PZ92] Man-Chi Pong and Yongguang Zhang. cxterm: A Chinese terminal emulator for the X Window System. Software—Practice and Experience, 22(10): 809–826, October 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ramati:2022:ABD

[Ram22] Ido Ramati. Aleph-bet, ditsand-dahs, zeros and ones: representing Hebrew in character code. Internet Histories, 6(3):280–297, 2022. CO-DEN ???? ISSN 2470-

1483.  URL http://www.
tandfonline.com/doi/full/
10.1080/24701475.2021.1953757.

Rannenberg:1996:SUU

[Ran96] W. Rannenberg. Supporting Unicode on a UNIX system. In Unicode Consortium [Uni96c], pages C5–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Raner:1997:NDI

[Ran97a] M. Raner. New developments for the implementation of Unicode word processing systems in Java. In Unicode Consortium [Uni97b], pages A13–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Raner:1997:UEC

[Ran97b]  M. Raner.  Universal and

efficient compression of Unicode character streams. In Unicode Consortium [Uni97b], pages C16–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Raner:1997:UJD

[Ran97c] Mirko Raner. Unicode and Java — design and implementation of a Unicode word processing system in Java. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Rauschmayer:2014:SJD

[Rau14] Axel Rauschmayer. Speaking JavaScript: [an in-depth guide for programmers]. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-6503-5. xvii + 437 pp. LCCN QA76.73.J39 R38 2014.

Ryner:1998:JKF

[RC98] Stephen Wood Ryner, Jr. and Nobuko Chikamatsu. Java Kanji Flashcard 500: Kanji, Java and the World Wide Web. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Reed:2017:PIU

[Ree17] Nathan Reed. A programmer’s introduction to Unicode. Web site, March 3, 2017.

                                        URL http://reedbeta.com/
                                        blog/programmers-intro-to-
                                        unicode/.

Resnick:1998:EUI

[Res98] Pete Resnick. Eudora and Unicode: Internationalizing the P.O. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Revanur:1993:WCM

[Rev93] N. Revanur. Win CALIS 2.0 and multi-lingual editor. In Borchardt and Johnson [BJ93], pages 130–131. ISBN ???? LCCN ????

Radzivilovsky:2016:UEM

[RGN16] Pavel Radzivilovsky, Yakov Galka, and Slava Novgorodov.

UTF-8 everywhere: Manifesto. Web site, September 28, 2016. URL http://

utf8everywhere.org/.

Richter:1994:AWN

[Ric94]  Jeffrey Richter.  Advanced

Windows NT: the developer’s guide to the Win32 application programming interface. Microsoft Press, Bellevue, WA, USA, 1994. ISBN 1-55615-567-0. xxxii + 700 pp. LCCN QA76.76.O63R545 1994.

Richter:1995:AWD

[Ric95] Jeffrey Richter. Advanced Windows: the developer’s guide to the Win32 API for Windows NT 3.5 and Windows 95. Microsoft Press, Bellevue, WA, USA, 1995. ISBN 1-55615-677-4. xxxiv + 930 pp. LCCN QA76.76.W56 R52 1995.

Richmond:1998:IUC

[Ric98] Charles M. Richmond. Internationalization of UNIX commands and utilities. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc13/program.
html.

Ritchie:19xx:IHQ

[Ritxx] Dennis Ritchie. An incomplete history of the QED text editor. Report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA,

19xx. URL http://plan9.
bell-labs.com/who/dmr/qed.
html.

Rowley:1997:AIR

[RM97] Chris A. Rowley and Frank Mittelbach. Applicationindependent representation of multilingual text. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ????

                                        URL http://www.unicode.
                                        org/unicode/iuc10/program.
                                        html.

Roshanbin:2016:AIU

[RM16] Narges Roshanbin and James Miller. ADAMAS: Interweaving Unicode and color to enhance CAPTCHA security. Future Generation Computer Systems, 55(??):289–310, February 2016. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0167739X14002386.

Roberts:1998:UIV

[Rob98a] Gary Roberts. Unicode inside: a view from the trenches. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Roberts:1998:UIT

[Rob98b] Gary Roberts. Unicode inside: The Teradata experience. In Unicode Consortium [Uni98b], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Robertson:2010:UML

[Rob10] Will Robertson. Unicode mathematics in LaTEX: Advantages and challenges. TUGboat, 31(2): 211–220, 2010. ISSN 0896-3207.

Robertson:2015:SML

[Rob15] Will Robertson. Single- and multi-letter identifiers in Unicode mathematics. TUGboat, 36(2):106–108, 2015. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb36-2/
tb113robertson.pdf.

Rodgers:2004:IMC

[RSLO04] R. P. Channing Rodgers, Ziying Sherwin, Henk Lamberts, and Inge M. Okkes. ICPC multilingual collaboratory: a Weband Unicode-based system for distributed editing/translating/ viewing of the multilingual international classification of primary care. In Fieschi et al. [FCJ04], pages 425–429. ISBN 1-58603-444-8. Two volumes.

Rublee:1996:AMU

[Rub96] C. Rublee. Acrobat meets Unicode. In Unicode Consortium [Uni96c], pages A2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Saadi:2008:OUV

[Saa08] Z. Saadi. Orthographic Unicode variations in Arabic: a case study of character occurrences in news corpora. In Dilworth B. Parkinson, editor, 21st Symposium on Arabic Linguistics: Perspectives on Arabic Linguistics XXI (Provo, UT), volume 301 of Amsterdam studies in the theory and history of linguistic science. Series IV, Current issues in linguistic theory, pages 1–12. John Benjamins Pub Co., Amsterdam, The Netherlands, 2008. ISBN 90-272-4817-6. ISSN 0304-0763. LCCN ????

Sajib:2019:TBS

[Saj19] Md Qutub Uddin Sajib. Typesetting the Bangla script in Unicode TEX engines — experiences and insights. TUGboat, 40(3):263–269, ???? 2019. CODEN ???? ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb40-3/tb126sajib-
bangla.pdf.

Sargent:1996:UTO

[Sar96] M. Sargent. A Unicode text object model (TOM). In Unicode Consortium [Uni96c], pages B2– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Sargent:1997:UTO

[Sar97] Murray Sargent, III. A Unicode text object model (TOM). In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Sargent:1998:FII

[Sar98a] Murray Sargent, III. A flexible IME implementation. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Sargent:1998:TTE

[Sar98b] Murray Sargent, III. Tips and tricks on editing and displaying Unicode text. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Sargent:2006:UNP

[Sar06]  Murray Sargent III.  Uni-

code nearly plain-text encoding of mathematics. Unicode Technical Note 28, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 12, 2006.

URL http://www.unicode.
org/notes/tn28/;   http://
www.unicode.org/notes/tn28/
tn28-1.html;   http://www.
unicode.org/notes/tn28/UTN28-
PlainTextMath.pdf.

Sasahara:1998:CCJ

[Sas98] Hiroyuki Sasahara. Coincidence and clash of jitai. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Storer:1998:DDC

[SC98] James A. Storer and Martin Cohn, editors. DCC ’98: Data Compression Conference, March 30–April 1, 1998, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. CO-DEN DDCCF9. ISBN 0-8186-8406-2, 0-8186-8408-9 (microfiche). ISSN 1068-0314. LCCN QA76.9.D33 D37 1998. IEEE Computer Society Press order number PR08406. IEEE Order Plan catalog number 98TB100225.

Scharfglass:1997:KA

[Sch97] Ira Scharfglass. Keynote address. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Scherer:2002:CEA

[Sch02] Markus W. Scherer. Canonical equivalence in applications. Unicode Technical Note 5, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, July 24, 2002.

URL http://www.unicode.
org/notes/tn5/.

Scherer:2004:UP

[Sch04] Markus Scherer. UTF-16 for processing. Unicode Technical Note 12, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, January 13,
2004.  URL http://www.
unicode.org/notes/tn12/tn12-
1.html;   http://www.unicode.
org/notes/tn13/.

Spitz:1995:IAP

[SD95] A. Lawrence Spitz and Andreas Dengel, editors. International Association for Pattern Recognition Workshop on Document Analysis Systems, volume 14 of Series in machine perception and artificial intelligence. World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1995. ISBN 981-02-2122-3, 981-02-2122-3. LCCN TK7882.P3I55 1994.

Scherer:2002:BMC

[SD02] Markus W. Scherer and Mark Davis. BOCU-1: MIMEcompatible Unicode compression. Unicode Technical Note 6, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, August 9, 2002.

URL http://www.unicode.
org/notes/tn6/. Last revision:

2006-02-04.

Scherer:2006:PDU

[SD06] Markus W. Scherer and Mark Davis. Proposed draft Unicode technical standard #40: BOCU-1: MIME-compatible Unicode compression. Technical report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, February 3, 2006. URL

                                        http://www.unicode.org/reports/
                                        tr40/tr40-1.html.

Seeger:2013:SUT

[See13] Ulrich Seeger. Semicode — ein Unicode-Tastaturtreiber für Mac OS X. (German) [Semicode — a Unicode keyboard driver for Mac OS X]. Seminar für Sprachen und Kulturen des Vorderen Orients, Semitistik, Universität Heidelberg, Heidelberg, Germany, 2013.

SHARE:1993:PSE

[SHA93] Proceedings SHARE Europe Spring Meeting. Distributed Applications (Hamburg, Germany, 19–22 April 1993). SHARE Europe, Carouge/ Geneva, Switzerland, 1993. ISBN ???? LCCN ????

Shimada:1998:IUS

[Shi98a] Shigeru Shimada. IBM AS/ 400 Unicode support. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.

unicode.org/unicode/iuc12/
program.html.

Shiratori:1998:CUH

[Shi98b] Takaaki Shiratori. Considerations for Unified Han character search operations, input, and output, in multi-lingual environment. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Sidhu:2005:GCN

[Sid05] Sukhjinder Sidhu. Gurmukhi character names. Unicode Technical Note 25, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, November 15, 2005.

URL http://www.unicode.
org/notes/tn15/;   http://
www.unicode.org/notes/tn25/
GurmukhiNames.pdf;   http:
//www.unicode.org/notes/tn25/
tn25-1.html.

Sidhu:2006:IUC

[Sid06] Sukhjinder Sidhu. ISCII to Unicode conversion issues for Gurmukhi. Unicode Technical Note 30, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 22, 2006.

URL http://www.unicode.

                                        org/notes/tn30/;   http://
                                        www.unicode.org/notes/tn30/
                                        tn30-1.html.

Sprung:2000:TSC

[SJ00] Robert C. (Robert Charles) Sprung and Simone Jaroniec, editors. Translating Into Success: Cutting-Edge Strategies for Going Multilingual in a Global Age, volume 11 of American Translators Association scholarly monograph series. John Benjamins, Amsterdam, The Netherlands and Philadelphia, PA, USA, 2000. ISBN 90-272-3187-7 (paperback), 1-55619-630-X. xxii + 239 pp. LCCN P306.A1 A53 v.11 2000.

Schulz:2018:BGF

[SK18] Herbert Schulz and Richard Koch. A beginner’s guide to file encoding and TEXShop. TUGboat, 39(3):173–176, ???? 2018. CODEN ???? ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb39-3/
                                        tb123schulz-encoding.pdf.

Smura:1987:TNB

[SP87] Edwin J. Smura and Archie D. Provan. Toward a new beginning: the development of a standard for font and character encoding to control electronic document interchange. IEEE Transactions on Professional Communication, PC-30 (4):259–264, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Spring:1996:SRM

[Spr96] Michael B. Spring. Standards:

Reference model for data interchange standards. Computer, 29(8):87–88, August 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses cost of conversion from ASCII to Unicode.

Smith:1998:GAT

[SS98] William Smith and Don Stinchfield. Global access to technical documentation. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc12/program.
html.

Segbert:1997:LAC

[SSB97] M. Segbert, K. Steinwachs, and P. Burnett, editors. Library automation in Central and Eastern Europe: International conference — April 1996, Budapest. European Commission, ????, 1997. ISBN 92-828-0305-8. LCCN ????

Seethalakshmi:2005:OCR

[SSB^{+}05] R. Seethalakshmi, T. R.

Sreeranjani, T. Balachandar, A. Singh, M. Singh, R. Ratan, and S. Kumar. Optical character recognition for printed Tamil text using Unicode. In Proceedings of the First International Congress, ICUDL2005, pages 1297–1305. Zhejiang University Press, ????, 2005. ISBN ???? ISSN 1009-3095. LCCN ???? Published in Journal: Zhejiang University Science, 6A, no. 11, (2005).

Steele:1996:TCF

[Ste96] Guy L. Steele, Jr. Trends & controversies: The future of Lisp: The Java programming language. IEEE expert: intelligent systems and their applications, 11(1):10–16, February 1996. CODEN IEEXE7. ISSN 0885-9000. Mentions Unicode support in Java.

Syropoulos:2003:DTU

[STS03] Apostolos Syropoulos, Antonis Tsolomitis, and Nick Sofroniou. Digital typography using LaTEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95217-9. xxix + 510 pp. LCCN Z253.4.L38 S97 2003.

Sugimoto:1996:CTE

[Sug96] S. Sugimoto. Component technologies and environment for digital libraries. Joho-Shori (J. Information Processing Soc. Japan), 37(9):820–825, September 1996. CODEN JOSHA4. ISSN 0447-8053.

Suignard:1998:WTH

[Sui98] Michel Suignard. Worldwide typography and how to apply JIS X-4051-1995 to Unicode. In Unicode Consortium [Uni98a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Suignard:2001:CTE

[Sui01] Michel Suignard. CSS3 text extensions. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/c5.pdf.

Sun:1996:DIC

[Sun96] Y. Sun. Design and implementation of the CJK unified character set on UNIX system. In Unicode Consortium [Uni96c], pages B6–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Sun:1998:IHS

[Sun98a] Sam X. Sun. Internationalization of the handle system — a persistent global name service. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Sun:1998:ALS

[Sun98b] Sun Microsystems, Inc. Asianlanguage support in the Solaris operating environment. White paper, Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, October 1998.

URL http://www.sun.com/
software/white-papers/wp-
asianlanguage/.

Sung:1998:ULS

[Sun98c] Ienup Sung. Unicode locale support in Solaris. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc12/program.html.

Sun:19xx:AAD

[Sunxxa] Sun Microsystems, Inc. Asian Application Developer’s Guide. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 19xx. ISBN ???? ???? pp. LCCN ????

                                        URL http://docs.sun.com/;
                                         http://www.sun.com/globalization/

.

Sun:19xx:SIG

[Sunxxb] Sun Microsystems, Inc. Solaris Internationalization Guide For Developers. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 19xx. ISBN ???? LCCN ???? URL http:

                                        //docs.sun.com/;   http://
                                        www.sun.com/globalization/

.

Sun:2000:ICU

[Sun00a] Sun Microsystems, Inc. International components for Unicode. World-Wide Web document, 2000. URL http://oss.

                                        software.ibm.com/icu/.

Sun:2000:ILE

[Sun00b] Sun Microsystems, Inc. International Language Environments Guide (Solaris 8). Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, February 2000. ISBN ???? 243 (est.) pp. LCCN ???? US$64.00. URL http:

//docs.sun.com/;   http://
www.sun.com/globalization/

. Product number DM10002730. Suzuki:1998:IEI

[Suz98] Takao Suzuki. Internet Explorer internationalization. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Swinnen:2009:APA

[Swi09] Gérard Swinnen. Apprendre á programmer avec Python: objet, multithreading, événements, bases de données, programmation web, programmation réseau, Unicode. Eyrolles, Paris, France, 2009. xviii + 341 pp. LCCN ????

Swoyer:2005:SPC

[Swo05] S. Swoyer. Staying power!

COBOL is not about to go away. COBOL has become UNICODE-compliant, object oriented and XML friendly. Application Development Trends, 12(2):14–18, 2005. ISSN 1073-9564.

Tan:1994:PIC

[Tan94]  S. T. Tan, editor.  Pro-

ceedings. International Conference on Chinese Computing ’94 (ICCC94). The Latest Technological Advancement and ’94 (ICCC94). The Latest Technological Advancement and Application. National University of Singapore, Singapore, 1994. ISBN 981-00-5648-6. LCCN ????

Tang:1996:IUN

[Tan96] F. Y.-F. Tang. Implementing Unicode in Netscape Navigator. In Unicode Consortium [Uni96c], pages A1– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

                                        //www.unicode.org/unicode/
                                        iuc9conf/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Tang:1997:UIN

[Tan97] Frank Yung-Fong Tang. Unicode and internationalization in Netscape Communicator. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Tang:1998:MMM

[Tan98] Frank Yung-Fong Tang. Multilingual Mozilla — Mozilla language enabling project. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Tanaka:2013:UUV

[Tan13] Takuji Tanaka. upTEX — Unicode version of pTEX with CJK extensions. TUGboat, 34(3): 285–288, 2013. ISSN 0896-3207.

Tauber:2016:PPP

[Tau16] J. K. Tauber. pyuca: a Python implementation of the Unicode Collation Algorithm. Journal of Open Source Software, 1(1): 21:1, May 2016. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/10.
21105/joss.00021.

Taylor:1992:GS

[Tay92]  Dave Taylor.  Global Soft-

ware: developing applications for the international market. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97706-6, 3-540-97706-6. xvi + 319 pp. LCCN QA76.76.D47T39 1992.

Tenney:1997:AUR

[Ten97a] M. Tenney. Apple’s Unicode road map. In Unicode Consortium [Uni97b], pages C4–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Tenney:1997:FUS

[Ten97b] Merle Tenney. The future of Unicode support on the Macintosh. In Unicode Consortium [Uni97a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
program.html.

Tenney:1998:BUA

[Ten98] Merle Tenney. Building a Unicode application on the Mac OS. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Texin:1998:MRDa

[Tex98a] Tex Texin. Migration of a 4GL and relational database to Unicode. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Texin:1998:MRDb

[Tex98b] Tex Texin. Migration of a 4GL and relational database to Unicode. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Texin:2001:UID

[Tex01] Tex Texin. Understanding ISO 8601 date and time representation formats. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/a13.ppt.

UC:1990:U

[The90] The Unicode Consortium. Unicode 1.0. The Unicode Consortium, Mountain View, CA, USA, December 1990. ISBN ?? various pp. LCCN ?? Draft standard. Final review document.

Unicode:1991:USW

[The91] The Unicode Consortium. The Unicode Standard: Worldwide Character Encoding. Version 1.0. Volumes 1 and 2. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-56788-1 (paperback, vol. 1), 0-201-60845-6 (paperback, vol. 2). LCCN QA268 .U55 1991, Z103 .U6 1991. US$39.95 (paperback), US$32.95 (hardcover).

Unicode:1996:USV

[The96] The Unicode Consortium. The Unicode Standard, Version 2.0. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48345-9 (paperback), 0-201-96623-9 (CD-ROM). various pp. LCCN QA268.U56 1996. US$54.95.

Unicode:2000:USV

[The00] The Unicode Consortium. The Unicode Standard, Version 3.0. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61633-5. xxx + 1040 pp. LCCN QA268.U56 1996. US$50.00. URL http://www.

unicode.org/book/;   http:
//www.unicode.org/unicode/
standard/versions/. Includes

CD-ROM. The principal authors and editors of The Unicode Standard, Version 3.0 are Joan Aliprand, Julie Allen, Joe Becker, Mark Davis, Michael Everson, Asmus Freytag, John Jenkins, Mike Ksar, Rick Mc-Gowan, Lisa Moore, Michel Suignard, and Ken Whistler.

Unicode:2003:USV

[The03] The Unicode Consortium. The Unicode Standard, Version 4.0. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-18578-1. xxxviii + 1462 pp. LCCN QA268 .U545 2004.

                                        URL http://www.unicode.
                                        org/versions/Unicode4.0.0/

. Includes CD-ROM.

Unicode:2006:UCD

[The06a] The Unicode Consortium. Unicode character database: Derived property data. Online report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, July 14, 2006. URL http:

                                        //www.unicode.org/Public/
                                        UNIDATA/DerivedAge.txt.

Unicode:2006:USV

[The06b] The Unicode Consortium. The Unicode Standard, Version 5.0. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-48091-0 (hardcover). l + 1417 pp. LCCN QA268 .U545 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0617/2006023526.html;
                                         http://www.unicode.org/versions/
                                        beta.html;   http://www.unicode.
                                        org/versions/Unicode5.0.0/

. Includes CD-ROM.

Unicode:2010:USV

[The10] The Unicode Consortium. The Unicode Standard, Version 5.0.

Tsinghua University Press, Tsinghua, China, 2010. ISBN 7-302-21244-9 (hardcover). 1358 pp. LCCN ???? URL http:/

/www.amazon.cn/gp/product/

B00328IJ46. Includes CD-ROM.

Unicode:2016:CUC

[The16a] The Unicode Consortium.

CLDR 30 — Unicode Common Locale Data Repository. Web document, October 5, 2016.

URL http://cldr.unicode.
org/index/downloads/cldr-

30. Unicode CLDR 30 provides an update to the key building blocks for software supporting the world’s languages. This data is used by all major software systems for their software internationalization and localization, adapting software to the conventions of different languages for such common software tasks.

Unicode:2016:USVa

[The16b] The Unicode Consortium. The Unicode(R) Standard: Version 9.0 — Core Specification. The Unicode Consortium, Mountain View, CA, USA, 2016. ISBN 1-936213-13-3. xxxvi + 998 + 1 pp. LCCN ???? URL http:/

/www.unicode.org/versions/
Unicode9.0.0/.

Unicode:2016:USVb

[The16c] The Unicode Consortium. The Unicode(R) Standard: Version 9.0 — Core Specification, volume 1. Lulu.com, ????, 2016. 474 pp. LCCN ????

                                        URL http://www.lulu.com/
                                        spotlight/unicode;   http:
                                        //www.unicode.org/versions/
                                        Unicode9.0.0/. Does not con-

tain Unicode code charts; they are available on the Web.

Unicode:2016:USVc

[The16d] The Unicode Consortium. The Unicode(R) Standard: Version 9.0 — Core Specification, volume 2. Lulu.com, ????, 2016. 572 pp. LCCN ????

                                        URL http://www.lulu.com/
                                        spotlight/unicode;   http:
                                        //www.unicode.org/versions/
                                        Unicode9.0.0/. Does not con-

tain Unicode code charts; they are available on the Web.

Unicode:2021:CRN

[The21a] The Unicode Consortium. CLDR 39 release note. Web document, April 7, 2021.

                                        URL http://cldr.unicode.
                                        org/index/downloads/cldr-
                                        39.

Unicode:2021:IIC

[The21b] The Unicode Consortium. ICU 69: International components for Unicode. Web document, April 7, 2021. URL http://

                                        site.icu-project.org/download/
                                        69.

Unicode:2021:U

[The21c] The Unicode Consortium. Unicode 14.0.0. Web document, September 14, 2021.

                                        URL http://www.unicode.
                                        org/versions/Unicode14.0.
                                        0/;   https://home.unicode.

org/announcing-the-unicode-
standard-version-14-0/.

Unicode:2021:USAa

[The21d] The Unicode Consortium. Unicode standard annex #14: Unicode line breaking algorithm. Web document, August 23, 2021. URL http:

//www.unicode.org/reports/
tr14/tr14-47.html.

Unicode:2021:USAb

[The21e] The Unicode Consortium. Unicode standard annex #29: Unicode text segmentation. Web document, August 24, 2021.

URL https://www.unicode.
org/reports/tr29/tr29-39.
html.

Unicode:2021:USAc

[The21f] The Unicode Consortium. Unicode standard annex #31: Unicode identifier and pattern syntax. Web document, August

26, 2021. URL http://www.
unicode.org/reports/tr31/
tr31-35.html.

Unicode:2021:USAd

[The21g] The Unicode Consortium. Unicode standard annex #38: Unicode Han database (Unihan). Web document, August 26,

2021.  URL https://www.
unicode.org/reports/tr38/
tr38-31.html.

Unicode:2021:USAe

[The21h] The Unicode Consortium. Unicode standard annex #45: U-source ideographs. Web document, August 26, 2021.

URL http://www.unicode.
org/reports/tr45/tr45-25.
html.

Consortium:2021:USVa

[The21i] The Unicode Consortium. The Unicode Standard, Version 14.0, Volume 1. lulu.com, ????, 2021. ISBN 1-936213-30-3. 496 pp. URL https:

//www.lulu.com/en/us/shop/
unicode-consortium/the-unicode-
standard-version-140-volume-
1/paperback/product-p8674j.
html.

Consortium:2021:USVb

[The21j] The Unicode Consortium. The Unicode Standard, Version 14.0, Volume 2. lulu.com, ????, 2021. ISBN 1-936213-31-1. 582 pp. URL https:

//www.lulu.com/en/us/shop/
unicode-consortium/the-unicode-
standard-version-140-volume-
1/paperback/product-p8674j.
html.

Unicode:2021:UTSa

[The21k] The Unicode Consortium. Unicode technical standard #10: Unicode collation algorithm. Web document, August 27,

2021.  URL http://www.
unicode.org/reports/tr10/
tr10-45.html.

Unicode:2021:UTSb

[The21l] The Unicode Consortium. Unicode technical standard #39: Unicode security mechanisms. Web document, August 25,

2021.  URL http://www.

unicode.org/reports/tr39/
tr39-24.html.

Unicode:2021:UTSc

[The21m] The Unicode Consortium. Unicode technical standard #46: Unicode IDNA compatibility processing. Web document, August 24, 2021. URL https:

//www.unicode.org/reports/
tr46/tr46-27.html.

Tivig:1997:UIH

[Tiv97] G. Tivig. Unicode implementation in the HP PicoVue patient monitor. In Unicode Consortium [Uni97b], pages C8–

??  ISBN ????  LCCN

????  URL http://www.
unicode.org/unicode/iuc11/
program.html.

Tang:1997:ICN

[TJ97] F. Y.-F. Tang and R. S.

Jung. International challenges for Netscape Communicator. In Unicode Consortium [Uni97b], pages C10–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Toledo:2001:FSD

[TK01] Sivan Toledo and Lars Knoll.

Font subsetting and downloading in the PostScript printer driver of Qt/X11. In USENIX [USE01], page ?? ISBN ????

LCCN ???? URL http://www.
linuxshowcase.org/toledo.
html.

Topping:2001:MDN

[Top01] Suzanne Topping. The multilingual domain name race: On your mark get set WAIT! World-Wide Web document, December 2001. URL http://

www-106.ibm.com/developerworks/
library/u-domains.html.

Torri:1996:ETS

[Tor96] K. Torri. Entire translation services. In Unicode Consortium [Uni96c], pages C6– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Tresman:1994:MPC

[Tre94] Ian Tresman. The multilingual PC directory: a guide to multilingual and foreign language products for IBM PCs and compatibles. Knowledge Computing, 9 Ashdown Drive, Borehamwood, Herts WD6 4LZ, UK, 1994. ISBN 1-873091-02-8, 0-87309-103-5. 256 pp. LCCN QA76.8.I2594 T735 1994. UK£35.00. URL http:/

/www.knowledge.co.uk/xxx/.

Tuthill:1996:CWS

[TS96] Bill Tuthill and David Smallberg. Creating Worldwide Software: Solaris International Developer’s Guide. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-494493-3. xxiii + 382 pp. LCCN QA76.76.O63 T888 1997. US$65.00.

Uren:1993:SIL

[UHP93] Emmanuel Uren, Robert Howard, and Tiziana Perinotti. Software Internationalization and Localization: an Introduction. Van Nostrand Reinhold, New York, NY, USA, 1993. ISBN 0-442-01498-8. xx + 300 pp. LCCN QA76.754.U84 1993. UK£32.00.

UI:1992:UIW

[UIW92] Proceedings of the Second Unicode Implementors Workshop, Merrimack, NH. The Unicode Consortium, Mountain View, CA, USA, 1992. ISBN ?? LCCN ??

Ullrich:2021:RWS

[Ull21] Torsten Ullrich. Real-world string comparison: How to handle Unicode sequences correctly. ACM Queue: Tomorrow’s Computing Today, 19(3):107–116, June 2021. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475965.3478522.

Umamaheswaran:1998:EFU

[Uma98] V. S. Umamaheswaran. EBCDICfriendly UCS transformation format. Unicode Technical Report 16, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, February

18, 1998. URL http:/

/www.unicode.org/unicode/
reports/techreports.html;
 http://www.unicode.org/unicode/
reports/tr16/. In progress.

Umamaheswaran:2002:UE

[Uma02] V. S. Umamaheswaran. UTF-EBCDIC. Unicode Technical Report 16, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 16, 2002.

URL http://www.unicode.
org/reports/tr16/;   http:
//www.unicode.org/unicode/
reports/tr16/tr16-8.html.

Unicode:1991:USV

[Uni91] Unicode, Inc. The Unicode Standard, Version 1.1. Unicode Technical Report 4, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 1991. Superseded by current version of the Unicode Standard.

Unicode:1992:EP

[Uni92]  Unicode, Inc.  Exploratory

proposals. Unicode Technical Report 3, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 1992.  URL http:/
/www.unicode.org/unicode/
reports/tr3.html.

UC:1994:PUIa

[Uni94a] Unicode Consortium, editor.

Proceedings of the Unicode Implementors Workshop 5, Tokyo, Japan, April 11 and 12, 1994. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, 1994.

UC:1994:PUIb

[Uni94b] Unicode Consortium, editor.

Proceedings of the Unicode Implementors Workshop 6, Santa Clara, California, September 8–9, 1994. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 1994.

UI:1994:UV

[Uni94c] The Unicode Consortium, P.O.

Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org. Unicode Version 1.1, 1994. US$10.00.

UC:1995:SIU

[Uni95a] Unicode Consortium, editor.

Seventh International Unicode Conference, San Jose, California, September 14–15, 1995: proceedings. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 1995. ISBN ???? LCCN QA268 .I56 1995. Two volumes.

Unicode:1995:USW

[Uni95b] Unicode Staff. Unicode Standard: Worldwide Character Encoding, Version 1.1. Addison-Wesley, Reading, MA, USA, second edition, 1995. ISBN 0-201-48345-9. ca. 950 pp. LCCN QA268.U56 1996. US$59.25. URL http:

//www.unicode.org/unicode/
uni2book/u2.html.

UC:1996:ASI

[Uni96a] Unicode Consortium, editor. Asia, software + the Internet: going global with Unicode: Eighth International Unicode/ISO 10646 Conference, Hong Kong, April 18– 19, 1996: proceedings. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 1996. ISBN ???? LCCN QA268 .I56 1996.

UC:1996:PCT

[Uni96b] Unicode Consortium, editor. Pre-conference tutorials proceedings: Software development + the Internet: going global with Unicode: Ninth International Unicode Conference, San Jose, CA, September 4–6 1996. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1996. ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

UC:1996:SDI

[Uni96c] Unicode Consortium, editor. Software Development + the Internet: Going Global with Unicode: Ninth International Unicode Conference, 4– 6 September, San Jose, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 1996. ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:

                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

UC:1997:ESI

[Uni97a] Unicode Consortium, editor. Europe, Software + the Internet: Going Global with Unicode: Tenth International Unicode Conference, March 10–12, 1997, Mainz, Germany. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 1997. ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        ;   http://www.unicode.org/
                                        unicode/uni2book/u2ord.html.

UC:1997:SDI

[Uni97b] Unicode Consortium, editor. Software Development + the Internet: Going Global with Unicode: The Eleventh International Unicode Conference, September 2–5, 1997, San Jose, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 1997. ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

UC:1998:ASI

[Uni98a] Unicode Consortium, editor. Asia, Software + the Internet:

Going Global with Unicode: The Twelfth International Unicode/ ISO10646 Conference (IUC12) April 8–10, 1998, Tokyo, Japan. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1998. ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

UC:1998:TIU

[Uni98b] Unicode Consortium, editor.

Thirteenth International Unicode Conference: Software + the Internet: Going Global with Unicode (R), September 8–11, 1998, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1998. ISBN ????

LCCN ???? URL http://www.
reuters.com/unicode/iuc13;
 http://www.unicode.org.

UC:1999:FIUb

[Uni99a] Unicode Consortium, editor.

Fifteenth International Unicode Conference: Unicode and the Web: the Global Connection, August 30 – September 2, 1999, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 1999. ISBN ????

                                        LCCN ???? URL http://www.
                                        reuters.com/unicode/iuc15;
                                         http://www.unicode.org.

UC:1999:FIUa

[Uni99b] Unicode Consortium, editor. Fourteenth International Unicode Conference: Unicode and the Web: the Global Connection, March 22–25, 1999, Boston, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, 1999. ISBN ???? LCCN

                                        ????  URL http://www.
                                        reuters.com/unicode/iuc14;
                                         http://www.unicode.org.

Unicode:19xx:BAR

[Unixxa] Unicode, Inc. The bidirectional algorithm reference implementation. Unicode Technical Report 9, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.
                                        org, 19xx.  URL http:/
                                        /www.unicode.org/unicode/
                                        reports/techreports.html.

In progress.

Unicode:19xx:FCD

[Unixxb] Unicode, Inc. Support for interlinear annotations. Unicode Technical Report 12, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.
org, 19xx.  URL http:/
/www.unicode.org/unicode/
reports/techreports.html.

In progress.

UC:2000:SIUb

[Uni00a] Unicode Consortium, editor.

Seventeenth International Unicode Conference (IUC17) Unicode and the Web: the Global Connection, September 5–8, 2000, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 2000. ISBN ????

LCCN ???? URL http://www.
unicode.org/iuc/iuc17.

UC:2000:SIUa

[Uni00b] Unicode Consortium, editor.

Sixteenth International Unicode Conference (IUC16) Unicode and the Web: the Global Connection, March 27–30, 2000, Amsterdam, Holland. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 2000. ISBN ???? LCCN

????  URL http://www.
unicode.org/iuc/iuc16.

UC:2001:EIU

[Uni01a] Unicode Consortium, editor.

Eighteenth International Unicode Conference (IUC18) Unicode and the Web: the Global Connection, April 24–27, 2001, Hong Kong. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 2001. ISBN ???? LCCN

????  URL http://www.
unicode.org/iuc/iuc18.

Unicode:2001:NIU

[Uni01b] Unicode Consortium, editor. Nineteenth International Unicode Conference (IUC19) Unicode and the Web: the Global Connection, September 10–14, 2001, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 2001. ISBN ???? LCCN

????  URL http://www.
unicode.org/iuc/iuc19.

Unicode:2002:TIU

[Uni02a] Unicode Consortium, editor. Twentieth International Unicode Conference (IUC20) Unicode and the Web: the Global Connection, January 28 – February 1, 2002, Washington, DC, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 2002. ISBN ????

LCCN ???? URL http://www.
unicode.org/iuc/iuc20.

Unicode:2002:TFI

[Uni02b] Unicode Consortium, editor.

Twenty-first International Unicode Conference (IUC21) Unicode, Localization and the Web: The Global Connection, May 14–17, 2002 Dublin, Ireland. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 2002. ISBN ????

LCCN ???? URL http://www.
unicode.org/iuc/iuc21.

Unicode:2002:TSI

[Uni02c] Unicode Consortium, editor.

Twenty-second International Unicode Conference (IUC22) Unicode and the Web: Evolution or Revolution? September 9–13, 2002, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 2002. ISBN ????

LCCN ???? URL http://www.
unicode.org/iuc/iuc22.

Unicode:2002:U

[Uni02d]  Unicode Consortium.  Uni-

code 3.2. Unicode Technical Report 28, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, March 27, 2002.

                                        URL http://www.unicode.
                                        org/unicode/reports/tr28/.

Unicode:2003:TTIb

[Uni03a] Unicode Consortium, editor. Twenty-fourth Internationalization and Unicode Conference (IUC24) Unicode, Internationalization, the Web: Powering Global Business, September 3– 5, 2003, Atlanta, Georgia, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 2003. ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/iuc/iuc24.

Unicode:2003:TTIa

[Uni03b] Unicode Consortium, editor. Twenty-third Internationalization and Unicode Conference (IUC23) Unicode, Internationalization, the Web: The Global Connection Week of March 24– 26, 2003 Prague, Czech Republic. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 2003. ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/iuc/iuc23.

Unicode:2004:TFI

[Uni04a] Unicode Consortium, editor. Twenty-fifth Internationalization and Unicode Conference (IUC25). Unicode in Government: Building a Multilingual Infrastructure, March 31– April 2, 2004, Washington, DC, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 2004. ISBN ???? LCCN ????

Unicode:2004:TSI

[Uni04b] Unicode Consortium, editor.

Twenty-sixth Internationalization and Unicode Conference (IUC26). Unicode and Internationalization: Best Practices in the Networked World, San Jose, California, USA, September 7–10, 2004. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 2004. ISBN ???? LCCN ????

Unicode:2005:TSIb

[Uni05a] Unicode Consortium, editor.

Twenty-eight Internationalization and Unicode Conference (IUC28). Multilingual Challenges and Solutions for 2006. September 7–9, 2005, Orlando, Florida, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 2005. ISBN ????

LCCN ???? URL http://

                                        www.unicodeconference.org/
                                        iuc28-pr.

Unicode:2005:TSIa

[Uni05b] Unicode Consortium, editor. Twenty-seventh Internationalization and Unicode Conference (IUC27). Unicode, Cultural Diversity, and Multilingual Computing, April 6–8, 2005, Berlin, Germany. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, 2005. ISBN ???? LCCN ????

Unicode:2005:U

[Uni05c] Unicode Consortium. Unicode 4.1.0. Technical report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, March 2005. URL http:/

                                        /www.unicode.org/versions/
                                        Unicode4.1.0/.

Unicode:2006:TNIa

[Uni06] Unicode Consortium, editor. Twenty-ninth Internationalization and Unicode Conference (IUC27). March 6–8, 2006, San Francisco, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 2006. ISBN ????

LCCN ???? URL http://

www.unicodeconference.org/
iuc29-pr.

Unicode:2003:UCH

[Uni07] Unicode. The Unicode Consortium home page, 2003, 2007.

URL http://unicode.org/.

Unicode:2011:USV

[Uni11] Unicode Consortium, editor.

The Unicode Standard, Version 6.0 — Core Specification. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, 2011.  ISBN

1-936213-01-X. ???? pp. LCCN QA268.U545 2011.

URL http://www.unicode.
org/versions/Unicode6.0.0/

.

Unicode:2012:USV

[Uni12] Unicode Consortium, editor.

The Unicode Standard, Version 6.1 — Core Specification. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 2012. ISBN ???? 692 (est.) pp. LCCN ???? US$15.96.

URL http://www.lulu.com/
content/paperback-book/the-
unicode-standard-version-
61---core-specification/12828817;
 http://www.lulu.com/shop/
unicode-consortium/the-unicode-
standard-version-61-core-

                                        specification/paperback/product-
                                        20082926.html;   http://www.
                                        unicode.org/versions/Unicode6.
                                        1.0/.

Unicode:2020:USV

[Uni20] Unicode Consortium. The Unicode Standard, Version 13.0. Lulu Press, Inc., ????, 2020. 488 (vol. 1), 572 (vol. 2) pp.

                                        URL https://www.lulu.com/
                                        search?q=Unicode. Two vol-

umes (chapters 1–11 and 12– 24). Code charts, Standard Annexes, and the Unicode Character Database are not included, but are available electronically on the Web.

Unsworth:1996:SHC

[Uns96] J. Unsworth. Software for humanities computing: Image annotation, Unicode, and SGML packages from the Institute for Advanced Technology in the Humanities. In Anonymous [Ano96a], pages 297–?? ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

USENIX:1993:PWU

[USE93] USENIX Association, editor. Proceedings of the Winter 1993 USENIX Conference: January 25–29, 1993, San Diego, California, USA. USENIX, San Francisco, CA, USA, 1993. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

USENIX:2001:PAL

[USE01] USENIX, editor. Proceedings of the 5th Annual Linux Showcase and Conference, November 5–10, 2001, Oakland, CA. USENIX, San Francisco, CA, USA, 2001. ISBN

????  LCCN ????  URL

http://www.linuxshowcase.
org/tech.html.

Umamaheswaran:1998:USE

[UY98] V. S. Umamaheswaran and Jianping Yang. Unicode support in EBCDIC based system. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Vacca:1991:UBA

[Vac91] John R. Vacca. Unicode breaks the ASCII barrier. Datamation, 37(15):55–??, August 1991. CO-DEN DTMNAT. ISSN 0011-6963.

VanStee:1991:PIM

[Van91] J. G. Van Stee. Programming implications of multibyte code pages. In Proceedings. SHARE Europe Anniversary Meeting. Amsterdam, The Netherlands, 7–11 October, 1991, volume 1110, pages 719– 740. SHARE Europe (SEAS), Geneva, Switzerland, 1991. Two volumes.

VanCamp:1994:USG

[Van94]  David Van Camp.  Uni-

code and software globalization. Dr. Dobbs Journal, 19 (3):46, 48–50, March 1994. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        davidvancamp.com/Articles_
                                        files/SoftwareGlobalization.
                                        DOC.

vanderPoel:1998:EWN

[van98] Erik van der Poel. Easier Web navigation. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc13/program.html.

Vieth:2008:DWN

[Vie08] Ulrik Vieth. Do we need a ‘Cork’ math font encoding? TUGboat, 29(3):426–434, 2008. ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb29-3/
                                        tb93vieth.pdf.

Wolf:2000:SCS

[W^{+}00] Misha Wolf et al. A standard compression scheme for Unicode. Unicode Technical Report 6, The Unicode Consortium, Mountain View, CA, USA, 2000. ?? pp.

                                        URL http://unicode.org/
                                        unicode/reports/tr6/index.
                                        html.

Whistler:1998:RLT

[WA98a] K. Whistler and G. Adams. RFC 2482: Language tagging in Unicode plain text, December 1998. URL ftp://ftp.

                                        internic.net/rfc/rfc2482.
                                        txt;   ftp://ftp.math.utah.
                                        edu/pub/rfc/rfc2482.txt. Sta-

tus: INFORMATIONAL.

Whistler:1998:PCL

[WA98b] Ken Whistler and Glenn Adams. Plane 14 characters for language tags. Unicode Technical Report 7, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, May 28, 1998. URL http:

//www.unicode.org/unicode/
reports/techreports.html;
http://www.unicode.org/unicode/
reports/tr7.html. In progress.

RFC2482

[WA99] K. Whistler and G. Adams.

RFC 2482: Language tagging in Unicode plain text, January 1999. URL ftp://ftp.

internic.net/rfc/rfc2482.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc2482.txt. Sta-

tus: INFORMATIONAL.

Wagner:2006:BSH

[Wag06] Zdeněk Wagner. Babel speaks Hindi. TUGboat, 27(2):176– 180, 2006. ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb27-2/tb87wagner.
pdf.

Wang:1997:UMS

[Wan97] R. Wang. Unicode and multilingual search and retrieval. In Unicode Consortium [Uni97b], pages A9–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Whistler:1998:CEM

[WD98a] Ken Whistler and Mark Davis.

Character encoding model. Unicode Technical Report 16, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org,  October

14, 1998. URL http:/

/www.unicode.org/unicode/
reports/techreports.html;
 http://www.unicode.org/unicode/
reports/tr17/.

Whistler:1998:UC

[WD98b] Ken Whistler and Mark Davis. Unicode collation. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ????

URL http://www.unicode.
org/unicode/iuc13/program.
html.

Whistler:1999:CEM

[WD99] Ken Whistler and Mark Davis. Character encoding model. Unicode Technical Report 17, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, November 23, 1999.

URL http://www.unicode.
org/unicode/reports/tr17/.

Wedel:1996:LAB

[Wed96] G. Wedel. Linking Arabic biographical data to reconstruct Arabic history: proposals to design a biographical information system. In Proceedings of the 5th International Conference and Exhibition on Multi-Lingual Computing, pages 7/411–7/414 (of xii + 582). Cambridge University, Cambridge, UK, 1996.

Weitz:2007:UUM

[Wei07]  Jay Weitz.  UNIMARC

and UNICODE — The 18th Meeting of the Permanent UNIMARC Committee 12–14 March 2007, National Library of Portugal, Lisbon. International cataloguing and bibliographic control: quarterly bulletin of the IFLA UBCIM Programme, 36(3):66, ???? 2007. ISSN 1011-8829.

Welch:1996:PIF

[Wel96] J. Welch. Platform independent font support. In Unicode Consortium [Uni96c], pages A3– ?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Welch:1997:UNIb

[Wel97a] J. Welch. Unicode and the new information appliances. In Unicode Consortium [Uni97b], pages B5–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Welch:1997:UNIa

[Wel97b] Jim Welch. Unicode and the new information appliances. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Wells:2001:ODM

[Wel01] J. C. Wells. Orthographic diacritics and multilingual computing. Language Problems & Language Planning, 24(3):249–272, February 2001. CODEN ???? ISSN 0272-2690.

Whistler:2004:UCM

[WF04] Ken Whistler and Asmus Freytag. Unicode conformance model. Unicode Technical Report 33, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org,  October

27, 2004. URL http:/

/www.unicode.org/reports/
tr33/tr33-1.html;   http:
//www.unicode.org/unicode/
reports/tr33/.

Wenyin:2008:EHO

[WFD08] L. Wenyin, A. Y. Fu, and X. Deng. Exposing homograph obfuscation intentions by coloring Unicode strings. In Yanchun Zhang, editor, Asia–Pacific web conference: Progress in WWW research and development: AP-Web 2008, volume 4976 of Lecture Notes in Computer Science, pages 275–286. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-78848-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Wien:2005:BUU

[WG05] Erik Wien and Lars Olav Gigstad. Boost.Unicode: a Unicode library for C++ = Boost.Unicode: et Unicode klassebibliotek for C++. Studentarbeid, Høgskolen i Gjøvik, Gjøvik, Norway, 2005.

Whistler:2005:SAE

[Whi05a] Ken Whistler. Sample American English translation of Unicode names list. Unicode Technical Note 24, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, July 15, 2005.

URL http://www.unicode.
org/notes/tn24/;   http://
www.unicode.org/notes/tn24/
tn24-1.html.

Whistler:2005:UNC

[Whi05b] Ken Whistler. Unicode named character sequences. Unicode Standard Annex 34, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, March 25,
2005.  URL http://www.
unicode.org/reports/tr34/
tr34-1.html;   http://www.
unicode.org/reports/tr34/
tr34-3.html;   http://www.
unicode.org/unicode/reports/
tr34/.

Whistler:2006:ELG

[Whi06] Ken Whistler. On the encoding of Latin, Greek, Cyrillic, and Han. Unicode Technical Note 26, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, February 14, 2006.

                                        URL http://www.unicode.
                                        org/notes/tn26/;   http://
                                        www.unicode.org/notes/tn26/
                                        tn26-1.html.

Whistler:2011:UCC

[Whi11] Ken Whistler. Unicode character categories. Technical Report 49, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

                                        mail: unicode-inc@unicode.

org, July 12, 2011. URL http:

                                        //www.unicode.org/reports/
                                        tr49/.

Wissink:2001:UNR

[Wis01]  Cathy Wissink.  Unicode

and the next release of Windows. In Unicode Consortium [Uni01a], page ?? ISBN

????  LCCN ????  URL

                                        http://www.microsoft.com/
                                        globaldev/presentations/Unicode18/
                                        uni+wxp.doc.

Wissink:2002:IIL

[Wis02] Cathy Wissink. Issues in Indic language collation. Unicode Technical Note 1, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, April 5, 2002.

URL http://www.unicode.
org/notes/tn1/;   http://
www.unicode.org/notes/tn1/
Wissink-IndicCollation.pdf.

Wissink:2001:UDW

[WK01] Cathy Wissink and Michael Kaplan. Unicode on downlevel Windows. In Unicode Consortium [Uni01a], page ?? ISBN ???? LCCN ???? URL http://

www.unicode.org/iuc/iuc18/
papers/b5.ppt.

Wahl:1997:RLD

[WKH97] M. Wahl, S. Kille, and T. Howes. RFC 2253: Lightweight Directory Access Protocol (v3): UTF-8 string representation of distinguished names, December 1997. URL ftp://ftp.

internic.net/rfc/rfc2253.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc2253.txt. Sta-

tus: PROPOSED STAN-DARD.

Wolf:1996:RCS

[Wol96] M. Wolf. Reuters compression scheme for Unicode (RCSU). In Unicode Consortium [Uni96c], pages B3–?? ISBN ???? LCCN QA268 .I56 1996. URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Wolf:1997:RCS

[Wol97] Misha Wolf. Reuters Compression Scheme for Unicode (RCSU). In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc10/program.html.

Wong:1998:DMJ

[Won98] Simon K. Wong. Developing multilingual Java applications for Oracle8. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc13/program.html.

Wood:1999:UMS

[Woo99] Alan Wood. Unicode and multilingual support in HTML, fonts, Web browsers and other applications. World-Wide Web document., February 3, 1999.

                                        URL http://www.alanwood.

net/unicode/. This site is updated frequently.

Wood:2010:ACP

[Woo10] James Wood. ABAP cookbook: programming recipes for everyday solutions. Galileo Press, Boston, MA, USA, 2010. ISBN 1-59229-326-3. 548 pp. LCCN QA76.73.A12 W66 2010.

Wright:2015:JAU

[Wri15] Joseph Wright. Joseph’s adventures in Unicodeland. TUGboat, 36(2):114–116, 2015. ISSN 0896-3207. URL https:

                                        //tug.org/TUGboat/tb36-2/
                                        tb113wright-unicode.pdf.

Wright:2020:CCT

[Wri20] Joseph Wright. Case changing: From TEX primitives to the Unicode algorithm. TUGboat, 41(1):39–42, ???? 2020. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb41-1/
tb127wright-case.pdf.

Wright:2022:CCL

[Wri22] Joseph Wright. Case changing: LaTEX reaches Unicodeland. TUGboat, 43(2):165– 166, 2022. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb43-2/
tb134wright-case.pdf.

Wright:2023:SUD

[Wri23]  Joseph Wright.  Storing

Unicode data in TEX engines. TUGboat, 44(1):134– 137, 2023. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb44-1/
tb136wright-unidata.pdf.

Wu:2000:TNS

[Wu00] Pei-Chi Wu. Transition from national standards to Unicode: multilingual support in operating systems and programming languages. Software—Practice and Experience, 30(7):765–774, June 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72001831/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72001831&PLACEBO=IE.pdf.

Wu:2001:BTF

[Wu01] Pei-Chi Wu. A base62 transformation format of ISO 10646 for multilingual identifiers. Software—Practice and Experience, 31(12):1125–1130, ???? 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85008992/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85008992&PLACEBO=IE.pdf.

Wu:2002:PST

[Wu02]  Pei-Chi Wu.  A page-shift

transformation format of ISO 10646. Software—Practice and Experience, 32(1):73–82, ???? 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88013124/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88013124&PLACEBO=IE.pdf.

Unicode:1991:HNS

[WWW^{+}91] Misha Wolf, Ken Whistler, Charles Wicksteed, Mark Davis, and Asmus Freytag. Handling non-spacing marks. Unicode Technical Report 5, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, 1991. Now part of The Unicode Standard, Version 2.0.

Wolf:1997:SCS

[WWW^{+}97] Misha Wolf, Ken Whistler, Charles Wicksteed, Mark Davis, and Asmus Freytag. A standard compression scheme for Unicode. Unicode Technical Standard 6, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-

mail: unicode-inc@unicode.

org, May 30, 1997. URL http:

//www.unicode.org/unicode/
reports/tr6/index.html. This

version supersedes earlier versions with the same report number. Version 2.0 is dated 30 September 1998.

Yang:1998:UDC

[Yan98] Jianping Yang. User defined character set support in Unicode. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Yasuoka:2005:AJV

[Yas05] K. Yasuoka. Abode-Japan1-6 vs. Unicode-character codes in Japan. Joho kanri = Journal of information processing and management., 58(8):487– 495, ???? 2005. ISSN 0021-7298.

Yergeau:1996:RUT

[Yer96] F. Yergeau. RFC 2044: UTF-8, a transformation format of Unicode and ISO 10646, October 1996. URL ftp://ftp.

                                        internic.net/rfc/rfc2044.
                                        txt;   ftp://ftp.math.utah.
                                        edu/pub/rfc/rfc2044.txt. Sta-

tus: INFORMATIONAL.

Yergeau:1998:RUT

[Yer98a]  F. Yergeau.  RFC 2279:

UTF-8, a transformation format of ISO 10646, January

                                        1998.  URL ftp://ftp.
                                        internic.net/rfc/rfc2044.
                                        txt;   ftp://ftp.internic.
                                        net/rfc/rfc2279.txt;   ftp:
                                        //ftp.math.utah.edu/pub/rfc/
                                        rfc2044.txt;   ftp://ftp.
                                        math.utah.edu/pub/rfc/rfc2279.

txt. Obsoletes RFC2044 [Yer96]. Status: PROPOSED STANDARD.

Yergeau:1998:LEU

[Yer98b] François Yergeau. Language engineering with Unicode. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

RFC3629

[Yer03]  F. Yergeau.  RFC 3629:

UTF-8, a transformation format of ISO 10646, November 2003. URL ftp://ftp.

                                        internic.net/rfc/rfc3629.
                                        txt;   ftp://ftp.math.utah.
                                        edu/pub/rfc/rfc3629.txt.

Yu:2005:MXD

[YLM^{+}05] Yijun Yu, Jianguo Lu, John Mylopoulos, Weiwei Sun, Jing-Hao Xue, and Erik H. D’Hollander. Making XML document markup international. Software—Practice and Experience, 35(1):1– 14, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yergeau:1997:RIH

[YNAD97] F. Yergeau, G. Nicol, G. Adams, and M. Duerst. RFC 2070: Internationalization of the hypertext markup language, January 6, 1997. URL ftp://ftp.

internic.net/rfc/rfc2070.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc2070.txt. This

Internet Request-for-Comment document describes an extension of the HTML 2.0 DTD to support the Unicode character set.

Yong:1997:MYW

[Yon97] Leong Kok Yong. Making your Web server render Unicode text for your client users. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Youatt:1998:MIP

[You98] Richard W. Youatt. Multilingual information processing in the Republic of Armenia. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

YKJJHIH:2000:YKJ

[Yun00] Yunikodo Kanji Joho Jiten Henshu Iinkai hen. Yunikodo kanji joho jiten = Sanseido’s Unicode kanji information dictionary. Sanseido, Tokyo, Japan, 2000. ISBN 4-385-13690-4. xvi + 467 + 122 pp. LCCN QA268 .Y86 2000.

Yang:1997:CSD

[YY97] Jianping Yang and Gail Yamanaka. Character set definition and conversion in Unicode. In Unicode Consortium [Uni97a], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc10/
                                        program.html.

Zhang:1997:LCE

[Zha97a] Z. Zhang. The Large-Scale Chinese Electronic Publications (LSCEP) with Unicode. In Unicode Consortium [Uni97b], pages B16–?? ISBN ????

                                        LCCN ???? URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Zhang:1997:NPC

[Zha97b] Z. Zhang. New progress of CJK encoding with ISO/IEC 10646 and Unicode. In Unicode Consortium [Uni97b], pages TB4–

??  ISBN ????  LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc11/
                                        program.html.

Ziff:2001:DUC

[Zif01] Donald Ziff. Documentum and UTF-8: Converting content management software product line to Unicode. In Unicode Consortium [Uni01a], page ??

ISBN ???? LCCN ????

URL http://www.unicode.
org/iuc/iuc18/papers/a17.
doc.

Zajac:1997:GBM

[ZV97] R. Zajac and M. Vanni.

Glossary-based MT engines in a multilingual analyst’s workstation architecture. Machine Translation, 12(1–2):131–151, ???? 1997. CODEN MACTEZ. ISSN 0922-6567.